### v2

# A new paradigm of mortality modeling via individual vitality dynamics

Abstract
The significance of mortality modeling extends across multiple research areas, including life insurance valuation, longevity risk management, life-cycle hypothesis, and retirement income planning. Despite the variety of existing approaches, such as mortality laws and factor-based models, they often lack compatibility or fail to meet specific research needs. To address these shortcomings, this study introduces a novel approach centered on modeling the dynamics of individual vitality and defining mortality as the depletion of vitality level to zero. More specifically, we develop a four-component framework to analyze the initial value, trend, diffusion, and sudden changes in vitality level over an individual’s lifetime. We demonstrate the framework’s estimation and analytical capabilities in various settings and discuss its practical implications in actuarial problems and other research areas. The broad applicability and interpretability of our vitality-based modeling approach offer an enhanced paradigm for mortality modeling.
1 Introduction
Mortality modeling is a critical research topic, serving as the foundation of numerous research questions and applications in various fields. In actuarial science, accurate mortality models are essential for understanding and managing risks associated with life insurance products. They enable life actuaries to assess the underlying mortality risk of insurance policies and ensure the financial stability of life insurance companies. In the realm of long-term care and disability insurance, mortality models are also contributing as an essential tool, for example, in a multi-state setup to capture the complexities of transition probabilities and claim distributions.
Beyond traditional insurance applications, mortality modeling plays an important role in retirement income planning as well. By estimating the survival time of individuals, mortality models can help to obtain optimal consumption and investment strategies, allowing for better financial planning and resource allocation throughout retirement. The ability to predict longevity and mortality trends is also crucial in pension fund management, public health planning, and demographic studies. In each of these areas, the reliability and interpretability of mortality modeling outcomes directly impact the quality of the decisions made.
In the literature, a variety of mortality modeling approaches exist, each with its strengths and limitations. Mortality laws are among the oldest and most fundamental approaches. These laws, such as the Gompertz-Makeham law (Gompertz,, 1825; Makeham,, 1860), aim to describe age-specific mortality rates using simple mathematical functions. Mortality laws are widely appreciated for their simplicity and ability to capture the general trend of increasing mortality with age. Despite their historical significance and ease of use, these laws often fall short in capturing the complexities of real-life mortality data, especially in the presence of cohort effects and changing mortality trends over time. More recent studies on this approach include, for example, Gavrilov and Gavrilova, (2019), Li et al., 2021a , Di Palo, (2023) and Dang et al., (2023).
In contrast to traditional mortality laws, the Cox proportional hazards model (Cox,, 1972) represents a significant advancement by not assuming a specific baseline hazard function. This semi-parametric model allows the hazard function to vary over time, influenced by covariates such as demographic, health-related, or environmental factors. The flexibility of the Cox model in incorporating various risk factors makes it a powerful tool in survival analysis and epidemiology. However, one of its limitations is the proportional hazards assumption, which implies that the covariates have a multiplicative effect on the hazard function. This assumption may not hold in all cases, leading to potential biases in the estimated effects of covariates. A comprehensive review of the Cox model and its recent developments can be found in Kalbfleisch and Schaubel, (2023).
Another category is developed from using continuous-time stochastic processes to describe the evolution of mortality intensities. The seminal work by Milevsky and Promislow, (2001) applied a continuous-time Cox model to survival analysis. Biffis, (2005) introduced the Markovian affine jump-diffusion process for modeling both financial and demographic risks. This model preserves analytical tractability and provides closed-form solutions to survival probabilities. Blackburn and Sherris, (2013) further validated the application of Markovian affine mortality models in longevity risk pricing and hedging. Bauer et al., (2010) focused on pricing longevity-linked securities using a forward force of mortality modeling approach. Other studies include, for example, Wong et al., (2017), Jevtić and Regis, (2019), Wang et al., (2021) and Zhou et al., (2022).
Factor-based mortality models, leading by the seminal works of Lee and Carter, (1992) and Cairns et al., (2006), have become prominent due to their ability to capture underlying trends and volatility in mortality data. The Lee-Carter (LC) model and the Cairns-Blake-Dowd (CBD) model express mortality rates as a combination of age-specific components and time-varying factors to capture mortality changes over time. Significant advancements have been made in this direction, including the works of Haberman and Renshaw, (2012), Kleinow, (2015), Shang, (2019), and Hunt and Villegas, (2023), with a review provided by Cairns et al., (2009). Despite their popularity, these models are based on discrete-time settings, which may not be suitable for continuous-time research questions requiring explicit solutions.
Similarly, spline-based mortality models offer a flexible non-parametric approach to mortality modeling by fitting smooth curves to mortality data. These models use splines, which are piece-wise polynomial functions, to capture the intricate patterns of mortality rates. Currie et al., (2004) demonstrated that the use of P-splines for smoothing mortality data provides a robust method for handling irregularities and fluctuations. Other studies include, for example, Dodd et al., (2018), Camarda, (2019) Hilton et al., (2019), Richards, (2020), Dodd et al., (2021), and Zhu and Zhou, (2023). The main advantage of spline-based models is their flexibility in fitting complex mortality patterns. However, their non-parametric nature can lead to overfitting and a lack of interpretability when compared to more structured models.
Recently, an innovative approach of modeling mortality via an individual’s remaining survival energy has been proposed, distinguishing itself from traditional models that typically focus on populations or cohorts. Specifically, Shimizu et al., (2021) introduced the Survival Energy Model (SEM), suggesting that each individual is born with a certain initial level of survival energy that depletes stochastically over time through a time-inhomogeneous diffusion process. Building on this, Shimizu et al., (2023) expanded the SEM framework by incorporating more sophisticated stochastic processes to better capture the complexities of mortality dynamics. By focusing on the stochastic nature of an individual’s survival energy, the SEM provides a more intuitive understanding of mortality dynamics.
We note that within the SEM framework, an individual’s death time is essentially the first time that the survival energy process reaches zero. Therefore, this modeling approach is closely connected to the field of first passage analysis, which is a well-established topic in applied probability with extensive applications in various fields such as operations management (see, e.g., Yamazaki, (2017) and Han et al., (2024)), quantitative finance (see, e.g., Davydov and Linetsky, (2001) and Cai and Kou, (2011)) and actuarial science (see, e.g., Lin and Willmot, (2000) and Li and Zhou, (2024)). We show in this paper that the depletion of an individual’s vitality (a.k.a., survival energy) exhibits similarities to the dynamics of an insurer’s cash flow. Specifically, the negative jumps in the vitality process represent decreases in vitality level due to accidents, analogous to how an insurer’s surplus process often takes a spectrally negative form to model claim payments.
The concept of modeling individual vitality is not new beyond actuarial literature. It was pioneered by Strehler and Mildvan, (1960), who defined vitality as “the capacity of an individual organism to stay alive”. For a discussion on the application of the Strehler-Mildvan model to human mortality, please refer to Finkelstein, (2012) and the references therein. Anderson, (1992) further contributed by incorporating Brownian motion to construct a stochastic vitality process, transforming the study of mortality into a first-passage time problem. Numerous studies on mortality modeling have been conducted in this direction, including the works by Skiadas, (1995); Anderson, (2000); Weitz and Fraser, (2001); Aalen and Gjessing, (2004); Steinsaltz and Evans, (2004); Li and Anderson, (2009); Li et al., (2013); Li and Anderson, (2013); Skiadas and Skiadas, (2010, 2014); Li and Anderson, (2015); Sharrow and Anderson, (2016), and references therein.
The existing studies demonstrate that vitality modeling is a robust approach to understanding mortality dynamics. The recent consideration of survival energy models in actuarial science highlights their potential to offer new insights into mortality modeling and forecasting. By integrating the stochastic nature of individual vitality and considering both deterministic and stochastic components, vitality models can provide a comprehensive framework for mortality modeling that is both flexible and intuitive. The main objective of this paper is to show that vitality modeling can become a valuable tool in actuarial science, addressing a number of research problems and enhancing the solutions provided by existing methods. Moreover, vitality modeling offers an alternative, more intuitive understanding of mortality dynamics, which can complement and improve upon traditional actuarial approaches.
The first contribution of this paper is the introduction of a vitality-based framework for mortality modeling. This framework conceptualizes mortality through the lens of individual vitality, where death is defined as the point at which an individual’s vitality depletes to zero. The fundamental idea is that each person is born with an initial level of vitality, which decreases over time due to both natural aging and external factors. The framework is built around four intuitive components: the initial vitality level, natural depletion trends, stochastic diffusion process, and sudden jumps. The resulting framework offers a comprehensive and flexible approach to mortality modeling and is capable of accommodating various features and ideas related to mortality dynamics. This flexibility sets the stage for our second contribution, where we demonstrate how different specifications of the proposed framework can lead to diverse and insightful modeling outcomes.
The second contribution of this paper lies in the various specifications of the proposed modeling framework, which is capable of representing existing mortality models, including the Gompertz law and the Cairns-Blake-Dowd model. By transforming these models into a unified vitality-based structure, we enhance their ability to capture more complex mortality dynamics, such as jumps to account for accidental deaths and diffusion to reflect the randomness in natural life decay. Under each specification, we derive survival probabilities and other relevant properties to showcase the insightful modeling outcomes, setting the stage for our third contribution on actuarial applications.
The third contribution of this paper demonstrates the practical applications of the vitality-based modeling framework in actuarial science. We showcase that this framework can be employed in various actuarial tasks and has the potential to solve existing problems and improve traditional methods. Specifically, we discuss how vitality can be incorporated in calculating life expectancies, pricing life insurance products, and deriving optimal consumption decisions. Additionally, our vitality-based framework is utilized to model disability and recovery probabilities, analyze different causes-of-death, and explain the under- and over-estimation of survival probabilities throughout an individual’s life course.
The structure of this paper is organized as follows. Section 2 details the proposed modeling framework, explaining its four key components: initial vitality level, natural depletion trends, stochastic diffusion processes, and sudden jumps. Section 3 covers the various specifications of the proposed framework, including static and dynamic cases, and offers specific examples to illustrate its estimation and analytical capabilities. Section 4 discusses the applications of the framework in actuarial science, such as life insurance valuation, disability modeling, and optimal consumption decisions. Section 5 presents an illustration on the numerical estimation of a vitality-based mortality model. Finally, Section 6 concludes the paper with a discussion on the limitations and suggestions for future research.
2 The Proposed Modeling Framework
In this section, we introduce the vitality-based mortality modeling framework, detailing its fundamental structure and components. The proposed framework conceptualizes mortality through the depletion of an individual’s vitality to zero and utilizes four components to describe the dynamics of the underlying vitality. Each component captures different aspects of vitality dynamics, jointly providing a comprehensive and flexible approach to mortality modeling.
Consider an individual who is age $x$ at time 00 (the time of modeling), and let $V(t)$ be the vitality of this individual at time $t\geq 0$. The individual is considered deceased when the vitality level is exhausted (i.e., when $V(t)=0$). Let $\tau$ be the first passage time of the vitality $V(t)$ crossing zero; that is,
$\tau=\inf\left\{t\geq 0:V(t)\leq 0\right\}.$
Then, the $T$-year survival probability for the individual is $\Pr(\tau>T)$, for which we aim to derive an expression in this paper.
Our modeling framework describes the dynamics of $V(t)$ as
$V(t)=V(0)-Y(t)-W(t)-J(t),\quad t\geq 0,$ (1)
where $V(0)$, $Y(t)$, $W(t)$ and $J(t)$ are the four components with the following interpretations:
  1. 1.
    $V(0)$ represents the vitality level at time 00, also called the initial vitality.
  2. 2.
    $Y(t)$ represents the trend at which the vitality level is naturally depleted.
  3. 3.
    $W(t)$ represents the diffusion underlying the depleting rate of the vitality level.
  4. 4.
    $J(t)$ represents the jump process that may affect the vitality level.
In the following subsections, We describe these components in detail and discuss their conceptual underpinnings and intuitive implications.
Component 1 (C1): The initial $V(0)$
The first component of our vitality-based modeling framework is the initial vitality of the individual being modeled, denoted as $V(0)$. This component is crucial as it sets the baseline from which all future vitality changes are measured. If the age $x$ is zero, then $V(0)$ represents the vitality level of a newborn. Otherwise, $V(0)$ can be interpreted as the remaining vitality level of the individual $x$ years after birth.
We can model $V(0)$ either as a fixed value or a random variable. A fixed value of $V(0)$ assumes that the level of initial vitality is known and also the same for all the individuals being modeled (i.e., all individuals who are age $x$ at time 00). This assumption might be reasonable if the cohort of individuals is homogeneous. For example, we can assume that all newborns from a certain birth year will have the exact same initial vitality. This simplification helps in creating a more streamlined model for analysis and is particularly useful in scenarios where there is little variation in initial health conditions within a cohort.
In a more realistic setting, the value of $V(0)$ should be treated as random to represent the inherent variability in the vitality level of individuals. This randomness reflects the diversity in individuals within a cohort, such as genetics, lifestyles, and other factors that affect an individual’s initial health condition. We should thus model $V(0)$ as a random variable with
$V(0)\sim F_{0},$
where $F_{0}$ is a distribution that could be, for example, exponential or Pareto. This approach acknowledges that even within a seemingly homogeneous cohort, individuals will have different initial vitality levels due to various reasons.
Furthermore, the remaining vitality level after $x$ years of living will differ among individuals due to their personal lifestyle, living environment, health status, and other individual-specific factors. Even for newborns, the initial vitality should vary within an inhomogeneous cohort to reflect different birth traits, such as genetic disorders, birth defects, and pregnancy conditions. This variability is essential for constructing a model that can accurately reflect individual differences, and also for building a linkage between existing mortality models and vitality-based models.
Component 2 (C2): The trend $Y(t)$
The second component describes the decline in an individual’s vitality over time due to the natural aging process. We denote $Y(t)$ to represent the depletion trend in vitality that occurs as the individual ages beyond time 00. To reflect the general trend of vitality depletion in the absence of stochastic variations or sudden shocks, $Y(t)$ can be modeled as a deterministic function of time. In particular, we assume that
$\mathop{}\!\mathrm{d}Y(t)=\mu(t)\,\mathop{}\!\mathrm{d}t,\quad t\geq 0,$
where $\mu(t)$ is the depletion rate of vitality at time $t$ when the individual is aged $x+t$. The exact choice of $\mu(t)$ should be informed by both theoretical considerations and empirical studies.
The function $\mu(t)$ can be specified in various forms, depending on the specific characteristics and assumptions made. Common choices include constant, exponential, or more complicated functional forms that better fit empirical observations. For instance, an exponential decline might be used to represent a situation where the rate of vitality loss accelerates with age. Another example would be using the Gompertz law to specify $\mu(t)$; that is, $\mu_{x}(t)=bc^{x+t}$ where $b$ and $c$ are the conventional Gompertz parameters. We added a subscript of age $x$ to $\mu(t)$ to align with the typical notation of mortality laws and to emphasize that the depletion rate, given by the Gompertz law, depends on the individual’s age at time $t$.
Alternatively, we may assume that the depletion rate is uniform over time; that is, $\mu(t)=\delta$ with $\delta$ being a constant. Under this setting, the rate at which vitality depletes is independent of time $t$, which may be reasonable when the duration of $\mu(t)=\delta$ is short. Moreover, we may further assume a piece-wise function of $\mu(t)$, where the depletion rate is flat over each integer age of the individual; that is, $\mu_{x}(t)=\delta_{x+\lfloor t\rfloor}$. This assumption is equivalent to the constant force of mortality assumption frequently used in mortality modeling.
Component 3 (C3): The diffusion $W(t)$
Acknowledging that an individual’s life is full of uncertainties and thus vitality should not be depleted in a perfectly predictable manner, the third component introduces a certain degree of uncertainty to the vitality dynamics of the individual. From the modeling perspective, given that the trend component $Y(t)$ exhibits a deterministic nature, it is reasonable to consider a diffusion component, denoted as $W(t)$, to reflect the fact that the vitality level of an individual can fluctuate randomly around its depletion trend. This stochastic variation will account for any minor health fluctuations that an individual may experience over time.
One simple way to model $W(t)$ is to use a Brownian motion; that is,
$\mathop{}\!\mathrm{d}W(t)=\sigma(t)\mathop{}\!\mathrm{d}B(t),\quad t\geq 0,$
where $B(t)$ is a standard Brownian motion and $\sigma(t)$ is a diffusion parameter at time $t$. A more advanced stochastic process can be considered to replace $B(t)$, such as a fractional Brownian motion, for capturing the long-term dependence in the individual’s health status over time. This setting is particularly useful when the individual has health conditions with a lingering effect on future vitality. Similar to the depletion rate, a piece-wise function of $W(t)$ can also be considered to reflect that different life periods may face different levels of vitality fluctuations.
With the diffusion component, we can simulate stochastic paths of individual vitality over time to provide insights into the variability and uncertainty inherent in the aging process. The simulations will generate a variety of potential future vitality trajectories, from which we can observe how different individuals’ path diverge from their deterministic trend. These simulated paths particularly help in understanding the probabilistic nature of vitality depletion and are crucial for applications such as deriving personalized optimal consumption decisions.
Component 4 (C4): The jump $J(t)$
Another natural phenomenon that could affect an individual’s health status is the occurrence of random events, such as an automobile incident or a medical surgery. These events can cause sudden and significant changes to an individual’s vitality level, but cannot be captured by the depletion trend or the stochastic diffusion. To address this, the last component of our proposed modeling framework is the jump component, denoted as $J(t)$. This component is designed to incorporate the impact of accidental events into the vitality dynamics via a jump process; that is,
$J(t)=\sum_{i=1}^{N(t)}Z_{i},\quad t\geq 0,$
where $N(t)$ counts the number of accidental events that have occurred up to time $t$, and $Z_{i}$ represents the size of the vitality jump caused by the $i$-th accident. We can interpret $J(t)$ as the total amount of vitality changes due to accidental events until time $t$ with random occurrence times and severity.
A straightforward example for modeling $J(t)$ would be a compound Poisson process, where $N(t)$ follows a Poisson process and $Z_{i}$ follows a certain severity distribution. This modeling choice is suitable because a Poisson process naturally models the occurrence of random events, while the size of each jump can follow various distributions depending on the nature of the accidental events. For instance, $Z_{i}$ can be modeled by a normal distribution to represent symmetric vitality changes or an exponential distribution for severe asymmetric impacts. A positive value of $Z_{i}$ will lead to a vitality loss, while a negative value of $Z_{i}$ will indicate a positive effect on vitality, such as a successful medical intervention. In a simplified setting, we may assume that $Z_{i}=\infty$ to represent the event of a fatal accident, where the individual’s vitality drops below zero instantaneously.
The jump component explains the variability in mortality due to accidental events that are not accounted for by gradual trends or random fluctuations. Together with the initial, trend and diffusion components, our proposed vitality-based framework is capable of providing insights into how a wide range of lifetime phenomena contribute to human mortality, including natural deaths, random health fluctuations, accidental events, and pre-existing medical conditions. By specifying these four components differently, our framework offers flexible modeling outcomes that encompass many existing mortality models. These different specifications and their implications will be discussed in the next section.
3 Model Specifications
In this section, we provide a number of specific cases in which the proposed vitality-based framework is implemented to arrive at various mortality modeling outcomes. The proposed framework is designed to be modulable, meaning that the four components can be selectively chosen based on different modeling objectives. For instance, by controlling certain components, the framework can replicate traditional mortality laws, offering an alternative approach to basic mortality analysis. Conversely, by incorporating all components, the resulting model is capable of capturing more complex phenomena that influence mortality dynamics over time. Through the specific cases, we aim to demonstrate the capabilities of the vitality-based framework across different modeling needs and highlight its potential to enhance our understanding of mortality patterns.
3.1 Static cases
A vitality mortality model is static if it only governs the mortality dynamics over an individual’s lifetime but does not account for general changes in population mortality over time. We discuss the dynamic cases in the next subsection, where period and cohort effects are captured and reflected by a vitality mortality model.
3.1.1 Traditional mortality laws
Let us consider the following specification of the vitality-based mortality modeling framework:
  • C1
    The initial vitality $V(0)$ follows an exponential distribution with unit scale; that is,
    $V(0)\sim\text{Exp}(1).$
  • C2
    The trend component $Y(t)$ is governed by a functional form of $\mu_{x}(t)>0$, specified by a mortality law. For instance, if the mortality law is Gompertz, then we have $\mu_{x}(t)=bc^{x+t}$ and
    $\mathop{}\!\mathrm{d}Y(t)=bc^{x+t}\mathop{}\!\mathrm{d}t,\quad t\geq 0,$
    where $b$ and $c$ are the two Gompertz parameters.
  • C3
    No diffusion component is specified (i.e., $W(t)=0$ for all $t\geq 0$).
  • C4
    No jump component is specified (i.e., $J(t)=0$ for all $t\geq 0$).
Conditioning on the initial vitality $V(0)$, $V(t)$ is a non-increasing function of $t$ since $\mu_{x}(t)>0$. We thus have
$\operatorname{Pr}(\tau>T)=\operatorname{Pr}(V(T)>0)=\Pr\left(V(0)>\int_{0}^{T}\mu_{x}(t)\mathop{}\!\mathrm{d}t\right).$
Given that $V(0)\sim\operatorname{Exp}(1)$, the $T$-year survival probability can be expressed as
$\Pr(\tau>T)=\exp\left(-\int_{0}^{T}\mu_{x}(s)\mathop{}\!\mathrm{d}s\right).$ (2)
We remark that the survival probability provided by equation (2) is the same as the survival probability implied by the mortality law $\mu_{x}(s)$ for $s\in(0,T)$. The vitality-based modeling approach is able to reformulate traditional mortality laws from describing the force of mortality $\mu_{x}$ to modeling the depletion speed of vitality with a random initial vitality level. A similar approach has been considered in biology research. For instance, Li and Anderson, (2009) considered a normal distribution for $V(0)$, whereas Aalen and Gjessing, (2004) considered an exponential distribution for $V(0)$ to draw a connection between hazard rate models and vitality models.
3.1.2 Plateau death models
One drawback of traditional mortality laws is that the force of mortality at extreme old ages would increase indefinitely, which contradicts the plateau effect observed in real morality data. Our proposed framework is capable of reproducing the plateau effect by specifying the first two components as
  • C1
    The initial vitality $V(0)$ follows a Type-II Pareto distribution with shape parameter $\alpha$ and scale parameter $\phi$. The survival function of this Pareto distribution is
    $S(v)=\left(1+\frac{v}{\phi}\right)^{-\alpha}.$
  • C2
    The trend component $Y(t)$ is governed by a mortality law $\mu_{x}(t)$:
    $\mathop{}\!\mathrm{d}Y(t)=\mu_{x}(t)\mathop{}\!\mathrm{d}t,\quad t\geq 0.$
The resulting $T$-year survival probability of the above modeling settings is
$\Pr(\tau>T)=\Pr\left(V(0)>\int_{0}^{T}\mu_{x}(t)\mathop{}\!\mathrm{d}t\right)=\left(1+\frac{\int_{0}^{T}\mu_{x}(t)\mathop{}\!\mathrm{d}t}{\phi}\right)^{-\alpha}.$ (3)
In particular, if $\mu_{x}(t)$ follows the Gompertz law, then the resulting survival probability is known as the Gamma-Gompertz model (see, e.g., Vaupel et al.,, 1979; Missov,, 2013). In the Gamma-Gompertz model, a frailty term $Z$ is specified such that $\mu_{x}(t|Z)=Z\cdot\mu_{x}(t)$ and $Z$ follows a Gamma distribution with shape parameter $\alpha$ and rate parameter $\phi$. The resulting $T$-year survival probability of the Gamma-Gompertz model will coincide with equation (3), which is derived from our vitality mortality model (see Appendix A.1). Missov, (2013) has demonstrated that the Gamma-Gompertz model reflects the plateau effect such that the implied force of mortality converges to a constant as age increases.
An alternative specification for capturing the plateau effect and resulting in equation (3) for the $T$-year survival probability is
  • C1
    The initial vitality $V(0)$ follows an exponential distribution with unit scale:
    $V(0)\sim\text{Exp}(1).$
  • C2
    The trend component $Y(t)$ is given by
    $\mathop{}\!\mathrm{d}Y(t)=Z\cdot\mu_{x}(t)\mathop{}\!\mathrm{d}t,\quad t\geq 0,$
    where $\mu_{x}(t)$ is specified by a mortality law and $Z$ follows a Gamma distribution with shape parameter $\alpha$ and rate parameter $\phi$.
These two specifications have distinct interpretations. The first one describes the initial vitality level of the underlying population to be more dispersed and the reason for the death plateau is due to a larger proportion of the population exhibiting a large initial vitality level. The second one implies that the death plateau is indeed due to the heterogeneity in the depletion rate.
3.1.3 A jump-diffusion vitality model
The traditional mortality laws, reformulated under our vitality-based framework, suffer two clear drawbacks. First, although the trend component $Y(t)$ is able to capture the natural decay of an individual’s vitality level, any accidental death that may occur randomly during the individual’s lifetime is not included. Second, the natural vitality decay specified by $\mu_{x}(t)$ is deterministic, which means the randomness that existed in the individual’s health status over time is ignored.
To introduce randomness, Shimizu et al., (2021) and Shimizu et al., (2023) have considered applying a Brownian motion and an inverse-Gaussian process, respectively, to govern an individual’s vitality. However, these studies assume a deterministic initial vitality and an arbitrarily chosen trend component, which consequently fails to connect the vitality-based approach with traditional mortality laws. One disadvantage of using inverse Gaussian processes is that their implied infinite number of small jumps may not be suitable for characterizing the decrement of vitality levels due to accidents.
Under the proposed framework, we now consider adding a Brownian motion in the diffusion component to reflect random fluctuations in vitality dynamics, and a compound Poisson process in the jump component to reflect the possibility of accidental events. In particular, the following specification is used:
  • C1
    The initial vitality $V(0)$ follows a distribution $F_{0}$ with positive support:
    $V(0)\sim F_{0}.$
  • C2
    The trend component $Y(t)$ is governed by
    $\mathop{}\!\mathrm{d}Y(t)=\mu_{x}(t)\mathop{}\!\mathrm{d}t,\quad t\geq 0,$
    where $\mu_{x}(t)$ is specified by a mortality law (e.g., $\mu_{x}(t)=bc^{x+t}$ for the Gompertz law).
  • C3
    The diffusion component $W(t)$ is specified as
    $W(t)=\sigma B(t),\quad t\geq 0,$
    where $B(t)$ is a Brownian motion and $\sigma$ is the diffusion parameter.
  • C4
    The jump component $J(t)$ is specified as
    $J(t)=\sum_{i=1}^{N(t)}Z_{i},\quad t\geq 0,$
    where $N(t)$ is a Poisson process with intensity rate $\lambda(t)$, and $\{Z_{i}\}_{i\geq 1}$ is a sequence of i.i.d. positive random variables with distribution function $F_{Z}(\cdot)$.
It follows that the dynamics of $V(t)$ is governed by
$V(t)=V(0)-\int_{0}^{t}\mu_{x}(s)\mathop{}\!\mathrm{d}s-\sigma B(t)-\sum_{i=1}^{N(t)}Z_{i}.$ (4)
We assume that $V(0)$, $W(t)$, and $J(t)$ are independent of each other. The resulting $T$-year survival probability can be expressed as
$\Pr(\tau>T)=\int_{0}^{\infty}\Pr\left(v-\sigma B(t)-\sum_{i=1}^{N(t)}Z_{i}\geq\int_{0}^{t}\mu_{x}(s)\mathop{}\!\mathrm{d}s,\;\forall t\leq T\right)\mathop{}\!\mathrm{d}F_{0}(v).$ (5)
We would like to point out that deriving analytical expressions for the survival probability is generally challenging in this case, unless specific forms of $Y(t)$ are assumed such as linear functions. To elaborate further, the main difficulty lies in deriving the law of the first passage time of $B(t)$ over a moving boundary, which is a non-trivial problem in applied probability (see, e.g., Salminen, (1988)). Furthermore, deriving analytical results about the death time using renewal arguments may also not be applicable if the trend does not preserve the renewal structure of the vitality dynamics, for instance, when $Y(t)$ is modeled under the Gompertz law. In such cases, one may resort to numerical methods to approximate the survival probability of a jump-diffusion vitality model. In this paper, we employed the Monte Carlo method as suggested in Jin and Wang, (2017) to numerically evaluate the survival probability as indicated in Eq. (5), and provided the details in Appendix B.
3.1.4 An alternative specification
In Section 3.1.2, we illustrated two model specifications that equally describe the death plateau. It is not surprising that more than one specification of the vitality-based modeling framework can provide outcomes that align with mortality laws. In this subsection, we formulate an alternative specification that also connects with mortality laws.
The following specification is assumed:
  • C1
    The initial vitality $V(0)$ follows a Gompertz distribution with shape parameter $\eta$ and scale parameter being one. The distribution function of this Gompertz distribution is
    $F(v)=1-e^{-\eta\left(e^{v}-1\right)}.$
  • C2
    The trend component $Y(t)$ has a linear decay at the rate of $\delta$; that is,
    $\mathop{}\!\mathrm{d}Y(t)=\delta\mathop{}\!\mathrm{d}t,\quad t\geq 0,$
We can show that the resulting $T$-year survival probability is
$\Pr(\tau>T)=\exp\left(-\eta\left(e^{T\delta}-1\right)\right),$ (6)
where, to obtain the Gompertz law’s expression of survival probability, the Gompertz parameters $b$ and $c$ need to be parameterised by $\eta=\frac{bc^{x}}{\ln(c)}$ and $\delta=\ln(c)$.
To introduce randomness in the evolution of an individual’s vitality over time, we can additionally assume that
  • C3
    The diffusion component $W(t)$ is specified as
    $W(t)=\sigma B(t),\quad t\geq 0,$
    where $B(t)$ is a Brownian motion and $\sigma>0$ is the diffusion parameter.
  • C4
    The jump component $J(t)$ is specified as
    $J(t)=\sum_{i=1}^{N(t)}Z_{i},\quad t\geq 0,$
    where $N(t)$ is a Poisson process with intensity rate $\lambda(t)$, and $\{Z_{i}\}_{i\geq 1}$ is a sequence of i.i.d. positive random variables.
It follows that the dynamics of $V(t)$ is governed by
$V(t)=V(0)-\delta t-\sigma B(t)-\sum_{i=1}^{N(t)}Z_{i},\quad t\geq 0.$ (7)
One special case worth noting is when the jump size $Z_{i}=\infty$ (i.e., each jump represents a fatal event). In this case, we would have
$\Pr(\tau>T)$ $={e}^{-\int_{0}^{T}\lambda(t)\mathop{}\!\mathrm{d}t}\cdot\Pr\left(V(0)-\delta t-\sigma B(t)>0\text{ for all }t\in[0,T]\right)$
$={e}^{-\int_{0}^{T}\lambda(t)\mathop{}\!\mathrm{d}t}\cdot\int_{0}^{\infty}\Phi\left(\frac{v-\delta T}{\sigma\sqrt{T}}\right)-\exp\left(\frac{2\delta v}{\sigma^{2}}\right)\Phi\left(\frac{-v-\delta T}{\sigma\sqrt{T}}\right)\mathop{}\!\mathrm{d}F_{0}(v),$ (8)
where the last equation follows from the well-known law of the first passage time of a Brownian motion with linear drift (see, e.g., Siegmund, (1986)).
In general, conditioning on $V(0)$, $V(t)$ is a spectrally negative Lévy process (SNLP), and the Laplace transform of $\tau$ is known explicitly. This allows for the use of either analytical or numerical methods of Laplace inversion to obtain the survival probability. To illustrate this, in Appendix A.2, we provide related results for the vitality mortality model described by equation (7) when the jump distribution is a mixture of exponential distributions.
We remark that it might be convenient to work with the exponential transform of a vitality mortality model, such that we define the exponentially-transformed $\tilde{V}(t)$ as
$\tilde{V}(t)=\exp\left(V(t)\right)=\exp\left(V(0)-Y(t)-W(t)-J(t)\right),\quad t\geq 0,$
and the individual dies if $\tilde{V}(t)$ cross the threshold $D=1$,
$\tau=\inf\left\{t\geq 0:\tilde{V}(t)\leq 1\right\}.$
[…]
3.2 Dynamic cases
A vitality mortality model is called dynamic when it adapts to changes in mortality over time, thereby explaining the period and cohort effects of mortality data. In stochastic mortality modeling, the period effect refers to changes in mortality over time across ages, while the cohort effect refers to changes in mortality across birth years that remain consistent over the cohort’s lifetime. To capture period and cohort effects, time-varying parameters are needed in a vitality mortality model, similar to the structure of the LC model and the CBD model.
Similar to the static cases where we established a linkage between vitality models and mortality laws, in this subsection we aim to build a bridge between the dynamic vitality mortality models and the popular stochastic mortality models. In particular, we begin by examining a dynamic extension of the Gompertz model, which is connected to the CBD model. We then expand our investigation to other mortality models and also consider cohort effects. We demonstrate that the proposed vitality-based framework is capable of providing new insights and enhancements to existing models.
3.2.1 A dynamic vitality model
We start from a straightforward example, where the depletion rate of vitality is dynamically specified with time-varying parameters. In particular, we first construct a dynamic Gompertz law with the following specifications:
  • C1
    The initial vitality $V(0)$ follows an exponential distribution with unit scale; that is, $V(0)\sim\text{Exp}(1)$.
  • C2
    The trend component $Y(t)$ is specified by a dynamic version of the Gompertz law with time-varying parameters $b(t)$ and $c(t)$:
    $\mathop{}\!\mathrm{d}Y(t)=\mu_{x}(t)\mathop{}\!\mathrm{d}t=b(t)c(t)^{x+t}\mathop{}\!\mathrm{d}t,\quad t\geq 0.$
    The log transforms of $c(t)$ and $b(t)$ are governed, respectively, by
    $\mathop{}\!\mathrm{d}\ln c(t)$ $=\mu_{c}\mathop{}\!\mathrm{d}t+\sigma_{c}\mathop{}\!\mathrm{d}B_{c}(t),\quad c(0)=c_{0}>0,$
    $\mathop{}\!\mathrm{d}\ln b(t)$ $=\mu_{b}\mathop{}\!\mathrm{d}t+\sigma_{b}\left(\sqrt{1-\rho^{2}}\mathop{}\!\mathrm{d}B_{b}(t)+\rho\mathop{}\!\mathrm{d}B_{c}(t)\right),\quad b(0)=b_{0}>0,$
    where $B_{c}(t)$ and $B_{b}(t)$ are two independent Brownian motions.
For the Gompertz law, a dynamic version of $b(t)$ can be interpreted as the general level of mortality for the individual at time $t$ (i.e., the intercept of mortality curve), and $c(t)$ can be interpreted as the rate of changes in mortality when the individual ages at time $t$ (i.e., the slope of mortality curve). Under the vitality mortality model, we may alternatively interpret $b(t)$ as the general rate of vitality depletion at time $t$ and $c(t)$ as the change in depletion rate when the individual ages at time $t$.
The solution of $Y(t)$ is an integral of correlated log-normally distributed random variables
$Y(t)$ $=Y(0)+\int_{0}^{t}\exp\left(\mu_{Y}(s)+\sigma_{Y}(s)B(s)\right)\mathop{}\!\mathrm{d}s,\quad t\geq 0,$ (9)
where $B(s)$ is a standard Brownian Motion, and
$\mu_{Y}(t)$ $=\mu_{b}\cdot t+(x+t)\cdot\mu_{c}\cdot t+\ln b_{0}+(x+t)\cdot\ln c_{0},$
$\sigma_{Y}(t)$ $=\sigma_{b}^{2}+(x+t)^{2}\cdot\sigma_{c}^{2}+2(x+t)\cdot\rho\cdot\sigma_{b}\cdot\sigma_{c},$
Since $Y(t)$ is a non-decreasing process, the $T$-year survival probability can be expressed as
$\Pr(\tau>T)$ $=\mathbb{E}\left[\exp\left(-Y(T)\right)\right].$
The dynamic Gompertz law presented above is closely related to the Cairns-Blake-Dowd (CBD) model, which describes mortality over time and age for a population of individuals. A continuous version of the CBD model (applied to $\ln\mu_{x,t}$ instead of $\operatorname{logit}q_{x,t}$) can be constructed as follows:
$\ln\mu(x,t)$ $=\kappa_{1}(t)+\kappa_{2}(t)\cdot(x-\bar{x}),\quad t\geq 0.$
The $T$-year survival probability of an individual aged $x$ at time 00 under this continuous version of the CBD model is
$\mathbb{E}\left[\exp\left(-\int_{0}^{T}\mu(x+t,t)\mathop{}\!\mathrm{d}t\right)\right],$
where
$\mu(x+t,t)=\exp\left(\kappa_{1}(t))-\kappa_{2}(t)\cdot\bar{x}\right)\cdot\left[\exp\left(\kappa_{2}(t)\right)\right]^{(x+t)}.$
Clearly, if $\kappa_{1}(t)$ and $\kappa_{2}(t)$ are modeled as a two-dimensional random walk with drifts (see, e.g., Li et al., 2021b, ), then the parameters in the process of $\ln b(t)$ and $\ln c(t)$ can be chosen to match with the continuous CBD model as
$b(t)=\exp\left(\kappa_{1}(t))-\kappa_{2}(t)\cdot\bar{x}\right),\quad c(t)=\exp\left(\kappa_{2}(t)\right).$
The resulting $T$-year survival probabilities would be the same between our vitality mortality model and the continuous version of the CBD model.
Recall from equation (9) that $Y(t)$ can be expressed as an integral of log-normal random variables. Although the distribution of $Y(t)$ cannot be explicitly determined, accurate numerical approximations can be made to evaluate the survival probabilities. For example, Bégin et al., (2023) demonstrated that the integral of log-normal random variables can be well-approximated by a single log-normal random variable (see also, Lo, (2013)). The approximated moment generating function from Asmussen et al., (2016) is used to value the expectation $\mathbb{E}[\exp(-Y(t))]$.
Our dynamic vitality mortality model can be constructed to have the same survival probability as other stochastic mortality models, whenever the force of mortality is guaranteed to be non-negative. For instance, consider the Lee-Carter model:
$\ln\mu(x,t)=\alpha(x)+\beta(x)\kappa(t),\quad t\geq 0.$
If $\alpha(x)$ and $\beta(x)$ are specified as continuous deterministic functions and $\kappa(t)$ is governed by a continuous random walk with drift, then a dynamic vitality mortality model can be established to match the likelihood of the Lee-Carter model. Another example is the Cox-Ingersoll-Ross (CIR) model for modeling the log force of mortality, as discussed in Huang et al., (2022) and Ungolo et al., (2023), where the likelihood aligns with our vitality model if we maintain the same specification for the depletion rate $\mu_{x}(t)$.
While the likelihood of our vitality models can align with well-known stochastic mortality models, we believe that the proposed modeling approach offers significant potential and enhanced capabilities. An immediate observation is that the Gompertz law fails to differentiate mortality events between natural causes and external causes. When modeling the death rate for all causes, it is common to model $\kappa_{1}(t)$ (or $\ln b(t)$ in the vitality model) as a random walk with a negative drift, implying a stable improvement in the general mortality. However, if we decompose mortality into different causes of death, distinct historical trends will be observed.
Figure 1 displays the accidental death rates from 1999 to 2016 across different age groups, obtained from the U.S. Centers for Disease Control and Prevention. Unlike the all-cause death rate, which typically shows a downward trend, the accidental death rate exhibits an upward trend for all of the age groups considered. This finding has significant implications for accidental death insurance, in which distinguishing the effects of accidental and non-accidental deaths is crucial. Under the vitality modeling framework, we are able to address this issue by specifying the jump component (C4) with a time-varying intensity that increases over time.
Figure 1: The observed accident death rates from 1999 to 2016 across different age groups.
Considering the model setting introduced at the beginning of this section, where $Y(t)$ can be represented by an integral of log-normally distributed random variables, the $T$-year survival probability is modified by incorporating $J(t)$ which follows a compound Poisson process as defined previously:
$\Pr(\tau>T)$ $=\Pr(J(T)<V(0)-Y(T))$
$=\int_{0}^{\infty}\left(\int_{0}^{v}\Pr(J(T)\leq v-u)\mathop{}\!\mathrm{d}F_{Y}(u)\right)\mathop{}\!\mathrm{d}F_{0}(v)$
where $F_{Y}(\cdot)$ is the distribution functions for $Y(T)$. Since $F_{0}(v)$ and $\Pr(J(T)\leq v-u)$ can be explicitly expressed and $F_{Y}(\cdot)$ can be approximated as the distribution function of a single log-normal random variable, the integral for obtaining $\Pr(\tau>T)$ can be approximated numerically.
3.2.2 A dynamic vitality model with cohort effect
In a factor-based mortality model, the cohort effect is captured by an additive term, often denoted by $\gamma$, such as in the Renshaw-Haberman model (M2), the Age-Period-Cohort model (M3), the CBD model with cohort effect (M6), the CBD model with quadratic and cohort effects (M7), and the CBD model with age-dependent cohort effect (M8), as outlined in Cairns et al., (2009). In a vitality-based mortality model, the cohort effect can be naturally embedded in the initial vitality to reflect changes in the average healthiness of each generation.
We illustrate the idea of a dynamic vitality model with cohort effect based on the dynamic Gompertz law with cohort effect, which corresponds to the CBD model with cohort effect (M6). Denote $V_{y}(t)$ as the time-$t$ vitality level of an individual born in year $y$ with an age of $x$ at time 00. The following specification is assumed:
  • C1
    The initial vitality $V_{y}(0)$ for cohort year $y$ follows an exponential distribution with rate parameter $\Gamma(y)$, where
    $\frac{\mathop{}\!\mathrm{d}\Gamma(y)}{\Gamma(y)}=\mu_{\gamma}\mathop{}\!\mathrm{d}y+\sigma_{\gamma}\mathop{}\!\mathrm{d}B_{\gamma}(y),\quad\Gamma(0)=1,$ (10)
    where $B_{\gamma}(y)$ is a standard Brownian motion.
  • C2
    The trend component $Y(t)$ is specified as in the dynamic Gompertz model:
    $\mathop{}\!\mathrm{d}Y(t)=b(t)c(t)^{x+t}\mathop{}\!\mathrm{d}t,\quad t\geq 0.$
    The log transforms of $c(t)$ and $b(t)$ are governed by
    $\mathop{}\!\mathrm{d}\ln c(t)$ $=\mu_{c}\mathop{}\!\mathrm{d}t+\sigma_{c}\mathop{}\!\mathrm{d}B_{c}(t),\qquad c(y)=c_{0}(y),$
    $\mathop{}\!\mathrm{d}\ln b(t)$ $=\mu_{b}\mathop{}\!\mathrm{d}t+\sigma_{b}\left(\sqrt{1-\rho^{2}}\mathop{}\!\mathrm{d}B_{b}(t)+\rho\mathop{}\!\mathrm{d}B_{c}(t)\right),\quad b(y)=b_{0}(y),$
    where $B_{c}(t)$ and $B_{b}(t)$ are two independent standard Brownian motions, that are also independent to $B_{\gamma}(y)$.
The time-$t$ vitality level for an individual born in year $y$ and aged $x$ at time 00 is therefore $V_{y}(t)=V_{y}(0)-Y(t)$. The $T-$year survival probability can be expressed as
$\Pr(\tau_{y}>T)$ $=\mathbb{E}\left[\exp\left(-\Gamma(y)Y(T)\right)\right],$
where $\tau_{y}$ is the first passage time for $V_{y}(t)$ crossing zero.
The cohort effect is reflected by the rate parameter $\Gamma(y)$, which can be intuitively interpreted as the average initial vitality level at the time of modeling for individuals born in birth year $y$. The initial condition $\Gamma(0)=1$ ensures identifiability between $\Gamma(y)$ and $b(t)$. Note that the dynamic of $\Gamma(y)$ can be chosen to satisfy certain properties, for example, setting $\mu_{\gamma}=0$ ensures that the future cohorts are expected to be the same as the current one, or modeling $\Gamma(y)$ as a CIR process to incorporate a mean-reverting cohort effect.
We now establish a link between this dynamic vitality model and the continuous version of the M6 model. We fix age $x$ and construct the M6 model with time $t$ and cohort $y$ as
$\ln\mu_{y}(t)$ $=\kappa_{1}(t)+\kappa_{2}(t)\cdot(x+t-\bar{x})+\gamma(y).$
We can again re-parameterize $b(t)$, $c(t)$, and $\Gamma(y)$ in our vitality model to match the M6 model by having
$b(t)=\exp\left(\kappa_{1}(t))-\kappa_{2}(t)\bar{x}\right),\quad c(t)=\exp\left(\kappa_{2}(t)\right),\quad\Gamma(y)=\exp(\gamma(y)).$
The $T$-year survival probability under the above re-parameterization of the vitality model will coincide with the M6 model.
3.2.3 A dynamic vitality model with age-cohort effect
The dynamic vitality models presented in Sections 3.2.1 and 3.2.2 begin the modeling of vitality with a fixed initial age. It would be interesting to formulate a vitality model that allows the vitality to start from an arbitrary age. We now consider a dynamic vitality model with an age-adjusted cohort effect, extended again from the Gompertz Law.
The following specification is applied:
  • C1
    The initial vitality $V_{x,y}(0)$ for an individual born in $y$ and age $x$ at time $t=0$ follows
    $V_{x,y}(0)\sim\text{Exp}\left(\Gamma(y)\cdot a^{x}\right),$
    where $\Gamma(y)\cdot a^{x}$ is the rate parameter. We assume $a>1$ is a constant value accounting for the fact that vitality decreases as the individual ages, and $\Gamma(y)$ is governed by a stochastic process, for example, the Geometric Brownian Motion specified in Section 3.2.2.
  • C2
    The trend component $Y(t)$ is specified as
    $\mathop{}\!\mathrm{d}Y(t)=b(t)\frac{c(t)^{x+t}}{a^{x}}\mathop{}\!\mathrm{d}t,\quad t\geq 0,$
    with the dynamics of $b(t)$ and $c(t)$ the same as in Sections 3.2.1 and 3.2.2.
In the above specification, $\Gamma(y)$ is designed to capture the general level of initial vitality at birth for cohort year $y$, with the term $\Gamma(y)\cdot a^{x}$ reflecting the expected vitality level when the individual reaches age $x$. On the other hand, $b(t)$ and $c(t)$ are the Gompertz parameters that capture the change in vitality depletion rate when the individual is aging at time $t$. It is immediately observed that the resulting survival probability matches with the M6 model.
It is worth noting that this specification provides more flexibility in modeling the interactions between cohorts and ages. For example, the initial vitality component can be modified to
$V_{x,y}(0)\sim\text{Exp}\left(\Gamma(y)^{x_{c}-x}\right),$
where $x_{c}$ is a large and fixed value, e.g., $x_{c}=100$. If $\Gamma(y)>1$, the cohort effect will diminish quickly as the individual ages, which reconciles with the cohort effect of the M8 model.
4 Model Applications
The proposed vitality-based modeling framework has great potential in actuarial research. In this section, we provide a number of illustrative examples to showcase the applications of a vitality mortality model. Specifically, we examine the following topics:
  • Pricing life annuity and insurance products based on individual vitality;
  • Deriving optimal investment and consumption decisions using a vitality model;
  • Explaining the biased estimation of survival rates from the vitality perspective;
  • Modeling disability and recovery probabilities using vitality thresholds;
  • Analyzing the cause of death using vitality components.
Applying the vitality mortality model differs from existing mortality models not only in terms of technical details but also by offering new interpretations and insights that cannot be explained using existing mortality models.
4.1 Life insurance valuation
The survival probabilities of a vitality mortality model often cannot be explicitly expressed, or are overly complicated to be used in practice. However, for the purpose of calculating life expectancies and pricing life insurance products, it is convenient to work with the death time $\tau(v)$, defined as the first passage time of vitality crossing zero for an individual with age $x$ and initial vitality $v$.
The life expectancy of an individual with given initial vitality $v$ can be expressed as
$\overset{\circ}{e}_{x}(v)=\mathbb{E}[\tau(v)]=\int_{0}^{\infty}tf_{\tau}(t;v)\mathop{}\!\mathrm{d}t,$
where $f_{\tau}(t;v)$ is the density function of $\tau(v)$. Using the specification from Section 3.1.1 (i.e., the Gompertz law), the death time is
$\tau(v)$ :=$:=\left\{t\geq 0:v-\frac{b}{\ln c}c^{x}(c^{t}-1)=0\right\}$
$=\frac{1}{\ln c}\ln\left(\frac{v\ln c}{bc^{x}}+1\right),$ (11)
and the life expectancy for a population of individuals with initial vitality $V(0)\sim\text{Exp}(1)$ is
$\overset{\circ}{e}_{x}=\mathbb{E}\left[\tau(V(0))\right]=\int_{0}^{\infty}\frac{1}{\ln c}\ln\left(\frac{v\ln c}{bc^{x}}+1\right)e^{-v}\mathop{}\!\mathrm{d}v.$
Similarly, the price of a life annuity with \$1 per year payable continuously to an individual with age $x$ and initial vitality $v$ is
[…]
where $\bar{a}_{\hbox{\vtop{\vbox{\hrule\kern 1.0pt\hbox{$\scriptstyle t$\kern 1.0pt}}\kern 1.0pt}\vrule\kern 1.0pt}}=\frac{1-e^{-\delta t}}{\delta}$ with $\delta$ being the force of interest rate. The annuity price for a population of individuals with initial vitality $V(0)\sim\text{Exp}(1)$ is
$\bar{a}_{x}$ $=\mathbb{E}\left[\bar{a}_{\hbox{\vtop{\vbox{\hrule\kern 1.0pt\hbox{$\scriptstyle\tau(V(0))$\kern 1.0pt}}\kern 1.0pt}\vrule\kern 1.0pt}}\right]=\int_{0}^{\infty}e^{-v}\int_{0}^{\infty}\bar{a}_{\hbox{\vtop{\vbox{\hrule\kern 1.0pt\hbox{$\scriptstyle t$\kern 1.0pt}}\kern 1.0pt}\vrule\kern 1.0pt}}f_{\tau}(t;v)\mathop{}\!\mathrm{d}t\mathop{}\!\mathrm{d}v.$
Using the relation between annuity and insurance, we immediately have $\bar{A}_{x}(v)=\mathbb{E}\left[e^{-\delta\tau(v)}\right]=1-\delta\bar{a}_{x}(v)$ and $\bar{A}_{x}=1-\delta\bar{a}_{x}$.
If the density function of $\tau(v)$ can be explicitly derived, then the life expectancy, annuity price, and insurance price can also be expressed explicitly in an integral form. This is the case for our proposed model if the diffusion component is not considered. When the density function of $\tau(v)$ cannot be explicitly written, an approximation method is then needed to provide accurate estimates for life expectancy and prices. For example, consider a vitality mortality model with initial vitality, trend, and diffusion (i.e., components C1-C3), then the problem becomes finding the first exit density of Brownian motion with a curved boundary. A number of studies, including Jennen and Lerche, (1981), Jennen, (1985), Skiadas, (1995), and Durbin and Williams, (1992), has provided accurate and easy-to-implement approximation methods to handle this problem. Skiadas and Skiadas, (2010) and Skiadas and Skiadas, (2014) have specifically applied approximation methods to vitality modeling. For more details on approximating the density function of the death time when C4 is absent, interested readers are referred to Appendix D.
We remark that, under the vitality modeling approach, the evaluation of life expectancy and mortality-related products can be applied to specific groups of individuals, assuming that the vitality information can be obtained or estimated. For instance, by altering the distribution of initial vitality (or other components as well), the mortality of a selected healthy population or a disabled population can be analyzed. Consequently, depending on the population’s health status, a more accurate price can be produced for different life insurance products. This pricing approach, which tailors the vitality mortality model to reflect specific health conditions, allows for a more personalized experience in actuarial valuation. Additionally, this approach can be extended to assess the impact of lifestyle changes, medical advancements, and other public health interventions on life expectancy and insurance prices.
4.2 Optimal investment and consumption strategies
One of the potential applications of vitality modeling is lifetime planning, e.g., Merton’s portfolio problem. Individuals, when making decisions about their investment portfolio, consumption, insurance purchases, medical expenses, educational investment, etc., will take into account their health level at the time. The vitality mortality model provides a natural way to frame the optimization problem. There exist a few studies that utilize the vitality-based approach. For example, Chen et al., (2022) modeled vitality as a geometric Brownian motion to study the optimal insurance decision, and Cheng et al., (2023) modeled vitality using a jump process to study the optimal medical investment. In this subsection, we provide an illustration of Merton’s portfolio consumption problem under our vitality mortality model.
4.2.1 The market and vitality processes
Following the seminal work of Merton, (1969), we construct a continuous-time model for the financial market and apply the vitality model in Section 3.1.4. Denote $(\Omega,\mathbb{F},\mathbb{P})$ as a complete probability space with a right-continuous filtration generated by a two-dimensional standard Brownian Motion $(B_{S}(t),B_{V}(t))$. An individual allocates two representative assets, one risk-free $S_{0}(t)$ and one risky $S_{1}(t)$ with the following dynamics:
$\mathop{}\!\mathrm{d}S_{0}(t)$ $=rS_{0}(t)\mathop{}\!\mathrm{d}t,\quad S_{0}(0)>0,$
$\mathop{}\!\mathrm{d}S_{1}(t)$ $=(r+\theta\sigma)S_{1}(t)\mathop{}\!\mathrm{d}t+\sigma_{S}S_{1}(t)\mathop{}\!\mathrm{d}B_{S}(t),\quad S_{1}(0)>0,$
for $t\geq 0$, where $r$ is the risk-free rate, $\theta$ is the market price of risk, and $\sigma_{S}$ is the vitality of the risky asset.
Let $\pi(t)$ be the portfolio weight invested in the risky asset, and $\zeta(t)$ be the consumption amount. Then the asset $A(t)$ of an individual has the following dynamic
$\mathop{}\!\mathrm{d}A(t)=A(t)\left[r+\pi(t)\theta\sigma_{S}\right]\mathop{}\!\mathrm{d}t+\sigma\pi(t)A(t)\mathop{}\!\mathrm{d}B_{S}(t)-\zeta(t)\mathop{}\!\mathrm{d}t,\quad A(0)=A_{0}.$
We now adopt the vitality mortality model from Section 3.1.4, which recovers the Gompertz law if the initial vitality follows the Gompertz distribution; that is,
$\mathop{}\!\mathrm{d}V(t)=-\delta\mathop{}\!\mathrm{d}t+\sigma_{V}\mathop{}\!\mathrm{d}B_{V}(t),\quad V(0)=V_{0}>0,$
for $t\geq 0$, and we assume an individual is able to track their vitality level over time.
4.2.2 The optimal investment and consumption problem
An individual is maximizing their lifetime utility by choosing the optimal investment and consumption strategies. Therefore, for any time $t\geq 0$, we define the value function $J(A(t)=a,V(t)=v)$ as
$J(A(t)=a,V(t)=v)$ $=\max_{\pi,\zeta\in\Pi}\mathbb{E}_{t,a,v}\left[\int_{t}^{\tau}e^{-\beta(s-t)}\cdot u(\zeta(s))\mathop{}\!\mathrm{d}s+\lambda\cdot u(A(\tau))\right],$ (12)
where $\tau$ is the first passage time that vitality level crosses zero, $\beta>0$ is the time-preference rate, $u(\zeta)=\ln\zeta$ is the utility function (log-utility is selected for illustration), $\lambda\geq 0$ is the weight given to the utility of terminal asset (i.e., bequest), $\mathbb{E}_{t,a,v}=\mathbb{E}[\cdot|t,A(t)=a,V(t)=v]$, and the investment/consumption strategy $(\pi=(\pi_{t})_{t\geq 0},\zeta=(\zeta_{t})_{t\geq 0})\in\Pi$ belong to the admissible set $\Pi$ if $\pi$ and $\zeta$ are progressively measurable and the resulting asset process $(A_{t})_{t\geq 0}$ is non-negative for all $t$. Notice that the optimization is time-independent since $V(t)$ is stationary and thus the value function does not depend on the time $t$.
The optimization problem given in equation (12) can be solved by the Hamiltonian-Jacobi-Bellman equation, as discussed in Chen et al., (2022). The solution of the optimal investment strategy $\pi^{*}(t)$ and consumption strategy $\zeta^{*}(t)$ are summarized in the remark below, where the details of derivations are presented in Appendix C.
Remark 1.
The optimal investment and consumption strategies of Problem (12) are
$\pi^{*}(t)$ $=\frac{\theta}{\sigma_{S}},\quad\zeta^{*}(t)=\frac{A(t)}{f(V(t))}.$
where
$f(v)$ $=\left(\lambda-\frac{1}{\beta}\right)e^{k_{1}v}+\frac{1}{\beta},\quad\text{with}\quad k_{1}=\left\{\begin{array}[]{ll}\frac{\delta-\sqrt{\delta^{2}+2\sigma^{2}_{V}\beta}}{\sigma^{2}_{V}}<0,&\sigma_{V}>0\\ \frac{-\beta}{\delta},&\sigma_{V}=0\end{array}\right.$
As for comparison, we have Merton’s optimal portfolio and consumption strategies presented in Appendix C. The optimal investment strategy follows the same as Merton’s optimal portfolio in an infinite time horizon, and the consumption strategy also collapsed to Merton’s case (w/o mortality) if vitality is infinite.
4.2.3 Further remarks
It is not surprising to observe that the individual’s consumption depends on the vitality level, but what is interesting here is that they are not always negatively related, as one would expect. When the bequest motive is moderate or low (i.e., $\lambda<1/\beta$) a reasonable consumption profile is observed such that a higher vitality represents a long time to live with, and thus a lower consumption must be made to make sure that an individual will not run of out money before they are deceased.
However, when the bequest motive is significantly large (i.e., $\lambda>1/\beta$) decreasing the vitality actually increases the consumption, a counter-intuitive result at first glance. This is due to the fact that individuals are overly concerned about their bequest, and thus leads to an overly conservative consumption initially. Once they are aging and the time for death becomes more certain, they become more aware of their over-saving and may wish to increase their consumption a bit.
We emphasize that by incorporating the vitality mortality model into the optimal investment and consumption problem, new insights can be observed. Of course, our stylized model is overly simplified and can be extended in numerous ways. First, although an individual is able to track their health status, it is impossible to measure their vitality level accurately. To better reflect reality, noises should be introduced for the vitality level observed and lead to an optimal investment and consumption problem with partial information (for references on partial information, see for example, Lakner, (1995); Pham and Quenez, (2001); Bäuerle and Rieder, (2007); Ceci, (2012), etc.).
Second, ambiguous beliefs regarding survival rates can influence individual decisions. In the model described above, individuals may face uncertainty in realizing their initial vitality levels $V(0)$ and the depletion rates $\delta$. For those averse to uncertainty, life-cycle investment and consumption planning will differ significantly. The concept of model uncertainty in life-cycle planning was pioneered by Maenhout, (2004) and has more recently been applied to ambiguous mortality risk by Young and Zhang, (2016); Shen and Su, (2019) within the context of traditional mortality laws.
Third, it is possible to model medical expenditure as a function of vitality when studying the optimal consumption problem. Similarly, it is possible to express the vitality depletion rate $\mu_{x}(t)$ as a decreasing function of the medical expenditure when studying the optimal investment problem on personal health, see for example Forster, (1989); Hugonnier et al., (2013); Cheng et al., (2023), etc. One can explore numerous other directions where the optimization problem can be extended, highlighting the significant potential of the proposed vitality model in the area of lifetime planning.
4.3 Subjective beliefs
Numerous empirical studies have documented that individuals tend to have strong biases in their subjective beliefs about survival probabilities. In particular, individuals often underestimate survival probabilities at younger ages but overestimate them at older ages, as indicated in Ludwig and Zimper, (2013), Groneck et al., (2016), and Kalwij and Koc, (2021). This empirical observation can be easily explained and understood from the vitality perspective. Specifically, when a person underestimates their initial vitality and underestimates their depletion rate, they will eventually underestimate the vitality level (and hence survival probability) at younger ages and, as time goes on, overestimate the vitality level (and hence survival probability) at older ages.
In addition, empirical studies indicate that individuals update the probability associated with pessimistic and optimistic sentiments when confronted with health shocks. This phenomenon can also be explained under our vitality-based modeling framework by the overestimation of a vitality jump, which results in an underestimation of the survival probability after the shock. Similar to a recent study by Apicella and De Giorgi, (2024), our framework does not need to make specific assumptions about the difference in belief bias between different ages or the difference in optimism between different ages. In contrast to Apicella and De Giorgi, (2024), where the authors construct their model solely to explain the belief bias, the proposed framework is constructed to accurately reflect the vitality experience of each individual while retaining the flexibility to explain belief bias.
Another interesting observation that can be reflected by our vitality model is the distinction between subjective belief in health status and subjective belief in life expectancy. Specifically, if a person correctly self-assesses as having an average health level among the whole population, they may still over- or under-estimate their life expectancy if they falsely believe that their life expectancy will also be the average among the whole population. Using the specification from Section 3.1.1 with $b=0.0001744$ and $c=1.082$ as an illustrative example, the population average life expectancy at age 60 is 17 years. However, a 60-year-old individual who believes that they have an average initial vitality level (which is one since the initial vitality follows an exponential distribution with a mean of one) will expect to live for 20.4 more years as given by equation (4.1).
According to Jensen’s Inequality, we can show that the average life expectancy is smaller than the life expectancy of an individual with an average vitality level:
[…]
The above result suggests that even if an individual can objectively determine their health level relative to the population, it does not guarantee an unbiased estimate of their survival probabilities. However, if the individual judges their vitality level based on quantiles rather than the average (e.g., the median of the population vitality level), they will be able to arrive at an unbiased estimate of their life expectancy.
4.4 Disability modeling
It is natural to incorporate disability risk into our vitality model, similar to how mortality risk is incorporated. As the vitality level depletes, the risk of becoming disabled should increase over time. By setting a threshold above zero, we can classify an individual as disabled once their vitality level falls below this threshold. This extension allows the vitality model to capture the probability of both mortality and disability, and provide new insights into the interplay of mortality and morbidity risks.
4.4.1 Permanent disability model
Consider that there exists a threshold $\omega>0$ such that when the vitality falls below $\omega$, the person becomes disabled. Consider the specification with an exponentially distributed initial vitality (C1) and a Gompertz law specified depletion rate (C2). The resulting vitality model will form a permanent disability model, since $Y(t)$ is a strictly increasing function of time. The probability that a healthy person with initial vitality above $\omega$ will remain healthy for $T$ years is
$\Pr\left(V(0)-Y(T)>w\big{|}V(0)>w\right)$ $=\frac{\int_{0}^{\infty}e^{-v}\cdot\mathbb{1}_{v-Y(T)>\omega}\mathop{}\!\mathrm{d}v}{\int_{0}^{\infty}e^{-v}\cdot\mathbb{1}_{v>\omega}\mathop{}\!\mathrm{d}v}$
$=\frac{\int_{Y(T)+\omega}^{\infty}e^{-v}\mathop{}\!\mathrm{d}v}{\int_{w}^{\infty}e^{-v}\mathop{}\!\mathrm{d}v}=\exp(-Y(T)),$
which is equivalent to the survival probability, and the disability probability is simply $1-\exp(-Y(T))$. This result holds even if we assume that each individual has a different disability threshold, such that the threshold $\Omega$ has a density function $\pi(\omega)$. In this case, the probability of remaining healthy for $T$ years is
$\frac{\int_{0}^{\infty}e^{-v}\cdot\int_{0}^{\infty}\pi(\omega)\cdot\mathbb{1}_{v-\omega-Y(T)>0}\mathop{}\!\mathrm{d}\omega\mathop{}\!\mathrm{d}v}{\int_{0}^{\infty}e^{-v}\cdot\int_{0}^{\infty}\pi(\omega)\cdot\mathbb{1}_{v-\omega>0}\mathop{}\!\mathrm{d}\omega\mathop{}\!\mathrm{d}v}$ $=\frac{\int_{0}^{\infty}\pi(\omega)e^{-(\omega+Y(T))}\mathop{}\!\mathrm{d}\omega}{\int_{0}^{\infty}\pi(\omega)\cdot e^{-\omega}\mathop{}\!\mathrm{d}\omega}=\exp(-Y(T)).$
The above setting of the disability threshold $\omega$ can be extended in other directions, such as considering an age- and time-dependent threshold. We leave this for future research.
4.4.2 Disability model with recovery
Vitality models that have only an initial component (C1) and a non-decreasing trend component (C2) cannot accommodate the possibility of recovery if the disability is determined by the relative position of the vitality level with respect to a disability threshold. However, when the diffusion component (C3) is included, the model becomes capable of describing disability with the potential for recovery.
For illustrative purposes, we consider the alternative Gompertz specification from Section 3.1.4 with a diffusion component (C3) modeled as a Brownian motion multiplied by a constant volatility $\sigma$, and set a constant threshold $\omega$ for disability. The $T$-year recovery probability for a disabled person aged $x$ is
$\frac{\int_{0}^{\omega}\Pr\left(v-Y(T)-W(T)>\omega,\text{ and }v-Y(t)-W(t)>0\text{ for any }t\in[0,T]\right)\mathop{}\!\mathrm{d}F_{0}(v)}{1-F_{0}(\omega)}.$
Since $Y(t)=\delta t$, the probability inside in the integral can be solved explicitly by setting $\widehat{W}(t)=W(t)+\frac{\delta}{\sigma}t$, which is a Brownian Motion with drift $\frac{\delta}{\sigma}$. Denote
$\widehat{M}(T)=\max_{0\leq t\leq T}\{W(t)+Y(t)\},$
the joint density of $(\widehat{M}(T),\widehat{W}(T))$ can be found as
$f(m,w)=\frac{2(2m-w)}{T\sqrt{2\pi T}}e^{\frac{\delta w}{\sigma}-\frac{\delta^{2}T}{2\sigma^{2}}-\frac{(2m-w)^{2}}{2T}},w\leq m,m\geq 0,$
and is zero otherwise (see Shreve et al., (2004) for a detailed derivation). Consequently, the probability inside the integral becomes
$\Pr\left(\widehat{M}(T)<\frac{v}{\sigma},\quad\widehat{W}(T)<\frac{v-\omega}{\sigma}\right)=\int_{0}^{v}\int_{0}^{v-\omega}f(m,w)\mathop{}\!\mathrm{d}m\mathop{}\!\mathrm{d}w.$
The recovery probability can then be determined by numerically solving the triple integral.
4.5 Cause-of-death modeling
In our vitality-based modeling framework (equation (1)), death can be triggered either by the natural depletion of vitality (represented by components C2 and C3) or by a sudden jump due to a fatal accident (represented by component C4). Therefore, it is of practical interest to distinguish the cause of death based on the components of $V(t)$. In this subsection, we demonstrate how the vitality model can provide probabilities of different causes of death.
For illustrative purposes, we focus on the vitality model with components C1, C2, and C4 only. Specifically, we assume that component C2 follows the Gompertz law, while component C4 is modeled as a compound Poisson process with jump intensity $\lambda>0$ and i.i.d. jump sizes $\left\{Z_{i}\right\}_{i\in\mathbb{N}_{+}}$. The jump sizes have a common cumulative distribution function denoted as ${F}_{Z}(z)=1-\mathrm{e}^{-\alpha z}$ and survival function $\bar{F}_{Z}(z)=1-F_{Z}(z)$. The resulting vitality model is expressed as follows:
$V(t)=V(0)-\frac{bc^{x}}{\ln c}\left(c^{t}-1\right)-\sum_{i=1}^{N(t)}Z_{i},\quad t\geq 0.$
Let us denote $\tau_{J}=\inf\{t\geq 0:V(t)<0\}$ and $\tau_{Y}=\inf\{t\geq 0:V(t)=0\}$. In other words, $\tau_{J}\cdot\mathbb{1}(\tau_{J}<\tau_{Y})$ represents the death time at which death is caused by an accident, while $\tau_{Y}\cdot\mathbb{1}(\tau_{Y}<\tau_{J})$ represents the death time at which death is caused by natural decay. It is clear that $\tau=\tau_{Y}\wedge\tau_{J}$. The next theorem gives the expressions of the Laplace transforms and densities of $\tau_{J}$ and $\tau_{Y}$, given that the initial vitality level is $v$. For ease of presentation, we let $\mathrm{Pr}_{v}:=\mathrm{Pr}(\cdot|V(0)=v)$ and $\mathbb{E}_{v}[\cdot]:=\mathbb{E}[\cdot|V(0)=v]$.
Theorem 4.1.
For $q\geq 0$, the Laplace transforms of the time of death due to an accident and the time of death due to natural decay in vitality level are given by the following expressions:
$\mathbb{E}_{v}\left[\mathrm{e}^{-q\tau_{J}}\mathbf{1}\left(\tau_{J}<\tau_{Y}\right)\right]=$ $\int_{0}^{t_{v}^{*}}\lambda\mathrm{e}^{-(q+\lambda)t-\alpha b^{*}_{v}}\cdot I_{0}\left(2\sqrt{\alpha\lambda tb^{*}_{v}}\right)\mathrm{d}t,$ (13)
and
$\mathbb{E}_{v}\left[\mathrm{e}^{-q\tau_{Y}}\mathbf{1}\left(\tau_{Y}<\tau_{J}\right)\right]=$ $\int_{0}^{t^{*}_{v}}\mathrm{e}^{-\left(q+\lambda\right)t-\alpha b^{*}_{v}}\cdot bc^{x+t}\sqrt{\frac{\alpha\lambda t}{b^{*}_{v}}}I_{1}\left(2\sqrt{\alpha\lambda tb^{*}_{v}}\right)\mathrm{d}t+\mathrm{e}^{-(q+\lambda)t^{*}_{v}},$ (14)
where $t^{*}_{v}=\frac{\ln(bc^{x}+v\ln c)-\ln b}{\ln c}-x$ (which refers to the remaining lifetime without any accidents/jumps), $b^{*}_{v}=v-bc^{x}\left(c^{t}-1\right)/\ln c$, and
$I_{k}(x):=\sum_{n=0}^{\infty}\frac{(x/2)^{k+2n}}{n!\Gamma(k+n+1)}$
is the modified Bessel function of the first kind. Moreover, by the uniqueness of the Laplace transform, it follows that
$\mathrm{Pr}_{v}\left(\tau_{J}\in\mathrm{d}t,\tau_{J}<\tau_{Y}\right)$ $=\mathbb{1}(t<t^{*})\lambda\mathrm{e}^{-\alpha b^{*}_{v}-\lambda t}I_{0}\left(2\sqrt{\lambda\alpha tb^{*}_{v}}\right)\mathrm{d}t,$
and
$\mathrm{Pr}_{v}\left(\tau_{Y}\in\mathrm{d}t,\tau_{Y}<\tau_{J}\right)=$ $\mathbb{1}(t<t^{*})\mathrm{e}^{-\lambda t-\alpha b^{*}_{v}}\cdot bc^{x+t}\sqrt{\frac{\alpha\lambda t}{b^{*}_{v}}}I_{1}\left(2\sqrt{\alpha\lambda tb^{*}_{v}}\right)\mathrm{d}t+\mathrm{e}^{-\lambda t}\delta_{t^{*}}(\mathrm{d}t),$
where $\delta_{t_{v}^{*}}(\cdot)$ is the Dirac mass at $t_{v}^{*}$.
We note that the proof of Theorem 4.1 can be modified to accommodate more general assumptions on the vitality dynamics $V(t)$. For instance, the trend component $Y(t)$ could be any general non-decreasing continuous function, and the jump component $J(t)$ could be a renewal process with positive jumps. Moreover, instead of a single jump process for the jump component, multiple jump processes can be incorporated to represent various causes of death, such as car accidents, murder, suicide, cancer, etc. In this way, a more detailed cause-of-death analysis can be performed.
5 A Numerical Illustration
In this section, we provide an illustrative example to discuss the numerical estimation procedure of the proposed vitality-based mortality modeling framework. We consider the jump-diffusion specification shown in Section 3.1.3. Recall from equations (4) and (5) that the dynamics of vitality is governed by
$V(t)=V(0)-\int_{0}^{t}\mu_{x}(s)\mathop{}\!\mathrm{d}s-\sigma B(t)-\sum_{i=1}^{N(t)}Z_{i},$
and the resulting $T$-year survival probability is
$\Pr(\tau>T)=\int_{0}^{\infty}\mathbb{P}\left(v-\sigma B(t)-\sum_{i=1}^{N(t)}Z_{i}\geq\int_{0}^{t}\mu_{x}(s)\mathop{}\!\mathrm{d}s,\;\forall t\leq T\right)\mathop{}\!\mathrm{d}F_{0}(v).$
5.1 Estimation simplifications
The computational requirements for evaluating the conditional probability in the equation of $\Pr(\tau>T)$ are quite intensive. For ease of computation, we make the following assumptions to achieve an approximation:
  • The force of mortality is constant between integer ages.
  • All of the jumps are fatal accidents (i.e., $Z_{i}=\infty$ for all $i$).
Under the constant force of mortality assumption, $Y(t)$ can be written as a piece-wise linear function:
$Y(t)=a_{i}(t-\lfloor t\rfloor)+b_{i},$
for $i\leq t<i+1$ and $i\in\{0,1,2,\ldots\}$, where
$b_{i}$ $=\int_{0}^{i}\mu_{x}(s)\mathop{}\!\mathrm{d}s\quad\text{and}\quad a_{i}=\int_{i}^{i+1}\mu_{x}(s)\mathop{}\!\mathrm{d}s.$
Using the well-known formula for the non-crossing probability for a linear boundary $at+b$ before time $s$, that is
$\Pr(B(t)<at+b\text{ for all }t\leq s\big{|}B(s)=x)=1-\exp\left\{-\frac{2b(as+b-x)}{s}\right\},$
together with the fatal jump assumption, the conditional survival probability is given by
${}_{k}p_{x}(v)=\mathbb{E}\left[\left(\prod_{s=1}^{k}\mathbb{1}_{B(s)<\frac{v-Y(s)}{\sigma}}\cdot\left(1-e^{-\frac{2(v-Y(s-1)-\sigma B(s-1))(v-Y(s)-\sigma B(s))}{\sigma^{2}}}\right)\right)\cdot\exp\left(-\int_{0}^{k}\lambda(s)\mathop{}\!\mathrm{d}s\right)\right],k\in\mathbb{N}^{+},$
for a given initial vitality $v>0$. It follows that the $T$-year survival probability becomes
$\Pr(\tau>T)=\int_{0}^{\infty}{}_{T}p_{x}(v)\mathop{}\!\mathrm{d}F_{0}(v).$
5.2 Maximum likelihood estimation
In this subsection, we consider the maximum likelihood estimation method for estimating the parameters of the vitality model. We begin by describing the likelihood function for a cohort mortality dataset.
Consider a cohort of individuals all aged $x$ at time $t=0$. Let $E_{x}$ be the initial exposure at time 00, and let $D_{x+t}$ be the observed death counts at age $x+t$ for $t\in\{0,1,2,\ldots\}$. We assume that $D_{x+t}$ follows a multinomial distribution, and thus the log-likelihood function of the cohort mortality data is
$l(\theta)=\sum_{t\geq 0}D_{x+t}\cdot\ln\left[\Pr(\tau>t)-\Pr(\tau>t+1)\right]+\ln E_{x}!-\sum_{t\geq 0}\ln D_{x+t}!,$
where $\theta$ is the set of all model parameters. Note that our estimation procedure differs from Shimizu et al., (2021) and Shimizu et al., (2023), who used the least squares method to estimate the parameters of their survival energy models.
Given the complexity of our vitality model, relying solely on a cohort mortality dataset from a national population is inadequate for estimation purposes. Specifically, a national cohort dataset does not distinguish between deaths caused by natural aging and those resulting from accidents. This limitation makes it challenging to differentiate between accidental deaths, captured by the jump component, and natural deaths, represented by the initial and depletion components. In terms of parameter estimation, the intensity rate $\lambda(t)$ of the Poisson process $N(t)$ cannot be easily estimated without information that explicitly categorizes accidental deaths.
To address this issue, we utilize mortality data specifically related to external causes of death. In particular, we use external death counts from the Centers for Disease Control and Prevention (CDC) of the United States to estimate the intensity rate $\lambda(t)$. The remaining parameters, $\theta_{-\lambda}=\theta\backslash\{\lambda(t)\}_{t=0,1,\cdots}$, are then estimated using the maximum likelihood estimation method. This approach allows us to accurately separate the jump component from the rest of the components. The numerical results of our estimation procedures are presented in the next subsection.
5.3 Numerical estimation results
To estimate the vitality model, we utilize the U.S. male mortality data from the Human Mortality Database (HMD) [^1] Human Mortality Database. Max Planck Institute for Demographic Research (Germany), University of California, Berkeley (USA), and French Institute for Demographic Studies (France). Available at www.mortality.org.. We focus on the cohort born in 1910, with ages 60 and above during the period from 1970 to 2016. Additionally, we employ data from the Centers for Disease Control and Prevention (CDC) that consists of the average accidental death rates for age groups in ten-year intervals, spanning from 1968 to 2016. The jump intensity $\lambda(t)$ is directly determined as a constant within each age group based on the CDC data. The remaining parameters are fitted to the HMD data described above.
Recall that the jump-diffusion vitality model from Section 3.1.3 uses the Gompertz law to specify the trend component; that is, $\mu_{x}(t)=bc^{x+t}$. It follows that the vitality dynamics are governed by
$V(t)=V(0)-\int_{0}^{t}bc^{x+s}\mathop{}\!\mathrm{d}s-\sigma B(t)-\sum_{i=1}^{N(t)}Z_{i},$
where $V(0)\sim\text{Exp}(1)$. Given that the intensity rate of $N(t)$ has been externally estimated and we assumed $Z_{i}=\infty$ for all $i$, the remaining parameters to be estimated are $b$, $c$, and $\sigma$. In addition to the vitality model, we also estimate the parameters of the Gompertz law for comparison purposes.
Table 1 reports the estimated values of $b$, $c$, and $\sigma$ for the vitality model and the estimated values of $b$ and $c$ for the Gompertz law. The maximized log-likelihood of each model is also reported in the last row. We observe that the estimated values of $b$ and $c$ are similar between the vitality model and the Gompertz law, indicating that both models accurately describe the underlying mortality curve via the assumed functional form, $\mu_{x}(t)=bc^{x+t}$. In addition to the Gompertz law, the vitality model has an estimated value of 0.0036 for $\sigma$, which reflects the random fluctuations in vitality over the lifetime of the cohort. Overall, the vitality model achieves a higher maximized log-likelihood of -2,181 compared to the Gompertz law’s -2,591.
Model Vitality Gompertz
$b$ 0.00015391 0.0001744
$c$ 1.0834 1.082
$\sigma$ 0.0036
$l(\theta)$2,181 -2,591
Table 1: Parameter estimates and maximized log-likelihood for the vitality model and the Gompertz law.
Figure 2 shows the fitted values of death rates and survival probabilities for ages 60-100 and periods 0-40 years under the vitality model and the Gompertz law. The observed values from the empirical data are also plotted for comparison. We observe that both the vitality model and the Gompertz law are able to depict a well-fitted mortality curve for the cohort considered and also produce reasonable survival probabilities at age 60 for the next 40 years. Comparing the two models, the vitality model appears to better fit the cohort’s mortality data and produce a mortality curve that is closer to the observed values.
Figure 2: Fitted log death rates $\ln q_{x}$ for $x\in(60,100)$ and survival probabilities ${}_{t}p_{60}$ for $t\in(0,40)$ under the vitality model (blue solid line) and the Gompertz law (red dash line). The observed values are depicted by dots.
6 Conclusion
In this paper, we introduced a vitality-based modeling framework that offers a new perspective on mortality modeling. The proposed framework integrates multiple components to capture the complexities of vitality dynamics over an individual’s lifetime. By incorporating the initial vitality level, natural depletion trends, stochastic diffusion processes, and sudden jumps, our framework provides a comprehensive and flexible approach to mortality modeling. The adaptability of the four components allows for the incorporation of various factors affecting mortality, providing deeper insights into population mortality trends and individual health trajectories.
Through both static and dynamic specifications of the proposed framework, we demonstrated its versatility in addressing various mortality modeling scenarios. Our framework not only replicates traditional mortality laws but also extends to more advanced stochastic mortality models, offering a unified structure for understanding mortality dynamics. An illustrative example is provided to showcase how a vitality model can be fitted using the maximum likelihood estimation method. Additionally, the framework’s modularity ensures that it can be tailored to specific research needs, making it a valuable tool for both theoretical exploration and practical applications.
Our findings suggest that the vitality-based modeling framework offers significant practical applications, such as life insurance valuation, disability modeling, cause-of-death modeling, explaining subjective beliefs about mortality, and deriving optimal consumption solutions. By capturing the initial level, natural aging, inherent randomness, and sudden changes in an individual’s vitality, the framework supports a better interpretation of mortality dynamics in the aforementioned applications. Lastly, we emphasize that the proposed framework can be further utilized in more nuanced scenarios to improve results and decision-making in various fields.
While we used four components to develop a comprehensive modeling framework for vitality dynamics, the specific structures considered for each component are not exhaustive in this paper. There are numerous possible extensions for articulating the four components. For instance, one could consider a multivariate stochastic process to govern the diffusion component, reflecting the fact that multiple underlying forces may drive the randomness in an individual’s vitality. The initial component could be modified to measure and explain healthy life expectancy. The jump component can also be expanded to consider jumps categorized by different causes of death, such as natural disasters, car accidents, heart attacks, and suicide.
We have demonstrated in this paper that the proposed modeling framework is capable of replicating mortality laws and stochastic mortality models. Beyond these existing models, further exploration can be conducted to show that continuous-time intensity-based models, such as the Vasicek process applied to the force of mortality, can also be mimicked by a vitality model. It would be interesting to understand how the first passage time problem connects with an continuous-time mortality model to produce explicit solutions for survival probabilities. This connection could potentially provide new insights and more accurate survival probability estimates.
Although we have explored various applications of the vitality-based modeling framework, there are several other potential applications that we did not focus on. For example, a dynamic vitality model can be used to manage longevity risk by developing dynamic hedging strategies and pricing longevity-linked securities. Additionally, since our framework models individual vitality, it can be employed to measure the basis risk between an individual insurance product and an index-based security. These extensions would provide further advancements to capital market solutions for longevity risk management.
Lastly, we acknowledge that the vitality dynamics of an individual cannot be directly observed, or at least not easily. To accurately estimate a vitality model using data beyond the national population data presented in this paper, one would need to use personal data over time or clinical data on vital measures. Potential sources of such data could include personal medical devices, such as smartwatches or mobile health trackers, which continuously monitor health metrics like heart rate, physical activity, and sleep patterns. Integrating these diverse data sources would allow for a more precise estimation of vitality dynamics. We remark that this research direction is highly related to InsurTech and could inspire future studies on personalized insurance products.
Acknowledgement
We extend our gratitude for the helpful and constructive comments received at the School of Actuarial Science at Université Laval, the 27th International Congress on Insurance: Mathematics and Economics, and the 59th Actuarial Research Conference. Xiaobai Zhu acknowledges the support from the Research Grants Council of the Hong Kong Special Administrative Region, China (CUHK 24615523), and National Natural Science Foundation of China (No. 12301613). Zijia Wang acknowledges the support from a start-up grant from The Chinese University of Hong Kong.
Appendix A Proofs
A.1 Death Plateau Models
Gamma-Gompertz Model
Consider a hazard rate model with the hazard rate function being $h(t)=Z\cdot\mu_{x}(t)$, where $Z$ follows a Gamma distribution with shape parameter $\alpha$ and rate parameter $\phi$. Note that the conditional $T$-year survival probability given $Z=z$ can be expressed as $\exp\left(-\int_{0}^{T}z\cdot\mu_{x}(t)\mathop{}\!\mathrm{d}t\right)$, then the unconditional $T$-year survival probability is
$\Pr(\tau>T)=$ $\int_{0}^{\infty}\left(e^{-\int_{0}^{T}z\cdot\mu_{x}(t)\mathop{}\!\mathrm{d}t}\right)\cdot\frac{\phi^{\alpha}\cdot z^{\alpha-1}\cdot e^{-z\cdot\phi}}{\Gamma(\alpha)}\mathop{}\!\mathrm{d}z$
[…]
$=$ $\left(1+\frac{\int_{0}^{T}\mu_{x}(t)\mathop{}\!\mathrm{d}t}{\phi}\right)^{-\alpha}.$
In particular, if $\mu_{x}(t)=bc^{x+t}$, it is known as the Gamma-Gompertz law.
Exponential Initial Vitality and Gamma Decay Rate
If we assume that the initial vitality level $V(0)\sim\text{Exp}(1)$ and $Y(t)$ follows
$\mathop{}\!\mathrm{d}Y(t)=Z\cdot\mu_{x}(t)\mathop{}\!\mathrm{d}t,\quad t\geq 0,$
where $Z$ follows a Gamma distribution with shape $\alpha$ and rate $\phi$ and independent of $V(0)$, then we have
$\Pr(\tau>T)$ $=\Pr\left(V(0)-Z\cdot\int_{0}^{T}\mu_{x}(s)\mathop{}\!\mathrm{d}s>0\right)$
$=\int_{0}^{\infty}\Pr\left(V(0)>z\cdot\int_{0}^{T}\mu_{x}(s)\mathop{}\!\mathrm{d}s\right)\cdot\frac{\phi^{\alpha}\cdot z^{\alpha-1}\cdot e^{-z\cdot\phi}}{\Gamma(\alpha)}\mathop{}\!\mathrm{d}z$
$=\int_{0}^{\infty}\exp\left(-z\cdot\int_{0}^{T}\mu_{x}(s)\mathop{}\!\mathrm{d}s\right)\cdot\frac{\phi^{\alpha}\cdot z^{\alpha-1}\cdot e^{-z\cdot\phi}}{\Gamma(\alpha)}\mathop{}\!\mathrm{d}z.$
The remaining steps follow exactly the same as before, and we will end up with the same survival function as in Appendix A.1.
Dagum Decay Rate
Assume that the initial vitality is a constant $v$, and
$\mathop{}\!\mathrm{d}Y(t)=Z\cdot\mu_{x}(t)\mathop{}\!\mathrm{d}t,\quad t\geq 0,$
where $Z$ follows a Dagum (Inverse-Burr) distribution with parameters $p$, $a$, $b$ and distribution function $F_{Z}(\cdot)$, then we have
$\Pr(\tau>T)$ $=\int_{0}^{\infty}\mathbb{1}\left(v-z\cdot\int_{0}^{T}\mu_{x}(s)\mathop{}\!\mathrm{d}s>0\right)\mathop{}\!\mathrm{d}F_{Z}(z)$
$=\int_{0}^{\infty}\mathbb{1}\left(z<\frac{v}{\int_{0}^{T}\mu_{x}(s)\mathop{}\!\mathrm{d}s}\right)\mathop{}\!\mathrm{d}F_{Z}(z)$
$=\left(1+\left(\frac{v}{b\cdot\left(\int_{0}^{T}\mu_{x}(s)\mathop{}\!\mathrm{d}s\right)}\right)^{-a}\right)^{-p}.$
If $a=1$, $p=\alpha$, and $b=\frac{v}{\phi}$, then we recover the same survival function as in Appendix A.1.
A.2 Alternative Expression of Mortality Laws
Noting that the vitality model given in equation (7) is a special case of the class of SNLPs, one can adopt the fluctuation theory of SNLP (see Kyprianou, (2014) for more details) to study $\tau$. For example, suppose the distribution of $\{Z_{i}\}_{i\geq 1}$ is a mixture of exponential distributions with probability density function
$f_{Z}(z)=\sum_{i=1}^{n}p_{i}\alpha_{i}\mathrm{e}^{-\alpha_{i}z},\quad z>0,$
where $\alpha_{i}>0$ and $\sum_{i=1}^{n}p_{i}=1$. In this case, it is known from Guérin and Renaud, (2016) that the Laplace transform of $\tau$ given that $V(0)=v$ is
$\mathbb{E}_{v}\left[\mathrm{e}^{-q\tau}\right]$ $=q\sum_{i=1}^{n+2}\frac{\mathrm{e}^{\theta_{i}^{(q)}v}}{\theta_{i}^{(q)}g\left(\theta_{i}^{(q)}\right)}-\frac{q}{\theta_{1}^{(q)}}\sum_{i=1}^{n+2}\frac{\mathrm{e}^{\theta_{i}^{(q)}v}}{g\left(\theta_{i}^{(q)}\right)},\quad q>0,$
where $\theta_{n+2}^{(q)}<\theta_{n+1}^{(q)}<\theta_{n}^{(q)}\ldots<\theta_{2}^{(q)}\leq 0\leq\theta_{1}^{(q)}$ are the roots (in $y$) to the equation
$\frac{1}{2}\sigma^{2}y^{2}-\delta y+\sum_{i=1}^{n}\frac{\lambda p_{i}\alpha_{i}}{y+\alpha_{i}}-\lambda-q=0,$
and
$g(y)=\sigma^{2}y-\delta-\sum_{i=1}^{n}\frac{\lambda p_{i}\alpha_{i}}{\left(y+\alpha_{i}\right)^{2}}.$
Then, by either analytically or numerically invert the above Laplace transform with respect to $q$, one can obtain the distribution function of $\tau$.
A.3 Dynamic Model
When $Y(t)$ is non-decreasing, for example, the integral of correlated log-normally distributed random variables as in Section 3.2.2, then
$\Pr(\tau>T)$ $=\Pr(V_{y}(0)>Y(T;y))$
$=\int_{0}^{\infty}\Pr(V_{y}(0)>Y(T;y)|\gamma)\cdot f_{y}(\gamma)\mathop{}\!\mathrm{d}\gamma$
$=\int_{0}^{\infty}\mathrm{Pr}\left(\mathrm{e}_{\gamma}>Y(T;y)\right)\cdot f_{y}(\gamma)\mathop{}\!\mathrm{d}\gamma$
$=\int_{0}^{\infty}\mathbb{E}\left[\mathrm{e}^{-\gamma Y(T;y)}\right]\cdot f_{y}(\gamma)\mathop{}\!\mathrm{d}\gamma$
$=\mathbb{E}\left[\exp\left(-\Gamma(y)\cdot Y(T;y)\right)\right],$
where $\mathrm{e}_{\gamma}$ is an exponential random variable with parameter $\gamma$, and $f_{y}(\gamma)$ is the density function for $\Gamma(y)$ of cohort $y$.
Appendix B Monte Carlo Valuation
Here we employ the Monte Carlo valuation to the survival probability of the model in Section 3.1.3. Denote $H(t,v)=v-Y(t)$ for ease of notation, we can apply the following procedure to approximate ${}_{t}p_{x}(v)$:
  1. 1.
    Select $n+1$ time points $\{t_{i}\}_{i=0,\cdots,n}$ such that $0=t_{0}<t_{1}<...<t_{n}=t$.
  2. 2.
    Given that $N(t)=k>0$, generate the first jump tim $v_{1}$ with distribution function at time $r_{1}<t$ being $\frac{\int_{0}^{r_{1}}\lambda(s)\mathop{}\!\mathrm{d}s}{\int_{0}^{t}\lambda(s)\mathop{}\!\mathrm{d}s}$, then we can iteratively draw the inter-arrival time between the $(i-1)$-th and the $i$-th jumps with distribution function for $r_{i}$, $r_{i-1}<r_{i}<t$, being $\frac{\int_{r_{i-1}}^{r_{i}}\lambda(s)\mathop{}\!\mathrm{d}s}{\int_{r_{i-1}}^{t}\lambda(s)\mathop{}\!\mathrm{d}s}$. Let $s_{1}=r_{1}$, $s_{2}=r_{1}+r_{2}$, $\cdots$, $s_{k}=\sum_{i=1}^{k}r_{i}$, which are the arrival time for each jump.
  3. 3.
    By combining the two sets of partitions $\{\tilde{u}_{i}\}_{i=0,\cdots,n+k}=\{t_{i}\}_{i=0,\cdots,n}\cup\{s_{i}\}_{i=1,\cdots,k}$ and re-ordering this set to have a new set denoted as $\{u_{i}\}_{i=0,\cdots,n+k}$, which is in ascending order such that $0=u_{0}<u_{1}<...<u_{n-1+k}<u_{n+k}=t$.
  4. 4.
    Generate $\{Z_{i}\}_{i=1,\cdots,k}$ jump sizes from $k$ i.i.d. random variables with the cdf $F_{Z}(\cdot)$, denote the simulated samples as $z_{1},z_{2},\ldots z_{k}$.
  5. 5.
    Compute the following quantities:
    $b_{0}^{+}:=\frac{H(0,v)}{\sigma},$
    $b_{1}^{+}:=\frac{H(u_{1},v)}{\sigma}-\sum_{j=1}^{k}\frac{z_{j}}{\sigma}\mathbb{1}_{\{s_{j}\leq u_{1}\}},\ \ldots,\ b_{i}^{+}:=\frac{H(u_{i},v)}{\sigma}-\sum_{j=1}^{k}\frac{z_{j}}{\sigma}\mathbb{1}_{\{s_{j}\leq u_{i}\}},$
    $b_{1}^{-}:=\frac{H(u_{1},v)}{\sigma}-\sum_{j=1}^{k}\frac{z_{j}}{\sigma}\mathbb{1}_{\{s_{j}<u_{1}\}},\ \ldots,\ b_{i}^{-}:=\frac{H(u_{i},v)}{\sigma}-\sum_{j=1}^{k}\frac{z_{j}}{\sigma}\mathbb{1}_{\{s_{j}<u_{i}\}},$
    for $i=1,2,...,n+k$.
  6. 6.
    Generate $(y_{1},y_{2},...y_{n+k-1})$ from the $(n+k-1)$-dimensional multivariate normal distribution of $B(u_{1}),B(u_{2}),...,B(u_{n+k-1})$ and let $y_{0}=0$.
  7. 7.
    Compute
    $Q_{1}=\prod_{i=1}^{n+k}\mathbb{1}\left(y_{i}<b_{i}^{+}\right)\left(1-\exp\left\{-\frac{2\left(b_{i-1}^{+}-y_{i-1}\right)\left(b_{i}^{-}-y_{i}\right)}{u_{i}-u_{i-1}}\right\}\right),$
    where $\Phi$ is the cumulative distribution function of a standard normal distribution.
  8. 8.
    Repeat Steps 1 – 7 for $m$ times, to obtain $Q_{i}$, $i=1,\cdots,m$ (from Step 7), and the survival probability (conditional on initial vitality of $v$) can be approximated by
    ${}_{t}p_{x}(v)\approx\frac{1}{m}\sum_{i=1}^{m}Q_{i}.$
Appendix C Optimal Investment and Consumption
C.1 Merton’s Problem - Infinite Time
The value function of Merton’s problem under an infinite time horizon at time $t$ is
$J(a)=\max_{\pi,\zeta\in\Pi}\mathbb{E}_{t,a}\left[\int_{t}^{\infty}e^{-\beta(s-t)}\cdot u(\zeta(s))\mathop{}\!\mathrm{d}s\right].$
The associated Hamilton–Jacobi–Bellman (HJB) equation can be expressed as
$\max_{\pi,\zeta}\left\{J_{a}\cdot\left[a\cdot r+a\cdot\pi\cdot\theta\cdot\sigma_{S}-\zeta\right]+J_{aa}\cdot\frac{a^{2}\cdot\sigma_{S}^{2}\cdot\pi^{2}}{2}+u(\zeta)\right\}-\beta\cdot J=0.$
The value function can be solved explicitly as
$J(a)=F\cdot\ln a+G,\quad\text{where }F=\frac{1}{\beta},\text{ }G=\exp\left(\frac{[r+\theta^{2}/2-\beta]/\beta+\ln\beta}{\beta}\right),$
and the optimal strategies are
$\pi^{*}(t)=\frac{\theta}{\sigma_{S}},\quad\zeta^{*}(t)=\beta\cdot A(t).$
C.2 Merton’s Problem - Mortality
Consider again Merton’s problem with finite planning horizon $T$ and with mortality modeled by a deterministic force of mortality function $\mu_{x}$. The value function at time $t$ (for a person age $x$ at time 00) is
$J(t,a)=\max_{\pi,\zeta\in\Pi}\mathbb{E}_{t,a}\left[\int_{t}^{T}e^{-\beta(s-t)}\cdot e^{-\int_{t}^{s}\mu_{x+y}\mathop{}\!\mathrm{d}y}\cdot\left(u(\zeta(s))+\mu_{x+s}\cdot\lambda_{1}\cdot u(A(s))\right)\mathop{}\!\mathrm{d}s+\lambda_{2}\cdot u(A(T))\right],$
where $\lambda_{1}$ and $\lambda_{2}$ are the weighting parameters for the bequest and the end-of-period wealth, respectively. The associated HJB equation is
$J_{t}+\max_{\pi,\zeta}\left\{J_{a}\cdot\left[a\cdot r+a\cdot\pi\cdot\theta\cdot\sigma_{S}-\zeta\right]+J_{aa}\cdot\frac{a^{2}\cdot\sigma_{S}^{2}\cdot\pi^{2}}{2}+u(\zeta)+\mu_{x+t}\cdot\lambda_{1}\cdot u(a)\right\}-(\beta+\mu_{x+t})\cdot J=0$
By conjecture that $J(t,a)=F(t)\cdot\ln a+G(t)$ with terminal condition
$J(T,a)=\lambda_{2}\cdot\ln a,\quad\implies F(T)=\lambda_{2}\quad\text{and}\quad G(T)=0,$
and by the first order condition, the optimal strategies are
$\pi^{*}$ $=\frac{-J_{a}\cdot\theta}{J_{aa}\cdot a\cdot\sigma_{S}}=\frac{\theta}{\sigma_{S}},\quad\zeta^{*}=\frac{1}{J_{a}}=\frac{a}{F(t)}.$
Substitute into the HJB equation and simplify:
$F^{\prime}(t)\ln a+G^{\prime}(t)+\frac{F(t)}{a}\cdot\left(a\cdot r+a\cdot\theta^{2}-\frac{a}{F(t)}\right)-\frac{F(t)}{a^{2}}\cdot\frac{a^{2}\theta^{2}}{2}$
$\qquad+\ln a-\ln F(t)+\mu_{x+t}\cdot\lambda_{1}\cdot u(a)-(\beta+\mu_{x+t})(F(t)\ln a+G(t))=0.$
Grouping terms with $a$ and without $a$, we have
$F^{\prime}(t)+1+\mu_{x+t}\cdot\lambda_{1}-(\beta+\mu_{x+t})\cdot F(t)=0$
$\implies$ $F(t)=\lambda_{2}\cdot e^{-\int_{t}^{T}\beta+\mu_{x+y}\mathop{}\!\mathrm{d}y}+\int_{t}^{T}e^{-\int_{t}^{s}\beta+\mu_{x+y}\mathop{}\!\mathrm{d}y}\cdot(1+\mu_{x+s}\cdot\lambda_{1})\mathop{}\!\mathrm{d}s$
$G^{\prime}(t)+F(t)\cdot r+F(t)\cdot\frac{\theta^{2}}{2}-1-\ln F(t)-(\beta+\mu_{x+t})\cdot G(t)=0$
$\implies$ $G(t)=\int_{t}^{T}e^{-\int_{t}^{s}\beta+\mu_{x+y}\mathop{}\!\mathrm{d}y}\cdot\left(F(s)\cdot\left(r+\frac{\theta^{2}}{2}\right)-1-\ln F(s)\right)\mathop{}\!\mathrm{d}s.$
In particular, when $T\rightarrow\infty$, $F(t)$ becomes
$F(t)$ $=\int_{t}^{\infty}e^{-\beta(t-s)}\cdot e^{-\int_{t}^{s}\mu_{x+y}\mathop{}\!\mathrm{d}y}\mathop{}\!\mathrm{d}s+\int_{t}^{\infty}e^{-\beta(t-s)}\cdot e^{-\int_{t}^{s}\mu_{x+y}\mathop{}\!\mathrm{d}y}\cdot\mu_{x+s}\cdot\lambda_{1}\mathop{}\!\mathrm{d}s$
$=\bar{a}_{x+t}+\lambda_{1}\cdot\bar{A}_{x+t},$
where $\bar{a}_{x}$ and $\bar{A}_{x}$ are annuity price and insurance price issued to a person aged $x$ with \$1 paid continuously, evaluated using a discount rate of $\beta$.
C.3 Merton’s Problem with Vitality
The associated HJB equation for the value function $J(a,v)$ is
$\max_{\pi,\zeta}\left\{J_{a}\cdot\left[a\cdot r+a\cdot\pi\cdot\theta\cdot\sigma_{S}-\zeta\right]+J_{aa}\cdot\frac{a^{2}\cdot\sigma_{S}^{2}\cdot\pi^{2}}{2}-J_{v}\cdot\delta+J_{vv}\cdot\frac{\sigma_{V}^{2}}{2}+u(\zeta)\right\}-\beta\cdot J=0$
with boundary conditions
$\lim_{v\rightarrow\infty}J(a,v)$ $=F\ln a+G$
$\lim_{v\rightarrow 0}J(a,v)$ $=\lambda\ln a,$
where $F$ and $G$ are defined in the value function in Section C.1. The first line corresponds to the case where the individual will never die because the vitality approaches infinite, and thus the problem collapses to Merton’s problem in Section C.1. The second line corresponds to when the individual is already dead since the vitality has been exhausted, and thus the individual’s utility is reflected only through the bequest.
By the first-order condition, the optimal strategies can be expressed as:
$\pi^{*}$ $=\frac{-J_{a}\cdot a\cdot\theta\cdot\sigma_{S}}{J_{aa}\cdot a^{2}\cdot\sigma_{S}^{2}}\quad\text{and}\quad\zeta^{*}=\frac{1}{J_{a}}.$
By conjecturing on the form of the value function $J(v,a)=f(v)\cdot\ln a+g(v)$, where $f(v)$ and $g(v)$ are functions of the vitality level, then
$\pi^{*}$ $=\frac{\theta}{\sigma_{S}}\quad\text{and}\quad\zeta^{*}=\frac{a}{f(v)}.$
Substituting the expression back to the HJB equation and collecting the terms, we have
$\delta\cdot f^{\prime}(v)-f^{\prime\prime}(v)\cdot\frac{\sigma_{V}^{2}}{2}-1+\beta\cdot f(v)=0$
$f(v)\cdot\left[r+\frac{\theta^{2}}{2}\right]-g^{\prime}(v)\cdot\delta+g^{\prime\prime}(v)\cdot\frac{\sigma_{V}^{2}}{2}-\ln f(v)-\beta\cdot g(v)=0$
Therefore, we have
$f(v)=\left(\lambda-\frac{1}{\beta}\right)\cdot e^{k_{1}v}+\frac{1}{\beta},\quad k_{1}=\left\{\begin{array}[]{ll}\frac{\delta-\sqrt{\delta^{2}+2\sigma^{2}_{V}\beta}}{\sigma^{2}_{V}}<0,&\sigma_{V}>0\\ \frac{-\beta}{\delta},&\sigma_{V}=0.\end{array}\right.$
In particular, when $\sigma_{V}=0$, the optimal consumption strategy becomes
$f(t)=\bar{a}_{\hbox{\vtop{\vbox{\hrule\kern 1.0pt\hbox{$\scriptstyle t^{*}$\kern 1.0pt}}\kern 1.0pt}\vrule\kern 1.0pt}}+\lambda\cdot e^{-\beta t^{*}},$
where $t^{*}=\frac{v}{\delta}$ is the remaining life, and $\bar{a}_{\hbox{\vtop{\vbox{\hrule\kern 1.0pt\hbox{$\scriptstyle t^{*}$\kern 1.0pt}}\kern 1.0pt}\vrule\kern 1.0pt}}$ is the annuity-certain evaluated with discount rate $\beta$. The expression of $f(t)$ has a similar interpretation with $F(t)$ in Section C.2 but based on the remaining life $t^{*}$ ($e^{-\beta t^{*}}$ is the present value of \$1 pays at time $t^{*}$).
Appendix D Approximation Methods
D.1 Tangent Approximation
If we set $J(t)=0$, and given the initial vitality level $V(0)=v$, then Jennen and Lerche, (1981); Skiadas and Skiadas, (2014) discussed a tangent approximation method to get the density of $\tau(v)$; that is,
$\Pr(\tau(v)\in\mathop{}\!\mathrm{d}t)=\frac{|H(t,v)-t\cdot H_{t}(t,v)|}{\sigma\sqrt{2\pi t^{3}}}e^{-\frac{\left(H(t,v)\right)^{2}}{2\sigma^{2}t}},$
where $H(t,v):=v-Y(t)$.
D.2 Series Expansion
The tangent approximation has the advantage of simplicity but may sacrifice the accuracy of the approximation. An extension that includes a series expansion has been proposed by Durbin and Williams, (1992). The density of $\tau(v)$ is approximated as
$\Pr(\tau(v)\in\mathop{}\!\mathrm{d}t)\approx\sum_{j=1}^{k}(-1)^{j-1}q_{j}(t),\quad k=1,2,\cdots,$
where
$q_{j}(t)=$ $\int_{0}^{t}\int_{0}^{t_{1}}\cdots\int_{0}^{t_{j-2}}\left[\frac{H\left(t_{j-1},v\right)}{\sigma t_{j-1}}-\sigma^{-1}H_{t}\left(t_{j-1},v\right)\right]$
×$\times\prod_{i=1}^{j-1}\left[\frac{H\left(t_{i-1},v\right)-H\left(t_{i},v\right)}{\sigma(t_{i-1}-t_{i})}-\sigma^{-1}H_{t}\left(t_{i-1},v\right)\right]f\left(t_{j-1},\cdots,t_{1},t\right)\mathop{}\!\mathrm{d}t_{j-1}\cdots\mathop{}\!\mathrm{d}t_{1},$
with $f\left(t_{j-1},\cdots,t_{1},t\right)$ being the joint density of $(B\left(t_{j-1}\right),\cdots,B\left(t_{1}\right),B(t))$ evaluated at the values of $(H(t_{j-1},v)/\sigma,\cdots,H(t_{1},v)/\sigma,H(t,v)/\sigma)$. One notable advantage of this approximation method is its rapid convergence when $H$ is concave in $t$. Durbin and Williams, (1992) demonstrated the accuracy of this series expansion using specific concave functions for $H$, and found that with a choice of $k=3$ (which leads to a double integration), the error lies within $0.001\%$ of the true value.
Appendix E Cause-of-Death Analysis
Proof of Theorem 4.1.
We begin by proving equation (13). Assume that the death occurs at the $n$th ($n\geq 1$) jump in the vitality level (at time $T_{n}$). Using the independence assumption between $N(t)$ and $\left\{Z_{i}\right\}_{i\geq 1}$, and conditional on the initial vitality $v$, we have
$\mathbb{E}_{v}\left[\mathrm{e}^{-q\tau_{J}}\mathbb{1}\left(\tau_{J}<\tau_{Y,}\tau_{J}=T_{n}\right)\right]=$ $\int_{-\infty}^{0}\int_{0}^{\infty}\mathbb{E}_{v}\left[\mathrm{e}^{-qT_{n}}\mathbb{1}\left(V\left(T_{n}^{-}\right)\in\mathrm{d}z,V\left(T_{n}\right)\in\mathrm{d}y\right)\right]$
$=$ $\int_{0}^{\infty}\bar{F}_{Z}(z)\mathbb{E}_{v}\left[\mathrm{e}^{-qT_{n}}\mathbb{1}\left(V\left(T_{n}^{-}\right)\in\mathrm{d}z\right)\right].$ (15)
Further conditioning on $T_{n}$, we have
$\mathbb{E}_{v}\left[\mathrm{e}^{-qT_{n}}\mathbb{1}\left(V\left(T_{n}^{-}\right)\in\mathrm{d}z\right)\right]$ $=\int_{0}^{\infty}\mathbb{1}(v-Y(t)>z)\mathbb{E}_{v}\left[\mathrm{e}^{-qt}\mathbb{1}\left(V\left(t\right)\in\mathrm{d}z\right)\right]\frac{\lambda^{n}t^{n-1}\mathrm{e}^{-\lambda t}}{(n-1)!}\mathrm{d}t$
$=\int_{0}^{\infty}\mathbb{1}(v-Y(t)>z)F^{*(n-1)}\left(v-Y(t)-\mathrm{d}z\right)\frac{\lambda^{n}t^{n-1}\mathrm{e}^{-\left(q+\lambda\right)t}}{(n-1)!}\mathrm{d}t.$
Substituting the above equation into equation (E) yields
$\mathbb{E}_{v}\left[\mathrm{e}^{-q\tau_{J}}\mathbb{1}\left(\tau_{J}<\tau_{Y,}\tau_{J}=T_{n}\right)\right]$
$=$ $\int_{0}^{\infty}\mathbb{1}\left(v-Y(t)>0\right)\left(\int_{0}^{v-Y(t)}f_{Z}^{*(n-1)}(v-Y(t)-z)\bar{F}_{Z}(z)\mathrm{d}z\right)\frac{\lambda^{n}t^{n-1}\mathrm{e}^{-(q+\lambda)t}}{(n-1)!}\mathrm{d}t$
$=$ $\int_{0}^{\infty}\mathbb{1}\left(v-Y(t)>0\right)\left(F_{Z}^{*(n-1)}\left(v-Y(t)\right)-F_{Z}^{*(n)}\left(v-Y(t)\right)\right)\frac{\lambda^{n}t^{n-1}\mathrm{e}^{-(q+\lambda)t}}{(n-1)!}\mathrm{d}t$
$=$ $\int_{0}^{\infty}\mathbb{1}(v-\frac{bc^{x}}{\ln c}\left(c^{t}-1\right)>0)\frac{\mathrm{e}^{-\alpha\left(v-\frac{bc^{x}}{\ln c}\left(c^{t}-1\right)\right)}\alpha^{n-1}\left(v-\frac{bc^{x}}{\ln c}\left(c^{t}-1\right)\right)^{n-1}}{(n-1)!}\frac{\lambda^{n}t^{n-1}\mathrm{e}^{-(q+\lambda)t}}{(n-1)!}\mathrm{d}t$
$=$ $\int_{0}^{t_{v}^{*}}\frac{\lambda\mathrm{e}^{-\alpha\left(v-\frac{bc^{x}}{\ln c}\left(c^{t}-1\right)\right)-(q+\lambda)t}\left(\lambda\alpha vt-\frac{\lambda\alpha bc^{x}t}{\ln c}\left(c^{t}-1\right)\right)^{n-1}}{(n-1)!(n-1)!}\mathrm{d}t,$ (16)
where the third equation follows from the fact that the $n$-fold convolution of $F_{Z}(z)$ with itself is given by
$F_{Z}^{*(n)}(z)$ $=1-\sum_{i=0}^{n-1}\frac{1}{i!}\mathrm{e}^{-\alpha z}(\alpha z)^{i}.$
Summing up equation (E) over $n$ leads to completes the proof of equation (13).
Now, suppose that the death occurs due to the decrement in $Y(\cdot)$ rather than a fatal accident, i.e., $\tau=\tau_{Y}<\tau_{J}$. Conditioning on $\{T_{n}=t,V(T_{n})=z>0\}$, the event that there are $n$ ($n\geq 0$) accidents occurred before death is equivalent to the event $\{\tau_{Y}=t+s,T_{n+1}-T_{n}>s\}$, where $s$ is the solution to $Y(t+s)-Y(t)=z$ (i.e., the vitality level $V$ decreases to 00 before the $(n+1)$th accident occurs). Solving the equation in $s$ leads to
$s=\frac{\ln\left(bc^{x+t}+z\ln c\right)-\ln\left(bc^{x+t}\right)}{\ln c}.$
Along with the fact that $T_{n+1}-T_{n}$ is exponentially distributed and independent of $(T_{n},V(T_{n}))$, we have
$\mathbb{E}_{v}\left[\mathrm{e}^{-q\tau_{Y}}\mathbb{1}\left(\tau_{Y}<\tau_{J,}N(\tau_{Y})=n\right)\right]$
$=$ $\int_{0}^{\infty}\mathbb{1}\left(v-Y(t)>0\right)\int_{0}^{v-Y(t)}\mathrm{e}^{-q(t+s)}\mathrm{Pr}_{v}\left(V\left(T_{n}\right)\in\mathrm{d}z,T_{n}\in\mathrm{d}t,T_{n+1}-T_{n}>s\right)$
$=$ $\int_{0}^{\infty}\mathbb{1}\left(v-Y(t)>0\right)\int_{0}^{v-Y(t)}\mathrm{e}^{-(q+\lambda)s}f_{Z}^{*(n)}(v-Y(t)-z)\frac{\lambda^{n}t^{n-1}\mathrm{e}^{-(q+\lambda)t}}{(n-1)!}\mathrm{d}z\mathrm{d}t.$
By substituting the representation of $s$ and the form of $Y$ into the above equation and performing some direct calculations, we obtain
$\mathbb{E}_{v}\left[\mathrm{e}^{-q\tau_{Y}}\mathbb{1}\left(\tau_{Y}<\tau_{J,}N(\tau_{Y})=n\right)\right]$
$=$ $\int_{0}^{t_{v}^{*}}\int_{0}^{v-\frac{bc^{x+t}-bc^{x}}{\ln c}}\left(\frac{\left(v-z\right)\ln c+bc^{x}}{bc^{x+t}}\right)^{-\frac{q+\lambda}{\ln c}}\frac{\alpha^{n}z^{n-1}\mathrm{e}^{-\alpha z}}{(n-1)!}\mathrm{d}z\frac{\lambda^{n}t^{n-1}\mathrm{e}^{-(q+\lambda)t}}{(n-1)!}\mathrm{d}t$
$=$ $\int_{0}^{t_{v}^{*}}\int_{0}^{v-\frac{bc^{x+t}-bc^{x}}{\ln c}}\left(\frac{\left(v-z\right)\ln c+bc^{x}}{bc^{x}}\right)^{-\frac{q+\lambda}{\ln c}}\frac{\alpha^{n}z^{n-1}\mathrm{e}^{-\alpha z}}{(n-1)!}\mathrm{d}z\frac{\lambda^{n}t^{n-1}}{(n-1)!}\mathrm{d}t$
$=$ $\int_{0}^{v}\left(\frac{\ln\frac{\left(v-z\right)\ln c+bc^{x}}{bc^{x}}}{\ln c}\right)^{n}\frac{\lambda^{n}}{n!}\left(\frac{(v-z)\ln c+bc^{x}}{bc^{x}}\right)^{-\frac{q+\lambda}{\ln c}}\frac{\alpha^{n}z^{n-1}\mathrm{e}^{-\alpha z}}{(n-1)!}\mathrm{d}z$
$=$ $\int_{0}^{t_{v}^{*}}\mathrm{e}^{-\left(q+\lambda\right)t-\alpha b_{v}^{*}}\frac{\left(b_{v}^{*}\right)^{n-1}\left(\alpha\lambda t\right)^{n}}{n!(n-1)!}bc^{x+t}\mathrm{d}t,$ (17)
for $n\geq 1$, where the third equation follows from interchanging the order of integrations, while the last equation is derived by changing of variable (letting $\frac{\ln\left(bc^{x}+z\ln c\right)-\ln b}{\ln c}-x=t$). Summing up equation (E) over $n$ ($n\geq 1$) and noting that
$\mathbb{E}_{v}\left[\mathrm{e}^{-q\tau_{Y}}\mathbf{1}\left(\tau_{Y}<\tau_{J,}N(\tau_{Y})=0\right)\right]=\mathrm{e}^{-(q+\lambda)t_{v}^{*}}$
completes that proof of equation (14).

### v3

@@
 Abstract
-The significance of mortality modeling extends across multiple research areas, including life insurance valuation, longevity risk management, life-cycle hypothesis, and retirement income planning.
-Despite the variety of existing approaches, such as mortality laws and factor-based models, they often lack compatibility or fail to meet specific research needs.
-To address these shortcomings, this study introduces a novel approach centered on modeling the dynamics of individual vitality and defining mortality as the depletion of vitality level to zero.
-More specifically, we develop a four-component framework to analyze the initial value, trend, diffusion, and sudden changes in vitality level over an individual’s lifetime.
-We demonstrate the framework’s estimation and analytical capabilities in various settings and discuss its practical implications in actuarial problems and other research areas.
-The broad applicability and interpretability of our vitality-based modeling approach offer an enhanced paradigm for mortality modeling.
+The significance of mortality modeling extends across multiple research areas, ranging from life insurance valuation to optimal lifetime decision-making.
+Existing approaches, such as mortality laws and factor-based models, often fall short in capturing the complexity of individual mortality, hindering their ability to address specific research needs.
+To overcome these limitations, this paper introduces a novel approach to mortality modeling centered on the dynamics of individual vitality.
+A four-component framework is developed to account for initial conditions, natural aging processes, stochastic fluctuations, and accidental events over an individual’s lifetime.
+We demonstrate the framework’s analytical capabilities across various settings and explore its practical implications in solving life insurance problems and deriving optimal lifetime decisions.
+Our results show that the proposed framework not only encompasses existing mortality models but also provides individualized mortality outcomes and offers an intuitive explanation for survival biases.
 1 Introduction
-Mortality modeling is a critical research topic, serving as the foundation of numerous research questions and applications in various fields.
-In actuarial science, accurate mortality models are essential for understanding and managing risks associated with life insurance products.
-They enable life actuaries to assess the underlying mortality risk of insurance policies and ensure the financial stability of life insurance companies.
-In the realm of long-term care and disability insurance, mortality models are also contributing as an essential tool, for example, in a multi-state setup to capture the complexities of transition probabilities and claim distributions.
-Beyond traditional insurance applications, mortality modeling plays an important role in retirement income planning as well.
-By estimating the survival time of individuals, mortality models can help to obtain optimal consumption and investment strategies, allowing for better financial planning and resource allocation throughout retirement.
-The ability to predict longevity and mortality trends is also crucial in pension fund management, public health planning, and demographic studies.
+Mortality modeling serves as a cornerstone for numerous research domains and practical applications, ranging from life insurance valuation and longevity risk management to optimal lifetime decision-making and public health policies.
+Accurate mortality models are essential for life insurers when evaluating and managing the demographic risks associated with life insurance products.
+These models are used to calculate prices, determine reserves, and fulfill regulatory requirements, ultimately facilitating better planning and management of life insurance products.
+Mortality modeling also plays an important role in optimal consumption and investment strategies.
+By estimating the survival time of individuals, mortality models can help develop optimal strategies for better financial planning and resource allocation.
+Additionally, the ability to predict longevity and mortality trends is crucial in pension fund management, public health planning, and demographic studies.
 In each of these areas, the reliability and interpretability of mortality modeling outcomes directly impact the quality of the decisions made.
-In the literature, a variety of mortality modeling approaches exist, each with its strengths and limitations.
-Mortality laws are among the oldest and most fundamental approaches.
+Traditional mortality modeling approaches, such as mortality laws and factor-based models, have been instrumental in performing the aforementioned tasks.
+However, they often fall short in capturing the intricate dynamics of individual mortality, particularly when dealing with the heterogeneity in health conditions and accidental events over a lifetime.
+This gap in modeling capabilities highlights the need for a more individualized approach to better account for the variability and complexity of individual mortality.
+This paper aims to address these limitations by introducing a novel mortality modeling framework centered on individual vitality dynamics.
+This framework conceptualizes mortality as the depletion of an individual’s vitality to zero and models the time of death as a first passage problem.
+By integrating multiple components into the underlying vitality process, our modeling approach delivers a more nuanced and intuitive understanding of mortality dynamics, which is important for improving actuarial practices and informing optimal decisions.
+1.1 Literature review
+Mortality laws are among the oldest and most fundamental approaches in mortality modeling.
 These laws, such as the Gompertz-Makeham law (Gompertz,, 1825; Makeham,, 1860), aim to describe age-specific mortality rates using simple mathematical functions.
-Mortality laws are widely appreciated for their simplicity and ability to capture the general trend of increasing mortality with age.
-Despite their historical significance and ease of use, these laws often fall short in capturing the complexities of real-life mortality data, especially in the presence of cohort effects and changing mortality trends over time.
-More recent studies on this approach include, for example, Gavrilov and Gavrilova, (2019), Li et al., 2021a , Di Palo, (2023) and Dang et al., (2023).
-In contrast to traditional mortality laws, the Cox proportional hazards model (Cox,, 1972) represents a significant advancement by not assuming a specific baseline hazard function.
-This semi-parametric model allows the hazard function to vary over time, influenced by covariates such as demographic, health-related, or environmental factors.
-The flexibility of the Cox model in incorporating various risk factors makes it a powerful tool in survival analysis and epidemiology.
-However, one of its limitations is the proportional hazards assumption, which implies that the covariates have a multiplicative effect on the hazard function.
-This assumption may not hold in all cases, leading to potential biases in the estimated effects of covariates.
-A comprehensive review of the Cox model and its recent developments can be found in Kalbfleisch and Schaubel, (2023).
-Another category is developed from using continuous-time stochastic processes to describe the evolution of mortality intensities.
+Mortality laws are widely appreciated for their simplicity and ability to capture the general trend of increasing mortality with age, but they often fall short in capturing the complexities of real-life mortality data.
+More recent studies on this approach include, for example, Gavrilov and Gavrilova, (2019), Li et al., 2021a , and Di Palo, (2023).
+In contrast to mortality laws, another category is developed from using stochastic processes to describe mortality intensities.
 The seminal work by Milevsky and Promislow, (2001) applied a continuous-time Cox model to survival analysis.
-Biffis, (2005) introduced the Markovian affine jump-diffusion process for modeling both financial and demographic risks.
-This model preserves analytical tractability and provides closed-form solutions to survival probabilities.
-Blackburn and Sherris, (2013) further validated the application of Markovian affine mortality models in longevity risk pricing and hedging.
-Bauer et al., (2010) focused on pricing longevity-linked securities using a forward force of mortality modeling approach.
-Other studies include, for example, Wong et al., (2017), Jevtić and Regis, (2019), Wang et al., (2021) and Zhou et al., (2022).
-Factor-based mortality models, leading by the seminal works of Lee and Carter, (1992) and Cairns et al., (2006), have become prominent due to their ability to capture underlying trends and volatility in mortality data.
-The Lee-Carter (LC) model and the Cairns-Blake-Dowd (CBD) model express mortality rates as a combination of age-specific components and time-varying factors to capture mortality changes over time.
-Significant advancements have been made in this direction, including the works of Haberman and Renshaw, (2012), Kleinow, (2015), Shang, (2019), and Hunt and Villegas, (2023), with a review provided by Cairns et al., (2009).
-Despite their popularity, these models are based on discrete-time settings, which may not be suitable for continuous-time research questions requiring explicit solutions.
-Similarly, spline-based mortality models offer a flexible non-parametric approach to mortality modeling by fitting smooth curves to mortality data.
-These models use splines, which are piece-wise polynomial functions, to capture the intricate patterns of mortality rates.
+This modeling approach preserves analytical tractability and provides closed-form solutions to survival probabilities.
+Other studies include, for example, Biffis, (2005), Bauer et al., (2010), Blackburn and Sherris, (2013), Wong et al., (2017), and Zhou et al., (2022).
+Factor-based models, led by the seminal works of Lee and Carter, (1992) and Cairns et al., (2006), have become prominent due to their ability to capture the trend and volatility in real mortality data through a combination of age-specific components and time-varying factors.
+Significant advancements have been made in this direction, including the works of Haberman and Renshaw, (2012), Kleinow, (2015), and Hunt and Villegas, (2023), with a comprehensive review provided by Cairns et al., (2009).
+Similarly, spline-based models offer a non-parametric approach to mortality modeling by fitting smooth spline curves to capture the intricate patterns of mortality data.
 Currie et al., (2004) demonstrated that the use of P-splines for smoothing mortality data provides a robust method for handling irregularities and fluctuations.
-Other studies include, for example, Dodd et al., (2018), Camarda, (2019) Hilton et al., (2019), Richards, (2020), Dodd et al., (2021), and Zhu and Zhou, (2023).
-The main advantage of spline-based models is their flexibility in fitting complex mortality patterns.
-However, their non-parametric nature can lead to overfitting and a lack of interpretability when compared to more structured models.
-Recently, an innovative approach of modeling mortality via an individual’s remaining survival energy has been proposed, distinguishing itself from traditional models that typically focus on populations or cohorts.
-Specifically, Shimizu et al., (2021) introduced the Survival Energy Model (SEM), suggesting that each individual is born with a certain initial level of survival energy that depletes stochastically over time through a time-inhomogeneous diffusion process.
-Building on this, Shimizu et al., (2023) expanded the SEM framework by incorporating more sophisticated stochastic processes to better capture the complexities of mortality dynamics.
-By focusing on the stochastic nature of an individual’s survival energy, the SEM provides a more intuitive understanding of mortality dynamics.
-We note that within the SEM framework, an individual’s death time is essentially the first time that the survival energy process reaches zero.
-Therefore, this modeling approach is closely connected to the field of first passage analysis, which is a well-established topic in applied probability with extensive applications in various fields such as operations management (see, e.g., Yamazaki, (2017) and Han et al., (2024)), quantitative finance (see, e.g., Davydov and Linetsky, (2001) and Cai and Kou, (2011)) and actuarial science (see, e.g., Lin and Willmot, (2000) and Li and Zhou, (2024)).
-We show in this paper that the depletion of an individual’s vitality (a.k.a., survival energy) exhibits similarities to the dynamics of an insurer’s cash flow.
-Specifically, the negative jumps in the vitality process represent decreases in vitality level due to accidents, analogous to how an insurer’s surplus process often takes a spectrally negative form to model claim payments.
-The concept of modeling individual vitality is not new beyond actuarial literature.
-It was pioneered by Strehler and Mildvan, (1960), who defined vitality as “the capacity of an individual organism to stay alive”.
-For a discussion on the application of the Strehler-Mildvan model to human mortality, please refer to Finkelstein, (2012) and the references therein.
-Anderson, (1992) further contributed by incorporating Brownian motion to construct a stochastic vitality process, transforming the study of mortality into a first-passage time problem.
-Numerous studies on mortality modeling have been conducted in this direction, including the works by Skiadas, (1995); Anderson, (2000); Weitz and Fraser, (2001); Aalen and Gjessing, (2004); Steinsaltz and Evans, (2004); Li and Anderson, (2009); Li et al., (2013); Li and Anderson, (2013); Skiadas and Skiadas, (2010, 2014); Li and Anderson, (2015); Sharrow and Anderson, (2016), and references therein.
-The existing studies demonstrate that vitality modeling is a robust approach to understanding mortality dynamics.
-The recent consideration of survival energy models in actuarial science highlights their potential to offer new insights into mortality modeling and forecasting.
-By integrating the stochastic nature of individual vitality and considering both deterministic and stochastic components, vitality models can provide a comprehensive framework for mortality modeling that is both flexible and intuitive.
-The main objective of this paper is to show that vitality modeling can become a valuable tool in actuarial science, addressing a number of research problems and enhancing the solutions provided by existing methods.
-Moreover, vitality modeling offers an alternative, more intuitive understanding of mortality dynamics, which can complement and improve upon traditional actuarial approaches.
+Other studies include, for example, Dodd et al., (2018), Camarda, (2019), Dodd et al., (2021), and Zhu and Zhou, (2023).
+1.2 Motivation
+Recently, a new approach of modeling mortality has been proposed by Shimizu et al., (2021) and Shimizu et al., (2023).
+Specifically, Shimizu et al., (2021) introduces the Survival Energy Model (SEM), where each individual is assumed to have a survival energy level that depletes stochastically over time through a time-inhomogeneous diffusion process.
+Shimizu et al., (2023) expanded the SEM framework by using an inverse Gaussian process with an explicit mortality function.
+By focusing on the stochastic nature of an individual’s survival energy, the SEM provides an intuitive way of understanding mortality dynamics.
+Within the SEM framework, an individual’s death time is essentially the first passage time that the survival energy process crosses zero.
+This modeling approach is closely connected to the field of first passage analysis, a well-established topic in applied probability with extensive applications in operations management (see, e.g., Yamazaki,, 2017; Han et al.,, 2024), quantitative finance (see, e.g., Davydov and Linetsky,, 2001; Cai and Kou,, 2011), and actuarial science (see, e.g., Lin and Willmot,, 2000; Asmussen and Albrecher,, 2010).
+For instance, the depletion of an individual’s survival energy is similar to the dynamics of an insurer’s cash flow, where a negative jump in survival energy can be equivalently viewed as an insurer’s surplus process taking a spectrally negative form.
+The concept of modeling individual survival energy, also known as vitality, is not new beyond financial literature.
+It was pioneered by Strehler and Mildvan, (1960), who defined vitality as “the capacity of an individual organism to stay alive.” Anderson, (1992) further contributed by using a Brownian motion to describe vitality dynamics.
+Along this direction, numerous studies have been conducted, including the works by Anderson, (2000), Aalen and Gjessing, (2004), Li and Anderson, (2009), Skiadas and Skiadas, (2010), Sharrow and Anderson, (2016), and Anderson et al., (2017).
+A discussion on the application of the Strehler-Mildvan model to mortality modeling is available in Finkelstein, (2012).
+1.3 Our contribution
+The objective of this paper is to explore a novel approach to mortality modeling through the lens of individual vitality dynamics.
+By conceptualizing death as the depletion of an individual’s vitality to zero, we aim to derive mortality modeling outcomes that are individualized for unique health conditions, lifestyle factors, and random life events.
+We also attempt to apply the vitality-based modeling approach to a variety of emerging research problems.
+Specifically, this paper makes three significant contributions.
 The first contribution of this paper is the introduction of a vitality-based framework for mortality modeling.
-This framework conceptualizes mortality through the lens of individual vitality, where death is defined as the point at which an individual’s vitality depletes to zero.
-The fundamental idea is that each person is born with an initial level of vitality, which decreases over time due to both natural aging and external factors.
-The framework is built around four intuitive components: the initial vitality level, natural depletion trends, stochastic diffusion process, and sudden jumps.
-The resulting framework offers a comprehensive and flexible approach to mortality modeling and is capable of accommodating various features and ideas related to mortality dynamics.
-This flexibility sets the stage for our second contribution, where we demonstrate how different specifications of the proposed framework can lead to diverse and insightful modeling outcomes.
-The second contribution of this paper lies in the various specifications of the proposed modeling framework, which is capable of representing existing mortality models, including the Gompertz law and the Cairns-Blake-Dowd model.
-By transforming these models into a unified vitality-based structure, we enhance their ability to capture more complex mortality dynamics, such as jumps to account for accidental deaths and diffusion to reflect the randomness in natural life decay.
-Under each specification, we derive survival probabilities and other relevant properties to showcase the insightful modeling outcomes, setting the stage for our third contribution on actuarial applications.
-The third contribution of this paper demonstrates the practical applications of the vitality-based modeling framework in actuarial science.
-We showcase that this framework can be employed in various actuarial tasks and has the potential to solve existing problems and improve traditional methods.
-Specifically, we discuss how vitality can be incorporated in calculating life expectancies, pricing life insurance products, and deriving optimal consumption decisions.
-Additionally, our vitality-based framework is utilized to model disability and recovery probabilities, analyze different causes-of-death, and explain the under- and over-estimation of survival probabilities throughout an individual’s life course.
+This framework assumes that each individual has an initial vitality that depletes to zero over time due to both natural causes and external factors.
+It comprises four components: initial vitality level, natural depletion trends, stochastic diffusion, and sudden jumps.
+Our modeling approach offers a comprehensive and flexible method for capturing various aspects of mortality dynamics.
+The second contribution lies in the various specifications of the proposed vitality-based framework.
+We demonstrate that certain specifications can represent existing mortality models, while others can generate more diverse modeling outcomes.
+Under each specification, we derive survival probabilities and other relevant properties to illustrate how the vitality-based framework can complement and improve existing models in describing complex mortality dynamics at the individual level.
+The third contribution of this paper relates to the practical applications of the vitality-based framework across several research domains.
+Specifically, we discuss how individual vitality can be incorporated into calculating life expectancies, pricing insurance products, and deriving optimal consumption decisions.
+Additionally, we derive disability and recovery probabilities, analyze different causes of death, and explain the under- and overestimation of survival probabilities under our vitality-based framework.
 The structure of this paper is organized as follows.
-Section 2 details the proposed modeling framework, explaining its four key components: initial vitality level, natural depletion trends, stochastic diffusion processes, and sudden jumps.
-Section 3 covers the various specifications of the proposed framework, including static and dynamic cases, and offers specific examples to illustrate its estimation and analytical capabilities.
-Section 4 discusses the applications of the framework in actuarial science, such as life insurance valuation, disability modeling, and optimal consumption decisions.
-Section 5 presents an illustration on the numerical estimation of a vitality-based mortality model.
-Finally, Section 6 concludes the paper with a discussion on the limitations and suggestions for future research.
-2 The Proposed Modeling Framework
-In this section, we introduce the vitality-based mortality modeling framework, detailing its fundamental structure and components.
-The proposed framework conceptualizes mortality through the depletion of an individual’s vitality to zero and utilizes four components to describe the dynamics of the underlying vitality.
-Each component captures different aspects of vitality dynamics, jointly providing a comprehensive and flexible approach to mortality modeling.
+Section 2 details the proposed vitality-based mortality modeling framework.
+Section 3 covers the various specifications of the proposed framework.
+Section 4 discusses how the vitality-based framework can be incorporated in various applications.
+Section 5 presents an illustration on the estimation of the proposed framework.
+Finally, Section 6 concludes the paper with a discussion on limitations and future research.
+2 The Proposed Framework
+In this section, we introduce the vitality-based mortality modeling framework along with its fundamental structure and components.
 Consider an individual who is age $x$ at time 00 (the time of modeling), and let $V(t)$ be the vitality of this individual at time $t\geq 0$.
@@
 The first component of our vitality-based modeling framework is the initial vitality of the individual being modeled, denoted as $V(0)$.
-This component is crucial as it sets the baseline from which all future vitality changes are measured.
-If the age $x$ is zero, then $V(0)$ represents the vitality level of a newborn.
+This component sets the baseline from which all future vitality changes are measured.
+If the individual’s age $x$ is zero, then $V(0)$ represents the vitality level of a newborn.
 Otherwise, $V(0)$ can be interpreted as the remaining vitality level of the individual $x$ years after birth.
 We can model $V(0)$ either as a fixed value or a random variable.
-A fixed value of $V(0)$ assumes that the level of initial vitality is known and also the same for all the individuals being modeled (i.e., all individuals who are age $x$ at time 00).
-This assumption might be reasonable if the cohort of individuals is homogeneous.
-For example, we can assume that all newborns from a certain birth year will have the exact same initial vitality.
-This simplification helps in creating a more streamlined model for analysis and is particularly useful in scenarios where there is little variation in initial health conditions within a cohort.
-In a more realistic setting, the value of $V(0)$ should be treated as random to represent the inherent variability in the vitality level of individuals.
-This randomness reflects the diversity in individuals within a cohort, such as genetics, lifestyles, and other factors that affect an individual’s initial health condition.
-We should thus model $V(0)$ as a random variable with
+A fixed value of $V(0)$ assumes that the level of initial vitality is known and also the same for all of the individuals being modeled.
+This setting may be reasonable if the group of individuals is homogeneous, for example, a portfolio of selected insurance policyholders.
+We may also assume that all newborns from a certain birth year will have the exact same initial vitality.
+In a more realistic setting, the value of $V(0)$ should be treated as random to represent the inherent variability in the initial vitality level of individuals.
+This randomness reflects the diversity in individuals within a group, such as genetics, lifestyles, and other factors that affect an individual’s initial health condition.
+We would thus model $V(0)$ as a random variable with
 $V(0)\sim F_{0},$
 where $F_{0}$ is a distribution that could be, for example, exponential or Pareto.
-This approach acknowledges that even within a seemingly homogeneous cohort, individuals will have different initial vitality levels due to various reasons.
-Furthermore, the remaining vitality level after $x$ years of living will differ among individuals due to their personal lifestyle, living environment, health status, and other individual-specific factors.
-Even for newborns, the initial vitality should vary within an inhomogeneous cohort to reflect different birth traits, such as genetic disorders, birth defects, and pregnancy conditions.
-This variability is essential for constructing a model that can accurately reflect individual differences, and also for building a linkage between existing mortality models and vitality-based models.
+Consider a group of individuals all aged $x>0$.
+Their remaining vitality after $x$ years of living will differ due to their personal lifestyle, living environment, and other individual-specific factors.
+Even for newborns with age $x=0$, the initial vitality should vary within a cohort year to reflect different birth traits, such as genetic disorders and birth defects.
+The random setting of $V(0)$ is useful for constructing a model that can reflect individual differences, and also for building a linkage between existing mortality models and vitality-based models.
 Component 2 (C2): The trend $Y(t)$
-The second component describes the decline in an individual’s vitality over time due to the natural aging process.
+The second component describes the decline in an individual’s vitality due to the natural aging process.
 We denote $Y(t)$ to represent the depletion trend in vitality that occurs as the individual ages beyond time 00.
-To reflect the general trend of vitality depletion in the absence of stochastic variations or sudden shocks, $Y(t)$ can be modeled as a deterministic function of time.
-In particular, we assume that
+First, $Y(t)$ can be modeled as a deterministic function of time.
+In particular, we can assume that
 $\mathop{}\!\mathrm{d}Y(t)=\mu(t)\,\mathop{}\!\mathrm{d}t,\quad t\geq 0,$
-where $\mu(t)$ is the depletion rate of vitality at time $t$ when the individual is aged $x+t$.
-The exact choice of $\mu(t)$ should be informed by both theoretical considerations and empirical studies.
+where $\mu(t)$ is the vitality depletion rate at time $t$ when the individual is aged $x+t$. The exact choice of $\mu(t)$ should be informed by both theoretical considerations and empirical studies.
 The function $\mu(t)$ can be specified in various forms, depending on the specific characteristics and assumptions made.
@@
 For instance, an exponential decline might be used to represent a situation where the rate of vitality loss accelerates with age.
-Another example would be using the Gompertz law to specify $\mu(t)$; that is, $\mu_{x}(t)=bc^{x+t}$ where $b$ and $c$ are the conventional Gompertz parameters.
-We added a subscript of age $x$ to $\mu(t)$ to align with the typical notation of mortality laws and to emphasize that the depletion rate, given by the Gompertz law, depends on the individual’s age at time $t$.
+Another example would be using the Gompertz law to specify $\mu(t)$; that is, $\mu(t)=bc^{x+t}$ where $b$ and $c$ are the conventional Gompertz parameters.
 Alternatively, we may assume that the depletion rate is uniform over time; that is, $\mu(t)=\delta$ with $\delta$ being a constant.
@@
 Component 3 (C3): The diffusion $W(t)$
-Acknowledging that an individual’s life is full of uncertainties and thus vitality should not be depleted in a perfectly predictable manner, the third component introduces a certain degree of uncertainty to the vitality dynamics of the individual.
-From the modeling perspective, given that the trend component $Y(t)$ exhibits a deterministic nature, it is reasonable to consider a diffusion component, denoted as $W(t)$, to reflect the fact that the vitality level of an individual can fluctuate randomly around its depletion trend.
-This stochastic variation will account for any minor health fluctuations that an individual may experience over time.
+The third component introduces a certain degree of uncertainty to the vitality dynamics of the individual.
+From the modeling perspective, given that the trend component $Y(t)$ has a deterministic nature, it is reasonable to consider a diffusion component, denoted as $W(t)$, to reflect the fact that the vitality level of an individual can fluctuate randomly around its depletion trend.
+This stochastic variation will account for any vitality fluctuations that an individual may experience over time.
 One simple way to model $W(t)$ is to use a Brownian motion; that is,
@@
 where $B(t)$ is a standard Brownian motion and $\sigma(t)$ is a diffusion parameter at time $t$.
-A more advanced stochastic process can be considered to replace $B(t)$, such as a fractional Brownian motion, for capturing the long-term dependence in the individual’s health status over time.
-This setting is particularly useful when the individual has health conditions with a lingering effect on future vitality.
-Similar to the depletion rate, a piece-wise function of $W(t)$ can also be considered to reflect that different life periods may face different levels of vitality fluctuations.
-With the diffusion component, we can simulate stochastic paths of individual vitality over time to provide insights into the variability and uncertainty inherent in the aging process.
-The simulations will generate a variety of potential future vitality trajectories, from which we can observe how different individuals’ path diverge from their deterministic trend.
-These simulated paths particularly help in understanding the probabilistic nature of vitality depletion and are crucial for applications such as deriving personalized optimal consumption decisions.
+A more advanced stochastic process can be considered to replace $B(t)$, such as a fractional Brownian motion for capturing the long-term dependence in the individual’s health status over time.
+Additionally, a piece-wise function of $W(t)$ can also be considered to reflect that different life periods may face different levels of vitality fluctuations.
+With the diffusion component, we can simulate random trajectories of vitality over time to provide insights into the variability and uncertainty inherent in the aging process.
+These simulated trajectories are particularly helpful in understanding the probabilistic nature of vitality depletion and are crucial for applications such as deriving personalized optimal consumption decisions.
+Moreover, we can use simulated trajectories to analyze how different individuals’ vitality diverge from their deterministic trend.
 Component 4 (C4): The jump $J(t)$
@@
 These events can cause sudden and significant changes to an individual’s vitality level, but cannot be captured by the depletion trend or the stochastic diffusion.
-To address this, the last component of our proposed modeling framework is the jump component, denoted as $J(t)$.
-This component is designed to incorporate the impact of accidental events into the vitality dynamics via a jump process; that is,
+To address this issue, we let the last component of our vitality-based modeling framework to be a jump component, denoted as $J(t)$.
+To incorporate the impact of accidental events into the vitality dynamics, we apply a jump process to $J(t)$; that is,
 $J(t)=\sum_{i=1}^{N(t)}Z_{i},\quad t\geq 0,$
@@
 In a simplified setting, we may assume that $Z_{i}=\infty$ to represent the event of a fatal accident, where the individual’s vitality drops below zero instantaneously.
-The jump component explains the variability in mortality due to accidental events that are not accounted for by gradual trends or random fluctuations.
-Together with the initial, trend and diffusion components, our proposed vitality-based framework is capable of providing insights into how a wide range of lifetime phenomena contribute to human mortality, including natural deaths, random health fluctuations, accidental events, and pre-existing medical conditions.
-By specifying these four components differently, our framework offers flexible modeling outcomes that encompass many existing mortality models.
-These different specifications and their implications will be discussed in the next section.
 3 Model Specifications
-In this section, we provide a number of specific cases in which the proposed vitality-based framework is implemented to arrive at various mortality modeling outcomes.
-The proposed framework is designed to be modulable, meaning that the four components can be selectively chosen based on different modeling objectives.
-For instance, by controlling certain components, the framework can replicate traditional mortality laws, offering an alternative approach to basic mortality analysis.
-Conversely, by incorporating all components, the resulting model is capable of capturing more complex phenomena that influence mortality dynamics over time.
-Through the specific cases, we aim to demonstrate the capabilities of the vitality-based framework across different modeling needs and highlight its potential to enhance our understanding of mortality patterns.
-3.1 Static cases
-A vitality mortality model is static if it only governs the mortality dynamics over an individual’s lifetime but does not account for general changes in population mortality over time.
-We discuss the dynamic cases in the next subsection, where period and cohort effects are captured and reflected by a vitality mortality model.
-3.1.1 Traditional mortality laws
-Let us consider the following specification of the vitality-based mortality modeling framework:
+In this section, we present several vitality models to illustrate how the proposed framework can be tailored to achieve different mortality modeling outcomes and also to enhance existing mortality models.
+Figure 1 outlines how the remainder of this section is structured and also highlights the connection between our vitality models and some existing models.
+Vitality Mortality ModelStatic ParametersDynamic Parameters Plateau Models Mortality Laws Jump-Diffusion Models M5 Model M6 Model M8 Model Section 3.1.1Section 3.1.2Section 3.1.3Section 3.2.1Section 3.2.2Section 3.2.3
+Figure 1: Connections between vitality mortality models and existing mortality models.
+3.1 Models with static parameters
+We begin from vitality models that are built by static parameters.
+The first goal is to establish a connection between traditional mortality laws and vitality models.
+We then extend those vitality models to include diffusion and jump components and derive analytical solutions for the resulting survival probabilities.
+3.1.1 Mortality laws
+Let us consider the following specification of the proposed framework:
   • C1
-    The initial vitality $V(0)$ follows an exponential distribution with unit scale; that is,
-    $V(0)\sim\text{Exp}(1).$
+    The initial vitality $V(0)$ follows an exponential distribution with unit scale; that is, $V(0)\sim\text{Exp}(1)$.
   • C2
@@
   • C2
-    The trend component $Y(t)$ is governed by a mortality law $\mu_{x}(t)$:
-    $\mathop{}\!\mathrm{d}Y(t)=\mu_{x}(t)\mathop{}\!\mathrm{d}t,\quad t\geq 0.$
+    The trend component $Y(t)$ is governed by a mortality law $\mu_{x}(t)$: $\mathop{}\!\mathrm{d}Y(t)=\mu_{x}(t)\mathop{}\!\mathrm{d}t$.
 The resulting $T$-year survival probability of the above modeling settings is
@@
 Missov, (2013) has demonstrated that the Gamma-Gompertz model reflects the plateau effect such that the implied force of mortality converges to a constant as age increases.
-An alternative specification for capturing the plateau effect and resulting in equation (3) for the $T$-year survival probability is
+An alternative specification for capturing the plateau effect is
   • C1
-    The initial vitality $V(0)$ follows an exponential distribution with unit scale:
-    $V(0)\sim\text{Exp}(1).$
+    The initial vitality follows an exponential distribution with unit scale: $V(0)\sim\text{Exp}(1)$.
   • C2
-    The trend component $Y(t)$ is given by
-    $\mathop{}\!\mathrm{d}Y(t)=Z\cdot\mu_{x}(t)\mathop{}\!\mathrm{d}t,\quad t\geq 0,$
-    where $\mu_{x}(t)$ is specified by a mortality law and $Z$ follows a Gamma distribution with shape parameter $\alpha$ and rate parameter $\phi$.
+    The trend component is given by $\mathop{}\!\mathrm{d}Y(t)=Z\mu_{x}(t)\mathop{}\!\mathrm{d}t$, where $\mu_{x}(t)$ is specified by a mortality law and $Z$ follows a Gamma distribution with shape parameter $\alpha$ and rate parameter $\phi$.
+The resulting $T$-year survival probability of the above specification is the same as equation (3) (see Appendix A.1 for more details).
 These two specifications have distinct interpretations.
-The first one describes the initial vitality level of the underlying population to be more dispersed and the reason for the death plateau is due to a larger proportion of the population exhibiting a large initial vitality level.
+The first one describes the initial vitality level of the underlying population as being more dispersed and the reason for the death plateau is due to a larger proportion of the population exhibiting a large initial vitality level.
 The second one implies that the death plateau is indeed due to the heterogeneity in the depletion rate.
+While it is possible to find other formulations that yield the same survival probabilities, our study does not aim to be exhaustive in covering all possible formulations.
+We provide an additional formulation at the end of Appendix A.1.
 3.1.3 A jump-diffusion vitality model
@@
 Second, the natural vitality decay specified by $\mu_{x}(t)$ is deterministic, which means the randomness that existed in the individual’s health status over time is ignored.
-To introduce randomness, Shimizu et al., (2021) and Shimizu et al., (2023) have considered applying a Brownian motion and an inverse-Gaussian process, respectively, to govern an individual’s vitality.
+To introduce randomness, Shimizu et al., (2021) and Shimizu et al., (2023) have considered applying a diffusion process and an inverse-Gaussian process, respectively, to govern an individual’s vitality.
 However, these studies assume a deterministic initial vitality and an arbitrarily chosen trend component, which consequently fails to connect the vitality-based approach with traditional mortality laws.
-One disadvantage of using inverse Gaussian processes is that their implied infinite number of small jumps may not be suitable for characterizing the decrement of vitality levels due to accidents.
+Moreover, one disadvantage of using inverse Gaussian processes is that their implied infinite number of small jumps may not be suitable for characterizing the decrement of vitality levels due to accidents.
 Under the proposed framework, we now consider adding a Brownian motion in the diffusion component to reflect random fluctuations in vitality dynamics, and a compound Poisson process in the jump component to reflect the possibility of accidental events.
-In particular, the following specification is used:
+In particular, a jump-diffusion vitality mortality model is specified as follows:
   • C1
-    The initial vitality $V(0)$ follows a distribution $F_{0}$ with positive support:
-    $V(0)\sim F_{0}.$
+    The initial vitality $V(0)$ follows a distribution $F_{0}$ with positive support: $V(0)\sim F_{0}$.
   • C2
-    The trend component $Y(t)$ is governed by
-    $\mathop{}\!\mathrm{d}Y(t)=\mu_{x}(t)\mathop{}\!\mathrm{d}t,\quad t\geq 0,$
-    where $\mu_{x}(t)$ is specified by a mortality law (e.g., $\mu_{x}(t)=bc^{x+t}$ for the Gompertz law).
+    The trend component $Y(t)$ is governed by $\mathop{}\!\mathrm{d}Y(t)=\mu_{x}(t)\mathop{}\!\mathrm{d}t$, where $\mu_{x}(t)$ is specified by a mortality law.
   • C3
-    The diffusion component $W(t)$ is specified as
-    $W(t)=\sigma B(t),\quad t\geq 0,$
-    where $B(t)$ is a Brownian motion and $\sigma$ is the diffusion parameter.
+    The diffusion component $W(t)$ is specified as $W(t)=\sigma B(t)$, where $B(t)$ is a Brownian motion and $\sigma$ is the diffusion parameter.
   • C4
-    The jump component $J(t)$ is specified as
-    $J(t)=\sum_{i=1}^{N(t)}Z_{i},\quad t\geq 0,$
-    where $N(t)$ is a Poisson process with intensity rate $\lambda(t)$, and $\{Z_{i}\}_{i\geq 1}$ is a sequence of i.i.d. positive random variables with distribution function $F_{Z}(\cdot)$.
-It follows that the dynamics of $V(t)$ is governed by
-$V(t)=V(0)-\int_{0}^{t}\mu_{x}(s)\mathop{}\!\mathrm{d}s-\sigma B(t)-\sum_{i=1}^{N(t)}Z_{i}.$ (4)
+    The jump component $J(t)$ is specified as $J(t)=\sum_{i=1}^{N(t)}Z_{i}$, where $N(t)$ is a Poisson process with intensity rate $\lambda(t)$, and $\{Z_{i}\}_{i\geq 1}$ is a sequence of independent and identically distributed (i.i.d.) positive random variables with distribution function $F_{Z}(\cdot)$.
+The dynamics of $V(t)$ is consequently governed by
+$V(t)=V(0)-\int_{0}^{t}\mu_{x}(s)\mathop{}\!\mathrm{d}s-\sigma B(t)-\sum_{i=1}^{N(t)}Z_{i},\quad t\geq 0.$ (4)
 We assume that $V(0)$, $W(t)$, and $J(t)$ are independent of each other. The resulting $T$-year survival probability can be expressed as
 $\Pr(\tau>T)=\int_{0}^{\infty}\Pr\left(v-\sigma B(t)-\sum_{i=1}^{N(t)}Z_{i}\geq\int_{0}^{t}\mu_{x}(s)\mathop{}\!\mathrm{d}s,\;\forall t\leq T\right)\mathop{}\!\mathrm{d}F_{0}(v).$ (5)
-We would like to point out that deriving analytical expressions for the survival probability is generally challenging in this case, unless specific forms of $Y(t)$ are assumed such as linear functions.
+We remark that deriving analytical expressions for the survival probability is generally challenging in this case, unless specific forms of $Y(t)$ are assumed such as linear functions.
 To elaborate further, the main difficulty lies in deriving the law of the first passage time of $B(t)$ over a moving boundary, which is a non-trivial problem in applied probability (see, e.g., Salminen, (1988)).
@@
 In such cases, one may resort to numerical methods to approximate the survival probability of a jump-diffusion vitality model.
-In this paper, we employed the Monte Carlo method as suggested in Jin and Wang, (2017) to numerically evaluate the survival probability as indicated in Eq.
-(5), and provided the details in Appendix B.
-3.1.4 An alternative specification
+In this paper, we employed the Monte Carlo method as suggested in Jin and Wang, (2017) to numerically evaluate the survival probability as indicated in equation (5), and provided the details in Appendix B.
 In Section 3.1.2, we illustrated two model specifications that equally describe the death plateau.
 It is not surprising that more than one specification of the vitality-based modeling framework can provide outcomes that align with mortality laws.
-In this subsection, we formulate an alternative specification that also connects with mortality laws.
+We now formulate an alternative specification that also connects with mortality laws.
 The following specification is assumed:
   • C1
-    The initial vitality $V(0)$ follows a Gompertz distribution with shape parameter $\eta$ and scale parameter being one. The distribution function of this Gompertz distribution is
-    $F(v)=1-e^{-\eta\left(e^{v}-1\right)}.$
+    The initial vitality $V(0)$ follows a Gompertz distribution with shape parameter $\eta$ and scale parameter being one.
+The distribution function of this Gompertz distribution is $F(v)=1-e^{-\eta\left(e^{v}-1\right)}.$
   • C2
-    The trend component $Y(t)$ has a linear decay at the rate of $\delta$; that is,
-    $\mathop{}\!\mathrm{d}Y(t)=\delta\mathop{}\!\mathrm{d}t,\quad t\geq 0,$
+    The trend component $Y(t)$ has a linear decay at the rate of $\delta$; that is, $\mathop{}\!\mathrm{d}Y(t)=\delta\mathop{}\!\mathrm{d}t$.
 We can show that the resulting $T$-year survival probability is
@@
 where, to obtain the Gompertz law’s expression of survival probability, the Gompertz parameters $b$ and $c$ need to be parameterised by $\eta=\frac{bc^{x}}{\ln(c)}$ and $\delta=\ln(c)$.
-To introduce randomness in the evolution of an individual’s vitality over time, we can additionally assume that
-  • C3
-    The diffusion component $W(t)$ is specified as
-    $W(t)=\sigma B(t),\quad t\geq 0,$
-    where $B(t)$ is a Brownian motion and $\sigma>0$ is the diffusion parameter.
-  • C4
-    The jump component $J(t)$ is specified as
-    $J(t)=\sum_{i=1}^{N(t)}Z_{i},\quad t\geq 0,$
-    where $N(t)$ is a Poisson process with intensity rate $\lambda(t)$, and $\{Z_{i}\}_{i\geq 1}$ is a sequence of i.i.d. positive random variables.
-It follows that the dynamics of $V(t)$ is governed by
+To introduce randomness in the evolution of an individual’s vitality over time, we can add the diffusion component and the jump component that were previously assumed in this subsection to formulate the jump-diffusion vitality mortality model again.
+The dynamics of $V(t)$ will be governed by
 $V(t)=V(0)-\delta t-\sigma B(t)-\sum_{i=1}^{N(t)}Z_{i},\quad t\geq 0.$ (7)
-One special case worth noting is when the jump size $Z_{i}=\infty$ (i.e., each jump represents a fatal event). In this case, we would have
+One special case worth noting is when the jump represents a fatal event and the survival probability can be analytically expressed in the following proposition.
+Proposition 1.
+Given the vitality model (7) with $Z_{i}=\infty$, the survival probability can be analytically solved as
 $\Pr(\tau>T)$ $={e}^{-\int_{0}^{T}\lambda(t)\mathop{}\!\mathrm{d}t}\cdot\Pr\left(V(0)-\delta t-\sigma B(t)>0\text{ for all }t\in[0,T]\right)$
 $={e}^{-\int_{0}^{T}\lambda(t)\mathop{}\!\mathrm{d}t}\cdot\int_{0}^{\infty}\Phi\left(\frac{v-\delta T}{\sigma\sqrt{T}}\right)-\exp\left(\frac{2\delta v}{\sigma^{2}}\right)\Phi\left(\frac{-v-\delta T}{\sigma\sqrt{T}}\right)\mathop{}\!\mathrm{d}F_{0}(v),$ (8)
-where the last equation follows from the well-known law of the first passage time of a Brownian motion with linear drift (see, e.g., Siegmund, (1986)).
+where $\Phi(\cdot)$ is the distribution function of a standard Gaussian distribution.
+Proposition 1 directly follows from the law of the first passage time of a Brownian motion with linear drift (see, e.g., Siegmund, (1986)), and it is thus omitted here.
 In general, conditioning on $V(0)$, $V(t)$ is a spectrally negative Lévy process (SNLP), and the Laplace transform of $\tau$ is known explicitly.
@@
 $\tilde{V}(t)=\exp\left(V(t)\right)=\exp\left(V(0)-Y(t)-W(t)-J(t)\right),\quad t\geq 0,$
-and the individual dies if $\tilde{V}(t)$ cross the threshold $D=1$,
-$\tau=\inf\left\{t\geq 0:\tilde{V}(t)\leq 1\right\}.$
 […]
-3.2 Dynamic cases
-A vitality mortality model is called dynamic when it adapts to changes in mortality over time, thereby explaining the period and cohort effects of mortality data.
-In stochastic mortality modeling, the period effect refers to changes in mortality over time across ages, while the cohort effect refers to changes in mortality across birth years that remain consistent over the cohort’s lifetime.
-To capture period and cohort effects, time-varying parameters are needed in a vitality mortality model, similar to the structure of the LC model and the CBD model.
-Similar to the static cases where we established a linkage between vitality models and mortality laws, in this subsection we aim to build a bridge between the dynamic vitality mortality models and the popular stochastic mortality models.
-In particular, we begin by examining a dynamic extension of the Gompertz model, which is connected to the CBD model.
-We then expand our investigation to other mortality models and also consider cohort effects.
-We demonstrate that the proposed vitality-based framework is capable of providing new insights and enhancements to existing models.
-3.2.1 A dynamic vitality model
+3.2 Models with dynamic parameters
+In this subsection, we build vitality models that use on dynamic parameters and consequently have a connection with stochastic mortality models.
+We begin by examining a dynamic extension of the Gompertz law, and then expand our investigation to models with age and cohort effects.
+3.2.1 A vitality model with period effect
 We start from a straightforward example, where the depletion rate of vitality is dynamically specified with time-varying parameters.
@@
   • C1
-    The initial vitality $V(0)$ follows an exponential distribution with unit scale; that is, $V(0)\sim\text{Exp}(1)$.
+    The initial vitality $V(0)$ follows an exponential distribution with unit scale: $V(0)\sim\text{Exp}(1)$.
   • C2
@@
 $\sigma_{Y}(t)$ $=\sigma_{b}^{2}+(x+t)^{2}\cdot\sigma_{c}^{2}+2(x+t)\cdot\rho\cdot\sigma_{b}\cdot\sigma_{c},$
-Since $Y(t)$ is a non-decreasing process, the $T$-year survival probability can be expressed as
-$\Pr(\tau>T)$ $=\mathbb{E}\left[\exp\left(-Y(T)\right)\right].$
+Since $Y(t)$ is a non-decreasing process, the $T$-year survival probability can be expressed as $\Pr(\tau>T)=\mathbb{E}\left[\exp\left(-Y(T)\right)\right]$.
 The dynamic Gompertz law presented above is closely related to the Cairns-Blake-Dowd (CBD) model, which describes mortality over time and age for a population of individuals.
@@
 $\ln\mu(x,t)$ $=\kappa_{1}(t)+\kappa_{2}(t)\cdot(x-\bar{x}),\quad t\geq 0.$
-The $T$-year survival probability of an individual aged $x$ at time 00 under this continuous version of the CBD model is
-$\mathbb{E}\left[\exp\left(-\int_{0}^{T}\mu(x+t,t)\mathop{}\!\mathrm{d}t\right)\right],$
-where
-$\mu(x+t,t)=\exp\left(\kappa_{1}(t))-\kappa_{2}(t)\cdot\bar{x}\right)\cdot\left[\exp\left(\kappa_{2}(t)\right)\right]^{(x+t)}.$
+The $T$-year survival probability of an individual aged $x$ at time 00 under this continuous version of the CBD model is $\mathbb{E}\left[\exp\left(-\int_{0}^{T}\mu(x+t,t)\mathop{}\!\mathrm{d}t\right)\right]$, where
+$\mu(x+t,t)=\exp\left(\kappa_{1}(t)-\kappa_{2}(t)\cdot\bar{x}\right)\cdot\left[\exp\left(\kappa_{2}(t)\right)\right]^{x+t}.$
 Clearly, if $\kappa_{1}(t)$ and $\kappa_{2}(t)$ are modeled as a two-dimensional random walk with drifts (see, e.g., Li et al., 2021b, ), then the parameters in the process of $\ln b(t)$ and $\ln c(t)$ can be chosen to match with the continuous CBD model as
-$b(t)=\exp\left(\kappa_{1}(t))-\kappa_{2}(t)\cdot\bar{x}\right),\quad c(t)=\exp\left(\kappa_{2}(t)\right).$
+$b(t)=\exp\left(\kappa_{1}(t)-\kappa_{2}(t)\cdot\bar{x}\right),\quad c(t)=\exp\left(\kappa_{2}(t)\right).$
 The resulting $T$-year survival probabilities would be the same between our vitality mortality model and the continuous version of the CBD model.
@@
 The approximated moment generating function from Asmussen et al., (2016) is used to value the expectation $\mathbb{E}[\exp(-Y(t))]$.
-Our dynamic vitality mortality model can be constructed to have the same survival probability as other stochastic mortality models, whenever the force of mortality is guaranteed to be non-negative.
+Our vitality mortality model can be constructed to have the same survival probability as other stochastic mortality models, whenever the force of mortality is guaranteed to be non-negative.
 For instance, consider the Lee-Carter model:
 $\ln\mu(x,t)=\alpha(x)+\beta(x)\kappa(t),\quad t\geq 0.$
-If $\alpha(x)$ and $\beta(x)$ are specified as continuous deterministic functions and $\kappa(t)$ is governed by a continuous random walk with drift, then a dynamic vitality mortality model can be established to match the likelihood of the Lee-Carter model.
+If $\alpha(x)$ and $\beta(x)$ are specified as continuous deterministic functions and $\kappa(t)$ is governed by a continuous random walk with drift, then a vitality mortality model can be established to match the likelihood of the Lee-Carter model.
 Another example is the Cox-Ingersoll-Ross (CIR) model for modeling the log force of mortality, as discussed in Huang et al., (2022) and Ungolo et al., (2024), where the likelihood aligns with our vitality model if we maintain the same specification for the depletion rate $\mu_{x}(t)$.
@@
 Since $F_{0}(v)$ and $\Pr(J(T)\leq v-u)$ can be explicitly expressed and $F_{Y}(\cdot)$ can be approximated as the distribution function of a single log-normal random variable, the integral for obtaining $\Pr(\tau>T)$ can be approximated numerically.
-3.2.2 A dynamic vitality model with cohort effect
+3.2.2 A vitality model with period-cohort effect
 In a factor-based mortality model, the cohort effect is captured by an additive term, often denoted by $\gamma$, such as in the Renshaw-Haberman model (M2), the Age-Period-Cohort model (M3), the CBD model with cohort effect (M6), the CBD model with quadratic and cohort effects (M7), and the CBD model with age-dependent cohort effect (M8), as outlined in Cairns et al., (2009).
-In a vitality-based mortality model, the cohort effect can be naturally embedded in the initial vitality to reflect changes in the average healthiness of each generation.
-We illustrate the idea of a dynamic vitality model with cohort effect based on the dynamic Gompertz law with cohort effect, which corresponds to the CBD model with cohort effect (M6).
+In a vitality mortality model, the cohort effect can be naturally embedded in the initial vitality to reflect changes in the average healthiness of each generation.
+We illustrate the idea of a vitality model with cohort effect based on the dynamic Gompertz law with cohort effect, which corresponds to the CBD model with cohort effect (M6).
 Denote $V_{y}(t)$ as the time-$t$ vitality level of an individual born in year $y$ with an age of $x$ at time 00.
@@
   • C2
-    The trend component $Y(t)$ is specified as in the dynamic Gompertz model:
-    $\mathop{}\!\mathrm{d}Y(t)=b(t)c(t)^{x+t}\mathop{}\!\mathrm{d}t,\quad t\geq 0.$
+    The trend component $Y(t)$ is specified as in the dynamic Gompertz model: $\mathop{}\!\mathrm{d}Y(t)=b(t)c(t)^{x+t}\mathop{}\!\mathrm{d}t$, $t\geq 0$.
 The log transforms of $c(t)$ and $b(t)$ are governed by
@@
 $\Pr(\tau_{y}>T)$ $=\mathbb{E}\left[\exp\left(-\Gamma(y)Y(T)\right)\right],$
-where $\tau_{y}$ is the first passage time for $V_{y}(t)$ crossing zero.
+where $\tau_{y}$ is the first passage time for $V_{y}(t)$ crossing zero, and one can refer to Appendix A.3 for more details.
 The cohort effect is reflected by the rate parameter $\Gamma(y)$, which can be intuitively interpreted as the average initial vitality level at the time of modeling for individuals born in birth year $y$.
 The initial condition $\Gamma(0)=1$ ensures identifiability between $\Gamma(y)$ and $b(t)$.
-Note that the dynamic of $\Gamma(y)$ can be chosen to satisfy certain properties, for example, setting $\mu_{\gamma}=0$ ensures that the future cohorts are expected to be the same as the current one, or modeling $\Gamma(y)$ as a CIR process to incorporate a mean-reverting cohort effect.
-We now establish a link between this dynamic vitality model and the continuous version of the M6 model. We fix age $x$ and construct the M6 model with time $t$ and cohort $y$ as
+Note that the dynamics of $\Gamma(y)$ can be chosen to satisfy certain properties, for example, setting $\mu_{\gamma}=0$ ensures that the future cohorts are expected to be the same as the current one, or modeling $\Gamma(y)$ as a CIR process to incorporate a mean-reverting cohort effect.
+We now establish a link between this vitality mortality model and the continuous version of the M6 model. We fix age $x$ and construct the M6 model with time $t$ and cohort $y$ as
 $\ln\mu_{y}(t)$ $=\kappa_{1}(t)+\kappa_{2}(t)\cdot(x+t-\bar{x})+\gamma(y).$
 We can again re-parameterize $b(t)$, $c(t)$, and $\Gamma(y)$ in our vitality model to match the M6 model by having
-$b(t)=\exp\left(\kappa_{1}(t))-\kappa_{2}(t)\bar{x}\right),\quad c(t)=\exp\left(\kappa_{2}(t)\right),\quad\Gamma(y)=\exp(\gamma(y)).$
+$b(t)=\exp\left(\kappa_{1}(t)-\kappa_{2}(t)\bar{x}\right),\quad c(t)=\exp\left(\kappa_{2}(t)\right),\quad\Gamma(y)=\exp(\gamma(y)).$
 The $T$-year survival probability under the above re-parameterization of the vitality model will coincide with the M6 model.
-3.2.3 A dynamic vitality model with age-cohort effect
-The dynamic vitality models presented in Sections 3.2.1 and 3.2.2 begin the modeling of vitality with a fixed initial age.
-It would be interesting to formulate a vitality model that allows the vitality to start from an arbitrary age.
-We now consider a dynamic vitality model with an age-adjusted cohort effect, extended again from the Gompertz Law.
+3.2.3 A vitality model with age-adjusted period-cohort effect
+The vitality models presented in Sections 3.2.1 and 3.2.2 begin the modeling of vitality with a fixed initial age $x$.
+It would be interesting to formulate a vitality model that allows the vitality to start from an arbitrary age and the cohort effect to vary as the individual ages.
+To achieve this, we consider a vitality model with an age-adjusted cohort effect, extended again from the Gompertz Law.
 The following specification is applied:
   • C1
-    The initial vitality $V_{x,y}(0)$ for an individual born in $y$ and age $x$ at time $t=0$ follows
-    $V_{x,y}(0)\sim\text{Exp}\left(\Gamma(y)\cdot a^{x}\right),$
-    where $\Gamma(y)\cdot a^{x}$ is the rate parameter.
-We assume $a>1$ is a constant value accounting for the fact that vitality decreases as the individual ages, and $\Gamma(y)$ is governed by a stochastic process, for example, the Geometric Brownian Motion specified in Section 3.2.2.
+    The initial vitality $V_{x,y}(0)$ for an individual born in year $y$ and age $x$ at time $t=0$ follows
+    $V_{x,y}(0)\sim\text{Exp}\left(\Gamma(y)^{\frac{x_{c}-x}{x_{c}}}\right),$
+    where $x_{c}$ is a large and fixed value (e.g., $x_{c}=100$) and $\Gamma(y)$ is governed by a stochastic process (e.g., the geometric Brownian motion given by equation (10)).
   • C2
-    The trend component $Y(t)$ is specified as
-    $\mathop{}\!\mathrm{d}Y(t)=b(t)\frac{c(t)^{x+t}}{a^{x}}\mathop{}\!\mathrm{d}t,\quad t\geq 0,$
-    with the dynamics of $b(t)$ and $c(t)$ the same as in Sections 3.2.1 and 3.2.2.
-In the above specification, $\Gamma(y)$ is designed to capture the general level of initial vitality at birth for cohort year $y$, with the term $\Gamma(y)\cdot a^{x}$ reflecting the expected vitality level when the individual reaches age $x$.
-On the other hand, $b(t)$ and $c(t)$ are the Gompertz parameters that capture the change in vitality depletion rate when the individual is aging at time $t$.
-It is immediately observed that the resulting survival probability matches with the M6 model.
-It is worth noting that this specification provides more flexibility in modeling the interactions between cohorts and ages. For example, the initial vitality component can be modified to
-$V_{x,y}(0)\sim\text{Exp}\left(\Gamma(y)^{x_{c}-x}\right),$
-where $x_{c}$ is a large and fixed value, e.g., $x_{c}=100$. If $\Gamma(y)>1$, the cohort effect will diminish quickly as the individual ages, which reconciles with the cohort effect of the M8 model.
+    The trend component $Y(t)$ is specified by $\mathop{}\!\mathrm{d}Y(t)=b(t)c(t)^{x+t}\mathop{}\!\mathrm{d}t$, $t\geq 0$, where the dynamics of $b(t)$ and $c(t)$ are the same as those presented in Sections 3.2.1 and 3.2.2.
+The above specification provides flexibility in modeling the interactions between cohorts and ages.
+More specifically, $\Gamma(y)$ captures the general level of initial vitality at birth for cohort year $y$, with the term $\frac{x_{c}-x}{x_{c}}$ reflecting an age-adjusted cohort effect as the individual ages from 0 to $x_{c}$.
+When $x=0$, the initial vitality follows $\text{Exp}(\Gamma(y))$ and the specification coincides with the one presented in Section 3.2.2.
+When $x=x_{c}$, the initial vitality follows $\text{Exp}(1)$ and the specification coincides with the one presented in Section 3.2.1.
+It is worth noting that, if $\Gamma(y)>1$, the cohort effect will diminish as the individual ages, which coincides with the cohort effect observed in the M8 model.
 4 Model Applications
-The proposed vitality-based modeling framework has great potential in actuarial research.
-In this section, we provide a number of illustrative examples to showcase the applications of a vitality mortality model.
+In this section, we provide a number of illustrative examples to showcase the application of a vitality mortality model in various research or practical problems.
 Specifically, we examine the following topics:
@@
   • Analyzing the cause of death using vitality components.
-Applying the vitality mortality model differs from existing mortality models not only in terms of technical details but also by offering new interpretations and insights that cannot be explained using existing mortality models.
+We believe that the vitality mortality model differs from existing mortality models not only in terms of technical details, but also by offering new interpretations and insights that cannot be explained using existing mortality models.
 4.1 Life insurance valuation
@@
 For example, consider a vitality mortality model with initial vitality, trend, and diffusion (i.e., components C1-C3), then the problem becomes finding the first exit density of Brownian motion with a curved boundary.
-A number of studies, including Jennen and Lerche, (1981), Jennen, (1985), Skiadas, (1995), and Durbin and Williams, (1992), has provided accurate and easy-to-implement approximation methods to handle this problem.
-Skiadas and Skiadas, (2010) and Skiadas and Skiadas, (2014) have specifically applied approximation methods to vitality modeling.
-For more details on approximating the density function of the death time when C4 is absent, interested readers are referred to Appendix D.
+A number of studies, including Jennen and Lerche, (1981), Jennen, (1985), and Durbin and Williams, (1992), has provided accurate and easy-to-implement approximation methods to handle this problem.
+Skiadas and Skiadas, (2010) has specifically applied approximation methods to vitality modeling.
+For more details on approximating the density function of the death time when C4 is absent, interested readers are referred to Appendix E.
 We remark that, under the vitality modeling approach, the evaluation of life expectancy and mortality-related products can be applied to specific groups of individuals, assuming that the vitality information can be obtained or estimated.
@@
 4.2 Optimal investment and consumption strategies
-One of the potential applications of vitality modeling is lifetime planning, e.g., Merton’s portfolio problem.
+One promising application of vitality modeling is in lifetime planning, e.g., Merton’s portfolio problem.
 Individuals, when making decisions about their investment portfolio, consumption, insurance purchases, medical expenses, educational investment, etc., will take into account their health level at the time.
@@
 4.2.1 The market and vitality processes
-Following the seminal work of Merton, (1969), we construct a continuous-time model for the financial market and apply the vitality model in Section 3.1.4.
+Following the seminal work of Merton, (1969), we construct a continuous-time model for the financial market and apply the vitality model (7) without jumps.
 Denote $(\Omega,\mathbb{F},\mathbb{P})$ as a complete probability space with a right-continuous filtration generated by a two-dimensional standard Brownian Motion $(B_{S}(t),B_{V}(t))$.
@@
 $\mathop{}\!\mathrm{d}S_{0}(t)$ $=rS_{0}(t)\mathop{}\!\mathrm{d}t,\quad S_{0}(0)>0,$
-$\mathop{}\!\mathrm{d}S_{1}(t)$ $=(r+\theta\sigma)S_{1}(t)\mathop{}\!\mathrm{d}t+\sigma_{S}S_{1}(t)\mathop{}\!\mathrm{d}B_{S}(t),\quad S_{1}(0)>0,$
+$\mathop{}\!\mathrm{d}S_{1}(t)$ $=(r+\theta\sigma_{S})S_{1}(t)\mathop{}\!\mathrm{d}t+\sigma_{S}S_{1}(t)\mathop{}\!\mathrm{d}B_{S}(t),\quad S_{1}(0)>0,$
 for $t\geq 0$, where $r$ is the risk-free rate, $\theta$ is the market price of risk, and $\sigma_{S}$ is the vitality of the risky asset.
-Let $\pi(t)$ be the portfolio weight invested in the risky asset, and $\zeta(t)$ be the consumption amount. Then the asset $A(t)$ of an individual has the following dynamic
-$\mathop{}\!\mathrm{d}A(t)=A(t)\left[r+\pi(t)\theta\sigma_{S}\right]\mathop{}\!\mathrm{d}t+\sigma\pi(t)A(t)\mathop{}\!\mathrm{d}B_{S}(t)-\zeta(t)\mathop{}\!\mathrm{d}t,\quad A(0)=A_{0}.$
-We now adopt the vitality mortality model from Section 3.1.4, which recovers the Gompertz law if the initial vitality follows the Gompertz distribution; that is,
+Let $\pi(t)$ be the portfolio weight invested in the risky asset, and $\zeta(t)$ be the consumption amount. Then the asset $A(t)$ of an individual has the following dynamics
+$\mathop{}\!\mathrm{d}A(t)=A(t)\left[r+\pi(t)\theta\sigma_{S}\right]\mathop{}\!\mathrm{d}t+\sigma_{S}\pi(t)A(t)\mathop{}\!\mathrm{d}B_{S}(t)-\zeta(t)\mathop{}\!\mathrm{d}t,\quad A(0)=A_{0}.$
+We now adopt the vitality mortality model from Eq.(7) without jumps, which generalizes the Gompertz law if the initial vitality follows the Gompertz distribution; that is,
 $\mathop{}\!\mathrm{d}V(t)=-\delta\mathop{}\!\mathrm{d}t+\sigma_{V}\mathop{}\!\mathrm{d}B_{V}(t),\quad V(0)=V_{0}>0,$
@@
 $J(A(t)=a,V(t)=v)$ $=\max_{\pi,\zeta\in\Pi}\mathbb{E}_{t,a,v}\left[\int_{t}^{\tau}e^{-\beta(s-t)}\cdot u(\zeta(s))\mathop{}\!\mathrm{d}s+\lambda\cdot u(A(\tau))\right],$ (12)
-where $\tau$ is the first passage time that vitality level crosses zero, $\beta>0$ is the time-preference rate, $u(\zeta)=\ln\zeta$ is the utility function (log-utility is selected for illustration), $\lambda\geq 0$ is the weight given to the utility of terminal asset (i.e., bequest), $\mathbb{E}_{t,a,v}=\mathbb{E}[\cdot|t,A(t)=a,V(t)=v]$, and the investment/consumption strategy $(\pi=(\pi_{t})_{t\geq 0},\zeta=(\zeta_{t})_{t\geq 0})\in\Pi$ belong to the admissible set $\Pi$ if $\pi$ and $\zeta$ are progressively measurable and the resulting asset process $(A_{t})_{t\geq 0}$ is non-negative for all $t$.
+where $\tau$ is the first passage time that vitality level crosses zero, $\beta>0$ is the time-preference rate, $u(\zeta)=\ln\zeta$ is the utility function (log-utility is selected for illustration), $\lambda\geq 0$ is the weight given to the utility of terminal asset (i.e., bequest), $\mathbb{E}_{t,a,v}=\mathbb{E}[\cdot|t,A(t)=a,V(t)=v]$, and $\Pi$ is the admissible set with definition given in Appendix C.
 Notice that the optimization is time-independent since $V(t)$ is stationary and thus the value function does not depend on the time $t$.
@@
 The solution of the optimal investment strategy $\pi^{*}(t)$ and consumption strategy $\zeta^{*}(t)$ are summarized in the remark below, where the details of derivations are presented in Appendix C.
-Remark 1.
+Theorem 4.1.
 The optimal investment and consumption strategies of Problem (12) are
@@
 As for comparison, we have Merton’s optimal portfolio and consumption strategies presented in Appendix C.
-The optimal investment strategy follows the same as Merton’s optimal portfolio in an infinite time horizon, and the consumption strategy also collapsed to Merton’s case (w/o mortality) if vitality is infinite.
+The optimal investment strategy follows the same as Merton’s optimal portfolio in an infinite time horizon, and the consumption strategy also collapsed to Merton’s case (without mortality) if vitality is infinite.
 4.2.3 Further remarks
-It is not surprising to observe that the individual’s consumption depends on the vitality level, but what is interesting here is that they are not always negatively related, as one would expect.
-When the bequest motive is moderate or low (i.e., $\lambda<1/\beta$) a reasonable consumption profile is observed such that a higher vitality represents a long time to live with, and thus a lower consumption must be made to make sure that an individual will not run of out money before they are deceased.
-However, when the bequest motive is significantly large (i.e., $\lambda>1/\beta$) decreasing the vitality actually increases the consumption, a counter-intuitive result at first glance.
+It is not surprising to observe that the individual’s consumption depends on the vitality level, but what is interesting here is that they are not always negatively correlated, as one would expect.
+When the bequest motive is moderate or low (i.e., $\lambda<1/\beta$), a reasonable consumption profile is observed such that a higher vitality represents a longer time to live with, and thus a lower consumption must be made to make sure an individual will not run out of money before they are deceased.
+However, when the bequest motive is significantly large (i.e., $\lambda>1/\beta$), decreasing the vitality actually increases the consumption, a counter-intuitive result at first glance.
 This is due to the fact that individuals are overly concerned about their bequest, and thus leads to an overly conservative consumption initially.
-Once they are aging and the time for death becomes more certain, they become more aware of their over-saving and may wish to increase their consumption a bit.
+Once they are aging and the time for death becomes more certain, they become more aware of their over-saving and may wish to increase their consumption a bit, as it would not jeopardize their bequest goals but will significantly enhance their lifetime utility.
+Figure 3 illustrates the sample paths for the optimal consumption rate (as a % of wealth) and the corresponding vitality process.
+The left panel represents a low bequest motive individual ($\lambda<1/\beta$), and the right panel represents a high bequest motive individual ($\lambda>1/\beta$).
+It is evident that a higher bequest motive results in a significantly lower consumption profile and a positive correlation between the consumption rate and vitality level.
+Conversely, when the bequest motive is low, the opposite pattern is observed.
+Figure 3: Sample trajectories of vitality process (transparent lines) and optimal consumption rate (solid lines) for relatively low bequest motive (left) and high bequest motive (right).
 We emphasize that by incorporating the vitality mortality model into the optimal investment and consumption problem, new insights can be observed.
@@
 However, a 60-year-old individual who believes that they have an average initial vitality level (which is one since the initial vitality follows an exponential distribution with a mean of one) will expect to live for 20.4 more years as given by equation (4.1).
-According to Jensen’s Inequality, we can show that the average life expectancy is smaller than the life expectancy of an individual with an average vitality level:
-[…]
-[…]
-The above result suggests that even if an individual can objectively determine their health level relative to the population, it does not guarantee an unbiased estimate of their survival probabilities.
+This observation can be summarized in the following proposition.
+Proposition 2.
+Given that $V(0)\sim F_{0}$ where $F_{0}$ is a distribution function with positive support, the deterministic vitality depletion rate is strictly positive $\mu(t)>0$, and there are no diffusion and jump components, then
+$\mathbb{E}[\tau]\leq\tau(\mathbb{E}[V(0)]),$
+where $\tau(\cdot)$ is the conditional death time defined in Section 4. The average life expectancy is smaller than the life expectancy of an individual with an average vitality level.
+The proof immediately follows from the fact that the inverse function of a strictly increasing function is strictly concave, followed by Jensen’s inequality.
+For example, for the vitality counterparts of traditional mortality law displayed in Section 3.1.1, $Y(t)$ is a strictly increasing function of time, then, the death time given the initial vitality value is a concave function.
+The proposition suggests that even if an individual can objectively determine their health level relative to the population, it does not guarantee an unbiased estimate of their survival probabilities.
 However, if the individual judges their vitality level based on quantiles rather than the average (e.g., the median of the population vitality level), they will be able to arrive at an unbiased estimate of their life expectancy.
@@
 4.4.1 Permanent disability model
-Consider that there exists a threshold $\omega>0$ such that when the vitality falls below $\omega$, the person becomes disabled.
+Consider that there exists a threshold $\psi>0$ such that when the vitality falls below $\psi$, the person becomes disabled.
 Consider the specification with an exponentially distributed initial vitality (C1) and a Gompertz law specified depletion rate (C2).
 The resulting vitality model will form a permanent disability model, since $Y(t)$ is a strictly increasing function of time.
-The probability that a healthy person with initial vitality above $\omega$ will remain healthy for $T$ years is
-$\Pr\left(V(0)-Y(T)>w\big{|}V(0)>w\right)$ $=\frac{\int_{0}^{\infty}e^{-v}\cdot\mathbb{1}_{v-Y(T)>\omega}\mathop{}\!\mathrm{d}v}{\int_{0}^{\infty}e^{-v}\cdot\mathbb{1}_{v>\omega}\mathop{}\!\mathrm{d}v}$
-$=\frac{\int_{Y(T)+\omega}^{\infty}e^{-v}\mathop{}\!\mathrm{d}v}{\int_{w}^{\infty}e^{-v}\mathop{}\!\mathrm{d}v}=\exp(-Y(T)),$
+The probability that a healthy person with initial vitality above $\psi$ will remain healthy for $T$ years is
+$\Pr\left(V(0)-Y(T)>\psi\big{|}V(0)>\psi\right)$ $=\frac{\int_{0}^{\infty}e^{-v}\cdot\mathbb{1}_{v-Y(T)>\psi}\mathop{}\!\mathrm{d}v}{\int_{0}^{\infty}e^{-v}\cdot\mathbb{1}_{v>\psi}\mathop{}\!\mathrm{d}v}$
+$=\frac{\int_{Y(T)+\psi}^{\infty}e^{-v}\mathop{}\!\mathrm{d}v}{\int_{\psi}^{\infty}e^{-v}\mathop{}\!\mathrm{d}v}=\exp(-Y(T)),$
 which is equivalent to the survival probability, and the disability probability is simply $1-\exp(-Y(T))$.
-This result holds even if we assume that each individual has a different disability threshold, such that the threshold $\Omega$ has a density function $\pi(\omega)$.
+This result holds even if we assume that each individual has a different disability threshold, such that the threshold $\Psi$ has a density function $\pi(\psi)$.
 In this case, the probability of remaining healthy for $T$ years is
-$\frac{\int_{0}^{\infty}e^{-v}\cdot\int_{0}^{\infty}\pi(\omega)\cdot\mathbb{1}_{v-\omega-Y(T)>0}\mathop{}\!\mathrm{d}\omega\mathop{}\!\mathrm{d}v}{\int_{0}^{\infty}e^{-v}\cdot\int_{0}^{\infty}\pi(\omega)\cdot\mathbb{1}_{v-\omega>0}\mathop{}\!\mathrm{d}\omega\mathop{}\!\mathrm{d}v}$ $=\frac{\int_{0}^{\infty}\pi(\omega)e^{-(\omega+Y(T))}\mathop{}\!\mathrm{d}\omega}{\int_{0}^{\infty}\pi(\omega)\cdot e^{-\omega}\mathop{}\!\mathrm{d}\omega}=\exp(-Y(T)).$
+$\frac{\int_{0}^{\infty}e^{-v}\cdot\int_{0}^{\infty}\pi(\psi)\cdot\mathbb{1}_{v-\psi-Y(T)>0}\mathop{}\!\mathrm{d}\psi\mathop{}\!\mathrm{d}v}{\int_{0}^{\infty}e^{-v}\cdot\int_{0}^{\infty}\pi(\psi)\cdot\mathbb{1}_{v-\psi>0}\mathop{}\!\mathrm{d}\psi\mathop{}\!\mathrm{d}v}$ $=\frac{\int_{0}^{\infty}\pi(\psi)e^{-(\psi+Y(T))}\mathop{}\!\mathrm{d}\psi}{\int_{0}^{\infty}\pi(\psi)\cdot e^{-\psi}\mathop{}\!\mathrm{d}\psi}=\exp(-Y(T)).$
 The above setting of the disability threshold $\omega$ can be extended in other directions, such as considering an age- and time-dependent threshold. We leave this for future research.
@@
 However, when the diffusion component (C3) is included, the model becomes capable of describing disability with the potential for recovery.
-For illustrative purposes, we consider the alternative Gompertz specification from Section 3.1.4 with a diffusion component (C3) modeled as a Brownian motion multiplied by a constant volatility $\sigma$, and set a constant threshold $\omega$ for disability.
-The $T$-year recovery probability for a disabled person aged $x$ is
-$\frac{\int_{0}^{\omega}\Pr\left(v-Y(T)-W(T)>\omega,\text{ and }v-Y(t)-W(t)>0\text{ for any }t\in[0,T]\right)\mathop{}\!\mathrm{d}F_{0}(v)}{1-F_{0}(\omega)}.$
-Since $Y(t)=\delta t$, the probability inside in the integral can be solved explicitly by setting $\widehat{W}(t)=W(t)+\frac{\delta}{\sigma}t$, which is a Brownian Motion with drift $\frac{\delta}{\sigma}$.
-Denote
-$\widehat{M}(T)=\max_{0\leq t\leq T}\{W(t)+Y(t)\},$
-the joint density of $(\widehat{M}(T),\widehat{W}(T))$ can be found as
-$f(m,w)=\frac{2(2m-w)}{T\sqrt{2\pi T}}e^{\frac{\delta w}{\sigma}-\frac{\delta^{2}T}{2\sigma^{2}}-\frac{(2m-w)^{2}}{2T}},w\leq m,m\geq 0,$
-and is zero otherwise (see Shreve et al., (2004) for a detailed derivation). Consequently, the probability inside the integral becomes
-$\Pr\left(\widehat{M}(T)<\frac{v}{\sigma},\quad\widehat{W}(T)<\frac{v-\omega}{\sigma}\right)=\int_{0}^{v}\int_{0}^{v-\omega}f(m,w)\mathop{}\!\mathrm{d}m\mathop{}\!\mathrm{d}w.$
-The recovery probability can then be determined by numerically solving the triple integral.
+For illustrative purposes, here we consider the simplest model that is able to capture the recovery probability, specifically, we consider the alternative Gompertz specification with a constant depletion rate and a diffusion component (C3) modeled as a Brownian motion multiplied by a constant volatility, along with a constant disability threshold.
+The result is summarized in the following proposition.
+Proposition 3.
+Consider the vitality model with an initial vitality level $V(0)\sim F_{0}$, a constant vitality depletion rate $\mu(t)=\delta$, and a diffusion component as a Brownian motion with volatility $\sigma$.
+Consider a constant disability threshold $\psi$, then the $T$-year disabled probability for a healthy person aged $x$ is
+$\int_{\psi}^{\infty}\frac{1}{1-F_{0}(\psi)}\cdot\left(\int_{0}^{\frac{v}{\sigma}}\int_{\frac{v-\psi}{\sigma}}^{\frac{v}{\sigma}}f(m,w;T)\mathop{}\!\mathrm{d}w\mathop{}\!\mathrm{d}m\right)\mathop{}\!\mathrm{d}F_{0}(v),$
+and the recovery probability for a disabled person aged $x$ is
+$\int_{0}^{\psi}\frac{1}{F_{0}(\psi)}\cdot\left(\int_{0}^{\frac{v}{\sigma}}\int_{0}^{\frac{v-\psi}{\sigma}}f(m,w;T)\mathop{}\!\mathrm{d}w\mathop{}\!\mathrm{d}m\right)\mathop{}\!\mathrm{d}F_{0}(v),$
+where
+$f(m,w;T)=\begin{cases}\frac{2(2m-w)}{T\sqrt{2\pi T}}e^{\frac{\delta w}{\sigma}-\frac{\delta^{2}T}{2\sigma^{2}}-\frac{(2m-w)^{2}}{2T}},&w\leq m,m\geq 0\\ \\ 0,&\text{elsewhere}.\end{cases}$
+The proof is given in Appendix D.
+The probabilities can be evaluated by numerically solving the triple-integral.
+Notice that the choice of the disability threshold here does not impact the death probability, suggesting an approach to seamlessly integrate disability modeling and mortality modeling, which contrasts with traditional multi-state modeling for disability.
+Of course, the model above might be overly simplified; further exploration in model extension and empirical examination with disability data would be necessary.
 4.5 Cause-of-death modeling
@@
 For $q\geq 0$, the Laplace transforms of the time of death due to an accident and the time of death due to natural decay in vitality level are given by the following expressions:
-$\mathbb{E}_{v}\left[\mathrm{e}^{-q\tau_{J}}\mathbf{1}\left(\tau_{J}<\tau_{Y}\right)\right]=$ $\int_{0}^{t_{v}^{*}}\lambda\mathrm{e}^{-(q+\lambda)t-\alpha b^{*}_{v}}\cdot I_{0}\left(2\sqrt{\alpha\lambda tb^{*}_{v}}\right)\mathrm{d}t,$ (13)
+$\mathbb{E}_{v}\left[\mathrm{e}^{-q\tau_{J}}\mathbb{1}\left(\tau_{J}<\tau_{Y}\right)\right]=$ $\int_{0}^{t_{v}^{*}}\lambda\mathrm{e}^{-(q+\lambda)t-\alpha b^{*}_{v}(t)}\cdot I_{0}\left(2\sqrt{\alpha\lambda tb^{*}_{v}(t)}\right)\mathrm{d}t,$ (13)
 and
-$\mathbb{E}_{v}\left[\mathrm{e}^{-q\tau_{Y}}\mathbf{1}\left(\tau_{Y}<\tau_{J}\right)\right]=$ $\int_{0}^{t^{*}_{v}}\mathrm{e}^{-\left(q+\lambda\right)t-\alpha b^{*}_{v}}\cdot bc^{x+t}\sqrt{\frac{\alpha\lambda t}{b^{*}_{v}}}I_{1}\left(2\sqrt{\alpha\lambda tb^{*}_{v}}\right)\mathrm{d}t+\mathrm{e}^{-(q+\lambda)t^{*}_{v}},$ (14)
-where $t^{*}_{v}=\frac{\ln(bc^{x}+v\ln c)-\ln b}{\ln c}-x$ (which refers to the remaining lifetime without any accidents/jumps), $b^{*}_{v}=v-bc^{x}\left(c^{t}-1\right)/\ln c$, and
+$\mathbb{E}_{v}\left[\mathrm{e}^{-q\tau_{Y}}\mathbb{1}\left(\tau_{Y}<\tau_{J}\right)\right]=$ $\int_{0}^{t^{*}_{v}}\mathrm{e}^{-\left(q+\lambda\right)t-\alpha b^{*}_{v}(t)}\cdot bc^{x+t}\sqrt{\frac{\alpha\lambda t}{b^{*}_{v}(t)}}I_{1}\left(2\sqrt{\alpha\lambda tb^{*}_{v}(t)}\right)\mathrm{d}t+\mathrm{e}^{-(q+\lambda)t^{*}_{v}},$ (14)
+where $t^{*}_{v}=\frac{\ln(bc^{x}+v\ln c)-\ln b}{\ln c}-x$ (which refers to the remaining lifetime without any accidents/jumps), $b^{*}_{v}(t)=v-bc^{x}\left(c^{t}-1\right)/\ln c$, and
 $I_{k}(x):=\sum_{n=0}^{\infty}\frac{(x/2)^{k+2n}}{n!\Gamma(k+n+1)}$
 is the modified Bessel function of the first kind. Moreover, by the uniqueness of the Laplace transform, it follows that
-$\mathrm{Pr}_{v}\left(\tau_{J}\in\mathrm{d}t,\tau_{J}<\tau_{Y}\right)$ $=\mathbb{1}(t<t^{*})\lambda\mathrm{e}^{-\alpha b^{*}_{v}-\lambda t}I_{0}\left(2\sqrt{\lambda\alpha tb^{*}_{v}}\right)\mathrm{d}t,$
+$\mathrm{Pr}_{v}\left(\tau_{J}\in\mathrm{d}t,\tau_{J}<\tau_{Y}\right)$ $=\mathbb{1}(t<t^{*})\lambda\mathrm{e}^{-\alpha b^{*}_{v}(t)-\lambda t}I_{0}\left(2\sqrt{\lambda\alpha tb^{*}_{v}(t)}\right)\mathrm{d}t,$
 and
-$\mathrm{Pr}_{v}\left(\tau_{Y}\in\mathrm{d}t,\tau_{Y}<\tau_{J}\right)=$ $\mathbb{1}(t<t^{*})\mathrm{e}^{-\lambda t-\alpha b^{*}_{v}}\cdot bc^{x+t}\sqrt{\frac{\alpha\lambda t}{b^{*}_{v}}}I_{1}\left(2\sqrt{\alpha\lambda tb^{*}_{v}}\right)\mathrm{d}t+\mathrm{e}^{-\lambda t}\delta_{t^{*}}(\mathrm{d}t),$
+$\mathrm{Pr}_{v}\left(\tau_{Y}\in\mathrm{d}t,\tau_{Y}<\tau_{J}\right)=$ $\mathbb{1}(t<t^{*})\mathrm{e}^{-\lambda t-\alpha b^{*}_{v}(t)}\cdot bc^{x+t}\sqrt{\frac{\alpha\lambda t}{b^{*}_{v}(t)}}I_{1}\left(2\sqrt{\alpha\lambda tb^{*}_{v}(t)}\right)\mathrm{d}t+\mathrm{e}^{-\lambda t}\delta_{t^{*}}(\mathrm{d}t),$
 where $\delta_{t_{v}^{*}}(\cdot)$ is the Dirac mass at $t_{v}^{*}$.
@@
 5 A Numerical Illustration
-In this section, we provide an illustrative example to discuss the numerical estimation procedure of the proposed vitality-based mortality modeling framework.
-We consider the jump-diffusion specification shown in Section 3.1.3.
-Recall from equations (4) and (5) that the dynamics of vitality is governed by
-$V(t)=V(0)-\int_{0}^{t}\mu_{x}(s)\mathop{}\!\mathrm{d}s-\sigma B(t)-\sum_{i=1}^{N(t)}Z_{i},$
-and the resulting $T$-year survival probability is
-$\Pr(\tau>T)=\int_{0}^{\infty}\mathbb{P}\left(v-\sigma B(t)-\sum_{i=1}^{N(t)}Z_{i}\geq\int_{0}^{t}\mu_{x}(s)\mathop{}\!\mathrm{d}s,\;\forall t\leq T\right)\mathop{}\!\mathrm{d}F_{0}(v).$
-5.1 Estimation simplifications
-The computational requirements for evaluating the conditional probability in the equation of $\Pr(\tau>T)$ are quite intensive.
-For ease of computation, we make the following assumptions to achieve an approximation:
+In this section, we provide an illustrative example on the numerical estimation procedure of the proposed framework.
+We consider the jump-diffusion vitality model discussed in Section 3.1.3.
+Recall that the vitality process $V(t)$ is described by equation (4) while the resulting survival probability $\Pr(\tau>T)$ is given in equation (5).
+5.1 Simplifications
+The computational requirements for evaluating equation (5) are quite intensive. For ease of computation, we make the following assumptions to achieve an approximation:
   • The force of mortality is constant between integer ages.
@@
 Under the constant force of mortality assumption, $Y(t)$ can be written as a piece-wise linear function:
-$Y(t)=a_{i}(t-\lfloor t\rfloor)+b_{i},$
-for $i\leq t<i+1$ and $i\in\{0,1,2,\ldots\}$, where
-$b_{i}$ $=\int_{0}^{i}\mu_{x}(s)\mathop{}\!\mathrm{d}s\quad\text{and}\quad a_{i}=\int_{i}^{i+1}\mu_{x}(s)\mathop{}\!\mathrm{d}s.$
-Using the well-known formula for the non-crossing probability for a linear boundary $at+b$ before time $s$, that is
-$\Pr(B(t)<at+b\text{ for all }t\leq s\big{|}B(s)=x)=1-\exp\left\{-\frac{2b(as+b-x)}{s}\right\},$
-together with the fatal jump assumption, the conditional survival probability is given by
-${}_{k}p_{x}(v)=\mathbb{E}\left[\left(\prod_{s=1}^{k}\mathbb{1}_{B(s)<\frac{v-Y(s)}{\sigma}}\cdot\left(1-e^{-\frac{2(v-Y(s-1)-\sigma B(s-1))(v-Y(s)-\sigma B(s))}{\sigma^{2}}}\right)\right)\cdot\exp\left(-\int_{0}^{k}\lambda(s)\mathop{}\!\mathrm{d}s\right)\right],k\in\mathbb{N}^{+},$
-for a given initial vitality $v>0$. It follows that the $T$-year survival probability becomes
+$Y(t)=\alpha_{i}(t-\lfloor t\rfloor)+\beta_{i},$
+for $i\leq t<i+1$ and $i\in\{0,1,2,\ldots\}$, where $\beta_{i}=\int_{0}^{i}\mu_{x}(s)\mathop{}\!\mathrm{d}s$ and $\alpha_{i}=\int_{i}^{i+1}\mu_{x}(s)\mathop{}\!\mathrm{d}s$.
+Using the well-known formula for the non-crossing probability of a linear boundary $\alpha t+\beta$ before time $s$; that is,
+$\Pr\left(B(t)<\alpha t+\beta\text{ for all }t\leq s\big{|}B(s)=x\right)=1-\exp\left\{-\frac{2\beta(\alpha s+\beta-x)}{s}\right\},$
+together with the fatal jump assumption, the survival probability conditional on an initial vitality $v>0$ is given by
+${}_{k}p_{x}(v)=\mathbb{E}\left[\left(\prod_{s=1}^{k}\mathbb{1}_{B(s)<\frac{v-Y(s)}{\sigma}}\left(1-e^{-\frac{2(v-Y(s-1)-\sigma B(s-1))(v-Y(s)-\sigma B(s))}{\sigma^{2}}}\right)\right)\exp\left(-\int_{0}^{k}\lambda(s)\mathop{}\!\mathrm{d}s\right)\right],$
+for $k\in\mathbb{N}^{+}$. It follows that the $T$-year survival probability becomes
 $\Pr(\tau>T)=\int_{0}^{\infty}{}_{T}p_{x}(v)\mathop{}\!\mathrm{d}F_{0}(v).$
-5.2 Maximum likelihood estimation
-In this subsection, we consider the maximum likelihood estimation method for estimating the parameters of the vitality model.
-We begin by describing the likelihood function for a cohort mortality dataset.
+5.2 Model estimation
+Given the complexity of vitality mortality models, using solely the mortality data from a national population is inadequate for estimation purposes.
+Specifically, such dataset does not distinguish between deaths caused by natural aging and those resulting from accidents.
+This limitation makes it challenging to differentiate between accidental deaths, captured by the jump component, and natural deaths, represented by the initial and trend components.
+In terms of parameter estimation, the intensity rate $\lambda(t)$ of the Poisson process $N(t)$ cannot be easily estimated without information that explicitly categorizes accidental deaths.
+To address this issue, we utilize mortality data that are specifically related to accidental causes of death.
+In particular, we use the WONDER dataset from the Centers for Disease Control and Prevention (CDC) to determine the intensity rate $\lambda(t)$.
+The remaining parameters, $\theta_{-\lambda}=\theta\backslash\{\lambda(t)\}_{t=0,1,\cdots}$, are then estimated using the maximum likelihood estimation method.
+This approach allows us to separate the jump component from the rest of the components.
+More specifically, we estimate $\theta_{-\lambda}$ using the U.S.
+male mortality data from the Human Mortality Database (HMD) [^1] Human Mortality Database.
+Max Planck Institute for Demographic Research (Germany), University of California, Berkeley (USA), and French Institute for Demographic Studies (France).
+Available at www.mortality.org..
+We focus on the cohort born in 1910, with age 60 and above during the period of 1970 to 2016.
+We estimate $\lambda(t)$ using the average accidental death rates for age groups in ten-year intervals during the period of 1968 to 2016 from the CDC WONDER dataset.
+The likelihood function of our maximum likelihood estimation is given as follows.
 Consider a cohort of individuals all aged $x$ at time $t=0$.
-Let $E_{x}$ be the initial exposure at time 00, and let $D_{x+t}$ be the observed death counts at age $x+t$ for $t\in\{0,1,2,\ldots\}$.
-We assume that $D_{x+t}$ follows a multinomial distribution, and thus the log-likelihood function of the cohort mortality data is
+Let $E_{x}$ be the initial exposure at time 00 and $D_{x+t}$ be the observed death counts at age $x+t$ for $t\in\mathbb{N}^{+}$.
+We assume that $D_{x+t}$ follows a multinomial distribution, which leads to the log-likelihood function as
 $l(\theta)=\sum_{t\geq 0}D_{x+t}\cdot\ln\left[\Pr(\tau>t)-\Pr(\tau>t+1)\right]+\ln E_{x}!-\sum_{t\geq 0}\ln D_{x+t}!,$
@@
 Note that our estimation procedure differs from Shimizu et al., (2021) and Shimizu et al., (2023), who used the least squares method to estimate the parameters of their survival energy models.
-Given the complexity of our vitality model, relying solely on a cohort mortality dataset from a national population is inadequate for estimation purposes.
-Specifically, a national cohort dataset does not distinguish between deaths caused by natural aging and those resulting from accidents.
-This limitation makes it challenging to differentiate between accidental deaths, captured by the jump component, and natural deaths, represented by the initial and depletion components.
-In terms of parameter estimation, the intensity rate $\lambda(t)$ of the Poisson process $N(t)$ cannot be easily estimated without information that explicitly categorizes accidental deaths.
-To address this issue, we utilize mortality data specifically related to external causes of death.
-In particular, we use external death counts from the Centers for Disease Control and Prevention (CDC) of the United States to estimate the intensity rate $\lambda(t)$.
-The remaining parameters, $\theta_{-\lambda}=\theta\backslash\{\lambda(t)\}_{t=0,1,\cdots}$, are then estimated using the maximum likelihood estimation method.
-This approach allows us to accurately separate the jump component from the rest of the components.
-The numerical results of our estimation procedures are presented in the next subsection.
-5.3 Numerical estimation results
-To estimate the vitality model, we utilize the U.S.
-male mortality data from the Human Mortality Database (HMD) [^1] Human Mortality Database.
-Max Planck Institute for Demographic Research (Germany), University of California, Berkeley (USA), and French Institute for Demographic Studies (France).
-Available at www.mortality.org..
-We focus on the cohort born in 1910, with ages 60 and above during the period from 1970 to 2016.
-Additionally, we employ data from the Centers for Disease Control and Prevention (CDC) that consists of the average accidental death rates for age groups in ten-year intervals, spanning from 1968 to 2016.
-The jump intensity $\lambda(t)$ is directly determined as a constant within each age group based on the CDC data.
-The remaining parameters are fitted to the HMD data described above.
-Recall that the jump-diffusion vitality model from Section 3.1.3 uses the Gompertz law to specify the trend component; that is, $\mu_{x}(t)=bc^{x+t}$.
-It follows that the vitality dynamics are governed by
+5.3 Numerical results
+Using the Gompertz law to specify the trend component of the vitality dynamics, the jump-diffusion vitality model is specified as
 $V(t)=V(0)-\int_{0}^{t}bc^{x+s}\mathop{}\!\mathrm{d}s-\sigma B(t)-\sum_{i=1}^{N(t)}Z_{i},$
 where $V(0)\sim\text{Exp}(1)$.
-Given that the intensity rate of $N(t)$ has been externally estimated and we assumed $Z_{i}=\infty$ for all $i$, the remaining parameters to be estimated are $b$, $c$, and $\sigma$.
+Given that the intensity rate of $N(t)$ has been externally determined and we assumed $Z_{i}=\infty$ for all $i$, the remaining parameters to be estimated are $b$, $c$, and $\sigma$.
 In addition to the vitality model, we also estimate the parameters of the Gompertz law for comparison purposes.
-Table 1 reports the estimated values of $b$, $c$, and $\sigma$ for the vitality model and the estimated values of $b$ and $c$ for the Gompertz law.
-The maximized log-likelihood of each model is also reported in the last row.
-We observe that the estimated values of $b$ and $c$ are similar between the vitality model and the Gompertz law, indicating that both models accurately describe the underlying mortality curve via the assumed functional form, $\mu_{x}(t)=bc^{x+t}$.
-In addition to the Gompertz law, the vitality model has an estimated value of 0.0036 for $\sigma$, which reflects the random fluctuations in vitality over the lifetime of the cohort.
-Overall, the vitality model achieves a higher maximized log-likelihood of -2,181 compared to the Gompertz law’s -2,591.
+Table 1 reports the estimated values of $b$, $c$, and $\sigma$ from the vitality model and the Gompertz law for both males (left panel) and females (right panel).
+The maximized log-likelihood of each model is reported at the last row of the corresponding table.
+We observe that the estimated values of $b$ and $c$ are similar between the vitality model and the Gompertz law, indicating that both models accurately describe the underlying mortality curve via their assumed functional form (i.e., $\mu_{x}(t)=bc^{x+t}$).
+Compared to the Gompertz law, the vitality model additionally has an estimated value of $\sigma$ at 0.0098 and 0.0034 for males and females, respectively.
+These values reflect the random fluctuations in vitality over the lifetime of the corresponding cohort of males and females.
+Overall, the vitality model achieved a higher maximized log-likelihood than the Gompertz law for both genders.
 Model Vitality Gompertz
-$b$ 0.00015391 0.0001744
-$c$ 1.0834 1.082
-$\sigma$ 0.0036
-$l(\theta)$2,181 -2,591
-Table 1: Parameter estimates and maximized log-likelihood for the vitality model and the Gompertz law.
-Figure 2 shows the fitted values of death rates and survival probabilities for ages 60-100 and periods 0-40 years under the vitality model and the Gompertz law.
-The observed values from the empirical data are also plotted for comparison.
-We observe that both the vitality model and the Gompertz law are able to depict a well-fitted mortality curve for the cohort considered and also produce reasonable survival probabilities at age 60 for the next 40 years.
-Comparing the two models, the vitality model appears to better fit the cohort’s mortality data and produce a mortality curve that is closer to the observed values.
-Figure 2: Fitted log death rates $\ln q_{x}$ for $x\in(60,100)$ and survival probabilities ${}_{t}p_{60}$ for $t\in(0,40)$ under the vitality model (blue solid line) and the Gompertz law (red dash line).
-The observed values are depicted by dots.
+$b$ $1.4819\times 10^{-4}$ $1.7440\times 10^{-4}$
+$c$ 1.0838 1.0820
+$\sigma$ $9.8267\times 10^{-3}$
+$l(\theta)$2,057 -2,591
+Model Vitality Gompertz
+$b$ $1.8107\times 10^{-5}$ $1.9915\times 10^{-5}$
+$c$ 1.1058 1.1048
+$\sigma$ $3.4139\times 10^{-3}$
+$l(\theta)$3,375 -3,901
+Table 1: Parameter estimates and maximized log-likelihood of the vitality model and the Gompertz law for males (left) and females (right).
 6 Conclusion
 In this paper, we introduced a vitality-based modeling framework that offers a new perspective on mortality modeling.
-The proposed framework integrates multiple components to capture the complexities of vitality dynamics over an individual’s lifetime.
-By incorporating the initial vitality level, natural depletion trends, stochastic diffusion processes, and sudden jumps, our framework provides a comprehensive and flexible approach to mortality modeling.
+The proposed framework integrates four unique components to capture the complexities of vitality dynamics over an individual’s life course.
 The adaptability of the four components allows for the incorporation of various factors affecting mortality, providing deeper insights into population mortality trends and individual health trajectories.
-Through both static and dynamic specifications of the proposed framework, we demonstrated its versatility in addressing various mortality modeling scenarios.
+Through both static and dynamic parameter specifications of the proposed framework, we demonstrated its versatility in addressing various mortality modeling scenarios.
 Our framework not only replicates traditional mortality laws but also extends to more advanced stochastic mortality models, offering a unified structure for understanding mortality dynamics.
 An illustrative example is provided to showcase how a vitality model can be fitted using the maximum likelihood estimation method.
-Additionally, the framework’s modularity ensures that it can be tailored to specific research needs, making it a valuable tool for both theoretical exploration and practical applications.
-Our findings suggest that the vitality-based modeling framework offers significant practical applications, such as life insurance valuation, disability modeling, cause-of-death modeling, explaining subjective beliefs about mortality, and deriving optimal consumption solutions.
-By capturing the initial level, natural aging, inherent randomness, and sudden changes in an individual’s vitality, the framework supports a better interpretation of mortality dynamics in the aforementioned applications.
-Lastly, we emphasize that the proposed framework can be further utilized in more nuanced scenarios to improve results and decision-making in various fields.
-While we used four components to develop a comprehensive modeling framework for vitality dynamics, the specific structures considered for each component are not exhaustive in this paper.
+Our investigation suggests that the vitality-based modeling framework offers significant practical applications, such as life insurance valuation, disability modeling, cause-of-death modeling, explaining subjective beliefs on mortality, and deriving optimal consumption solutions.
+We emphasize that the proposed framework can be further utilized in more nuanced scenarios to improve results and decision-making in various fields.
+While we used four components in the proposed framework, the specific structures considered for each component are not exhaustive in this paper.
 There are numerous possible extensions for articulating the four components.
 For instance, one could consider a multivariate stochastic process to govern the diffusion component, reflecting the fact that multiple underlying forces may drive the randomness in an individual’s vitality.
-The initial component could be modified to measure and explain healthy life expectancy.
-The jump component can also be expanded to consider jumps categorized by different causes of death, such as natural disasters, car accidents, heart attacks, and suicide.
-We have demonstrated in this paper that the proposed modeling framework is capable of replicating mortality laws and stochastic mortality models.
-Beyond these existing models, further exploration can be conducted to show that continuous-time intensity-based models, such as the Vasicek process applied to the force of mortality, can also be mimicked by a vitality model.
-It would be interesting to understand how the first passage time problem connects with an continuous-time mortality model to produce explicit solutions for survival probabilities.
+The jump component can also be expanded to consider jumps categorized by different causes of death.
+We have demonstrated in this paper that the proposed framework is capable of replicating mortality laws and factor-based models.
+Further exploration can be conducted to show that intensity-based models, such as the Vasicek process applied to the force of mortality, can also be mimicked by the proposed framework.
+It would be interesting to understand how the first passage time problem connects with an continuous-time intensity-based model.
 This connection could potentially provide new insights and more accurate survival probability estimates.
-Although we have explored various applications of the vitality-based modeling framework, there are several other potential applications that we did not focus on.
-For example, a dynamic vitality model can be used to manage longevity risk by developing dynamic hedging strategies and pricing longevity-linked securities.
+There are several other potential applications that we did not consider in this paper.
+For instance, a vitality model with dynamic parameters can be used to manage longevity risk by developing hedging strategies and pricing longevity-linked securities.
 Additionally, since our framework models individual vitality, it can be employed to measure the basis risk between an individual insurance product and an index-based security.
 These extensions would provide further advancements to capital market solutions for longevity risk management.
-Lastly, we acknowledge that the vitality dynamics of an individual cannot be directly observed, or at least not easily.
-To accurately estimate a vitality model using data beyond the national population data presented in this paper, one would need to use personal data over time or clinical data on vital measures.
-Potential sources of such data could include personal medical devices, such as smartwatches or mobile health trackers, which continuously monitor health metrics like heart rate, physical activity, and sleep patterns.
-Integrating these diverse data sources would allow for a more precise estimation of vitality dynamics.
+Lastly, we acknowledge that the vitality dynamics of an individual cannot be directly observed and easily estimated.
+To improve the estimation of a vitality model, one would need to use personal health data or clinical data on vital measures.
+Potential sources of such data could include personal medical devices, such as smartwatches or mobile health trackers.
 We remark that this research direction is highly related to InsurTech and could inspire future studies on personalized insurance products.
 Acknowledgement
-We extend our gratitude for the helpful and constructive comments received at the School of Actuarial Science at Université Laval, the 27th International Congress on Insurance: Mathematics and Economics, and the 59th Actuarial Research Conference.
 Xiaobai Zhu acknowledges the support from the Research Grants Council of the Hong Kong Special Administrative Region, China (CUHK 24615523), and National Natural Science Foundation of China (No.
 12301613).
-Zijia Wang acknowledges the support from a start-up grant from The Chinese University of Hong Kong.
-Appendix A Proofs
-A.1 Death Plateau Models
+Zijia Wang acknowledges the support from internal grants from The Chinese University of Hong Kong and National Natural Science Foundation of China (No.
+12401623).
+Appendix A Technical Details
+A.1 Plateau Death Models
 Gamma-Gompertz Model
@@
 $=\int_{0}^{\infty}\exp\left(-z\cdot\int_{0}^{T}\mu_{x}(s)\mathop{}\!\mathrm{d}s\right)\cdot\frac{\phi^{\alpha}\cdot z^{\alpha-1}\cdot e^{-z\cdot\phi}}{\Gamma(\alpha)}\mathop{}\!\mathrm{d}z.$
-The remaining steps follow exactly the same as before, and we will end up with the same survival function as in Appendix A.1.
+The remaining steps follow exactly the same as before, and we will end up with the same survival function as in equation (3).
 Dagum Decay Rate
@@
 $=\left(1+\left(\frac{v}{b\cdot\left(\int_{0}^{T}\mu_{x}(s)\mathop{}\!\mathrm{d}s\right)}\right)^{-a}\right)^{-p}.$
-If $a=1$, $p=\alpha$, and $b=\frac{v}{\phi}$, then we recover the same survival function as in Appendix A.1.
+If $a=1$, $p=\alpha$, and $b=\frac{v}{\phi}$, then we recover the same survival function as in equation (3).
 A.2 Alternative Expression of Mortality Laws
@@
 Then, by either analytically or numerically invert the above Laplace transform with respect to $q$, one can obtain the distribution function of $\tau$.
-A.3 Dynamic Model
+A.3 Vitality Models with Dynamic Parameters
 When $Y(t)$ is non-decreasing, for example, the integral of correlated log-normally distributed random variables as in Section 3.2.2, then
@@
 C.3 Merton’s Problem with Vitality
+Definition C.1.
+(Admissible Strategies) The pair of strategies $(\pi(\cdot),\zeta(\cdot))$ is called an admissible strategy, if it satisfies the following conditions:
+  1. 1.
+    $\pi(\cdot)$ and $\zeta(\cdot)$ are progressively measurable;
+  2. 2.
+    $\mathbb{E}_{0,a,v}\left[\int_{0}^{t}\pi^{2}(s)+\zeta^{2}(s)\mathop{}\!\mathrm{d}s\right]<\infty$ and $A(t)>0$ for any $t\in[0,\infty)$ and $v>0$.
+  3. 3.
+    associated with $(\pi(\cdot),\zeta(\cdot))$, the state equation $A(\cdot)$ has a unique strong solution.
+We denote the set of all admissible strategies by $\Pi$.
 The associated HJB equation for the value function $J(a,v)$ is
@@
 The expression of $f(t)$ has a similar interpretation with $F(t)$ in Section C.2 but based on the remaining life $t^{*}$ ($e^{-\beta t^{*}}$ is the present value of \$1 pays at time $t^{*}$).
-Appendix D Approximation Methods
-D.1 Tangent Approximation
-If we set $J(t)=0$, and given the initial vitality level $V(0)=v$, then Jennen and Lerche, (1981); Skiadas and Skiadas, (2014) discussed a tangent approximation method to get the density of $\tau(v)$; that is,
+Appendix D Disability with Recovery
+Proof.
+The $T$-year recovery probability for a disabled person aged $x$ can be expressed as
+[…]
+$=$ $\int_{0}^{\psi}\frac{\Pr\left(\left\{v-Y(T)-W(T)>\psi\right\}\text{ and }\left\{v-Y(t)-W(t)>0\text{ for any }t\in[0,T]\right\}\right)}{F_{0}(\psi)}\mathop{}\!\mathrm{d}F_{0}(v).$ (15)
+By assumption, the trend component is $Y(t)=\delta\cdot t$, and we may define
+$\widehat{W}(t)$ $=\frac{1}{\sigma}\cdot\left(W(t)+Y(t)\right)=\frac{1}{\sigma}\cdot\left(W(t)+\delta\cdot t\right),$
+$\widehat{M}(T)$ $=\frac{1}{\sigma}\cdot\max_{0\leq t\leq T}\left\{W(t)+Y(t)\right\}=\max_{0\leq t\leq T}\left\{\widehat{W}(t)\right\},$
+where $\widehat{W}(t)$ is a Brownian motion with drift $\delta/\sigma$, and $\sigma\cdot\widehat{M}(t)$ tracks the maximum vitality depletion until $t$.
+If $\sigma\cdot\widehat{M}(t)$ is larger than the initial vitality level, then the individual is no longer alive at time $t$.
+Denote $f(m,w;T)$ as the joint density of $(\widehat{M}(T),\widehat{W}(T))$.
+Using results from Shreve et al., (2004), one can derive
+$f(m,w;T)=\frac{2(2m-w)}{T\sqrt{2\pi T}}e^{\frac{\delta w}{\sigma}-\frac{\delta^{2}T}{2\sigma^{2}}-\frac{(2m-w)^{2}}{2T}},w\leq m,m\geq 0,$
+and is zero otherwise. Notice that the numerator of the expression inside equation (15) can be written as a function of $(\widehat{M}(T),\widehat{W}(T))$:
+$\Pr\left(\left\{v-Y(T)-W(T)>\psi\right\}\text{ and }\left\{v-Y(t)-W(t)>0\text{ for any }t\in[0,T]\right\}\right)$
+$=$ $\Pr\left(v-\sigma\cdot\widehat{W}(T)>\psi,v-\sigma\cdot\widehat{M}(t)>0\right)$
+$=$ $\Pr\left(\widehat{W}(T)<\frac{v-\psi}{\sigma},\ \widehat{M}(T)<\frac{v}{\sigma}\right).$
+Therefore, substitute the density function $f(m,v;T)$, we have the recovery probability as:
+$\int_{0}^{\psi}\frac{1}{F_{0}(\psi)}\cdot\left(\int_{0}^{\frac{v}{\sigma}}\int_{0}^{\frac{v-\psi}{\sigma}}f(m,w;T)\mathop{}\!\mathrm{d}w\mathop{}\!\mathrm{d}m\right)\mathop{}\!\mathrm{d}F_{0}(v).$
+A similarly procedure follows for the disability probability and therefore we omit the details.
+Appendix E Approximation Methods
+E.1 Tangent Approximation
+If we set $J(t)=0$, and given the initial vitality level $V(0)=v$, then Jennen and Lerche, (1981) discussed a tangent approximation method to get the density of $\tau(v)$; that is,
 $\Pr(\tau(v)\in\mathop{}\!\mathrm{d}t)=\frac{|H(t,v)-t\cdot H_{t}(t,v)|}{\sigma\sqrt{2\pi t^{3}}}e^{-\frac{\left(H(t,v)\right)^{2}}{2\sigma^{2}t}},$
 where $H(t,v):=v-Y(t)$.
-D.2 Series Expansion
+E.2 Series Expansion
 The tangent approximation has the advantage of simplicity but may sacrifice the accuracy of the approximation.
@@
 Durbin and Williams, (1992) demonstrated the accuracy of this series expansion using specific concave functions for $H$, and found that with a choice of $k=3$ (which leads to a double integration), the error lies within $0.001\%$ of the true value.
-Appendix E Cause-of-Death Analysis
+Appendix F Cause-of-Death Analysis
 Proof of Theorem 4.2.
@@
 $=\int_{0}^{\infty}\mathbb{1}(v-Y(t)>z)F^{*(n-1)}\left(v-Y(t)-\mathrm{d}z\right)\frac{\lambda^{n}t^{n-1}\mathrm{e}^{-\left(q+\lambda\right)t}}{(n-1)!}\mathrm{d}t.$
-Substituting the above equation into equation (E) yields
+Substituting the above equation into equation (F) yields
 $\mathbb{E}_{v}\left[\mathrm{e}^{-q\tau_{J}}\mathbb{1}\left(\tau_{J}<\tau_{Y,}\tau_{J}=T_{n}\right)\right]$
@@
 $F_{Z}^{*(n)}(z)$ $=1-\sum_{i=0}^{n-1}\frac{1}{i!}\mathrm{e}^{-\alpha z}(\alpha z)^{i}.$
-Summing up equation (E) over $n$ leads to completes the proof of equation (13).
+Summing up equation (F) over $n$ leads to completes the proof of equation (13).
 Now, suppose that the death occurs due to the decrement in $Y(\cdot)$ rather than a fatal accident, i.e., $\tau=\tau_{Y}<\tau_{J}$.
@@
 for $n\geq 1$, where the third equation follows from interchanging the order of integrations, while the last equation is derived by changing of variable (letting $\frac{\ln\left(bc^{x}+z\ln c\right)-\ln b}{\ln c}-x=t$).
-Summing up equation (E) over $n$ ($n\geq 1$) and noting that
-$\mathbb{E}_{v}\left[\mathrm{e}^{-q\tau_{Y}}\mathbf{1}\left(\tau_{Y}<\tau_{J,}N(\tau_{Y})=0\right)\right]=\mathrm{e}^{-(q+\lambda)t_{v}^{*}}$
+Summing up equation (F) over $n$ ($n\geq 1$) and noting that
+$\mathbb{E}_{v}\left[\mathrm{e}^{-q\tau_{Y}}\mathbb{1}\left(\tau_{Y}<\tau_{J,}N(\tau_{Y})=0\right)\right]=\mathrm{e}^{-(q+\lambda)t_{v}^{*}}$
 completes that proof of equation (14).